### v1

# Galois Self-dual 2-quasi Constacyclic Codes over Finite Fields

Abstract
[…]
Key words: Finite fields; $2$-quasi constacyclic codes; Galois self-dual codes; Hermitian self-dual codes; asymptotic property.
Email address: yfan@mail.ccnu.edu.cn (Yun Fan);
1 Introduction
Let $F$ be a finite field with cardinality $|F|=q=p^{\ell}$ where $p$ is a prime and $\ell$ is a positive integer, let $F^{\times}=F\setminus\{0\}$. Any ${\bf a}=(a_{0},a_{1},\dots,a_{n-1})\in{F}^{n}$ is called a word over $F$ of length $n$, where $n$ is a positive integer. The Hamming weight ${\rm w}({\bf a})$ is defined as the number of the indexes $i$ with $a_{i}\neq 0$. The Hamming distance between two words ${\bf a},{\bf a}^{\prime}\in F^{n}$ is defined as ${\rm d}({\bf a},{\bf a}^{\prime})={\rm w}({\bf a}-{\bf a}^{\prime})$. Any nonempty subset $C\subseteq F^{n}$ is called a code of length $n$ over $F$, and the words in the code are called codewords. The minimum distance ${\rm d}(C):=\min\limits_{{\bf c}\neq{\bf c}^{\prime}\in C}{\rm d}({\bf c},{\bf c}^{\prime})$. For any linear subspace $C$ of $F^{n}$, called a linear code, the minimum weight ${\rm w}(C):=\min\limits_{0\neq{\bf c}\in C}{\rm w}({\bf c})$; and it is known that ${\rm w}(C)={\rm d}(C)$. The fraction $\Delta(C)=\frac{{\rm d}(C)}{n}=\frac{{\rm w}(C)}{n}$ is called the relative minimum distance of $C$, and ${\rm R}(C)=\frac{\dim_{F}C}{n}$ is called the rate of $C$. A code sequence $C_{1},C_{2},\dots$ is said to be asymptotically good if the length $n_{i}$ of $C_{i}$ goes to infinity and there is a real number $\delta>0$ such that ${\rm R}(C_{i})>\delta$ and $\Delta(C_{i})>\delta$ for $i=1,2,\dots$. A class of codes is said to be asymptotically good if there is an asymptotically good sequence of codes in the class; otherwise, we say that the class of codes is asymptotically bad.
A linear code $C$ of $F^{n}$ is called a cyclic code if $C$ is invariant under the cyclic permutation on items, i.e.,
$(c_{0},c_{1},\ldots,c_{n-1})\in{C}\implies(c_{n-1},c_{0},c_{1},\dots,c_{n-2})\in{C}.$
A linear code $C$ of $F^{n}\times F^{n}$ is called a quasi-cyclic code of index $2$, abbreviated as $2$-quasi-cyclic code, if $C$ is invariant under the double cyclic permutation on items, i.e.,
$\begin{array}[]{l}(c_{0},c_{1},\dots,c_{n-1},c^{\prime}_{0},c^{\prime}_{1},\dots,c^{\prime}_{n-1})\in{C}\\[1.0pt] \implies(c_{n-1},c_{0},c_{1},\dots,c_{n-2},\,c^{\prime}_{n-1},c^{\prime}_{0},c^{\prime}_{1},\dots,c^{\prime}_{n-2})\in{C}.\end{array}$
The Euclidean inner product of words ${\bf a}=(a_{0},\dots,a_{n-1}),{\bf b}=(b_{0},\dots,b_{n-1})$ of $F^{n}$ is defined to be $\langle{\bf a},{\bf b}\rangle=\sum_{i=0}^{n-1}a_{i}b_{i}$. For a code $C\subseteq F^{n}$, the $C^{\bot}=\{\,a\in{F^{n}\,|\,\langle c,a\rangle=0,\,\forall c\in{C}}\,\}$ is the dual code of $C$. A code $C$ is said to be self-dual if $C=C^{\bot}$. Obviously, the rate $R(C)=\frac{1}{2}$ if $C$ is self-dual.
[…]
Cyclic codes and $2$-quasi-cyclic codes had been extended widely. Let $\lambda\in F^{\times}=F\setminus\{0\}$. A linear code $C$ of $F^{n}$ is called a $\lambda$-constacyclic code if $C$ is invariant under the $\lambda$-constacyclic permutation on items, i.e.,
$(c_{0},c_{1},\ldots,c_{n-1})\in C\implies(\lambda c_{n-1},c_{0},c_{1},\dots,c_{n-2})\in C.$ (1.1)
If $\lambda=-1$, the $\lambda$-constacyclic codes are called negacyclic codes. Further, a linear code $C$ of $F^{n}\times F^{n}$ is called a $2$-quasi $\lambda$-constacyclic code if $C$ is invariant under the double $\lambda$-constacyclic permutation on items, i.e.,
$\begin{array}[]{l}(c_{0},c_{1},\dots,c_{n-1},c^{\prime}_{0},c^{\prime}_{1},\dots,c^{\prime}_{n-1})\in{C}\\[1.0pt] \implies(\lambda c_{n-1},c_{0},c_{1},\dots,c_{n-2},\,\lambda c^{\prime}_{n-1},c^{\prime}_{0},c^{\prime}_{1},\dots,c^{\prime}_{n-2})\in{C}.\end{array}$ (1.4)
If $q$ is odd and $q\,{\not\equiv}\pm 1~{}({\rm mod}~{}8)$, the self-dual $2$-quasi negacyclic codes over $F$ are proved asymptotically good in [2]. While in [25] for $q\,{\equiv}-1~{}({\rm mod}~{}4)$ it is shown, based on Artin’s primitive root conjecture, that the $q$-ary self-dual $2$-quasi negacyclic codes are asymptotically good. Recently, for any $q$ and any $\lambda\in F^{\times}$ the $2$-quasi $\lambda$-constacyclic codes over $F$ are proved asymptotically good, see [10, Corollary I.3].
About the self-dualities, in the semisimple case (i.e., $\gcd(n,q)=1$), the self-dual cyclic codes over $F$ does not exist. Leon et al. [18] and many references, e.g. [7, 22, 26], devoted to the study on various generalizations, e.g., duadic codes, extended self-dual cyclic codes, etc. On the other hand, Dinh and Lopez-Permouth [9], Dinh [8] studied $\lambda$-constacyclic codes, and showed that in the semisimple case the self-dual $\lambda$-constacyclic codes exist only if $\lambda=-1$.
Extending the Euclidean inner product and the Hermitian inner product, Fan and Zhang [13] introduced the so-called Galois inner products. Recall that $|F|=q=p^{\ell}$. For $0\leq h<\ell$, the map $\sigma_{p^{h}}:F\to F$, $\alpha\mapsto\alpha^{p^{h}}$, is a Galois automorphism of $F$, which induces an automorphism $F^{n}\to F^{n}$, $(a_{0},\dots,a_{n-1})\mapsto(a_{0}^{p^{h}},\dots,a_{n-1}^{p^{h}})$. The following
$\langle{\bf a},{\bf b}\rangle_{h}=\sum_{i=0}^{n-1}a_{i}b_{i}^{p^{h}},~{}~{}~{}~{}\forall\,{\bf a}=(a_{0},\cdots,a_{n-1}),{\bf b}=(b_{0},\cdots,b_{n-1})\in F^{n},$ (1.5)
is called the Galois $p^{h}$-inner product on $F^{n}$. And for any code $C\subseteq F^{n}$, the following code
$C^{\bot{h}}=\big{\{}\,{\bf a}\in F^{n}\,\big{|}\,\langle{\bf c},{\bf a}\rangle_{h}=0,\;\forall\,{\bf c}\in C\big{\}}$ (1.6)
is called the Galois $p^{h}$-dual code of $C$. The code $C$ is said to be Galois $p^{h}$-self-dual (or Galois self-dual when $h$ is known from context) if $C=C^{\bot h}$. It is also obvious that $R(C)=\frac{1}{2}$ if $C$ is Galois $p^{h}$-self-dual. When $h=0$ ($h=\ell/2$ when $\ell$ is even, respectively), $\langle{\bf a},{\bf b}\rangle_{h}$ is just the Euclidean (Hermitian, respectively) inner product, and $C^{\bot{h}}$ is the Euclidean (Hermitian, respectively) dual code, and Galois $p^{h}$-self-dual codes are just the Euclidean self-dual (Hermitian self-dual, respectively) codes. The existence and the structure of Galois $p^{h}$-self-dual $\lambda$-constacyclic codes are studied in [13].
In this paper we study the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ and their asymptotic properties. The main contributions of this paper are the following.
  • We characterize the algebraic structure of the $2$-quasi $\lambda$-constacyclic codes and their Galois $p^{h}$-dual codes. We find that the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes behave very differently depending on whether ${\lambda^{1+p^{h}}=1}$ or not. In both the cases we obtain necessary and sufficient conditions for $2$-quasi $\lambda$-constacyclic codes being Galois $p^{h}$-self-dual.
[…]
In Section 2, some preliminaries are sketched.
In Section 3 we characterize the algebraic structure of the $2$-quasi $\lambda$-consta-cyclic codes over the finite field $F$ and their Galois $p^{h}$-dual codes.
In Section 4 we study the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$. Our discussion divide into two cases: $\lambda^{1+p^{h}}\!\neq 1$ or $\lambda^{1+p^{h}}\!=1$. In both the cases, we exhibit the necessary and sufficient conditions for a $2$-quasi $\lambda$-constacyclic code being Galois $p^{h}$-self-dual. And we show that if $\lambda^{1+p^{h}}\!\neq 1$, then Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes are asymptotically bad.
In Section 5, the Hermitian self-dual $2$-quasi-cyclic codes over $F$ are proved asymptotically good.
[…]
Finally, we end this paper by a conclusion in Section 7.
2 Preliminaries
In this paper, $F$ is always a finite field of cardinality $|F|=q=p^{\ell}$ (by $|S|$ we denote the cardinality of any set $S$), where $p$ is a prime and $\ell$ is a positive integer; and $h$ is an integer such that $0\leq h<\ell$; and $n>1$ is an integer. Any ring $R$ in this paper has identity $1_{R}$ (or denoted by $1$ for short); and ring homomorphisms and subrings are identity preserving. By $R^{\times}$ we denote the multiplication group consisting of all units (invertible elements) of $R$. In particular, $F^{\times}=F\,{\setminus}\,\{0\}$. If a ring $R$ is also an $F$-vector space, then $R$ is said to be an $F$-algebra. In that case, $F\to R$, $\alpha\mapsto\alpha 1_{R}$, is an embedding, so that we write that $F\subseteq R$.
Let $R$ and $S$ be $F$-algebras. A map $\psi:R\to S$ is called an $F$-algebra homomorphism if it is both a ring homomorphism and an $F$-linear map, i.e., $\psi(1_{R})=1_{S}$ and for any $a,b\in R$ and any $\alpha\in F$,
$\psi(ab)=\psi(a)\psi(b),~{}~{}\psi(a+b)=\psi(a)+\psi(b),~{}~{}\psi(\alpha a)=\alpha\psi(a),$
where the first two mean that $\psi$ is a ring homomorphism, and the last two mean that it is a linear map. Recall that for $0\leq h<\ell$, the map $\sigma_{p^{h}}:F\to F$, $\sigma_{p^{h}}(\alpha)=\alpha^{p^{h}}$ for $\alpha\in F$, is a Galois automorphism of the field $F$. If $\psi:R\to S$ is bijective and satisfies that for any $a,b\in R$ and any $\alpha\in F$,
$\psi(ab)=\psi(a)\psi(b),~{}~{}\psi(a+b)=\psi(a)+\psi(b),~{}~{}\psi(\alpha a)=\alpha^{p^{h}}\psi(a),$ (2.1)
then $\psi$ is called a $\sigma_{p^{h}}$-algebra isomorphism, or $p^{h}$-isomorphism for short. Note that the last two equalities of Eq.(2.1) mean that $\psi$ is a $p^{h}$-linear map. In particular, if $R=S$, then $\psi$ is called a $\sigma_{p^{h}}$-algebra automorphism, or a $p^{h}$-automorphism for short. And, if $\psi:R\to S$ is a $p^{h}$-isomorphism, then for any ideal $I$ of $R$, the image $\psi(I)$ is an ideal of $S$ and $\dim_{F}\psi(I)=\dim_{F}I$. There are two typical examples as follows.
Example 2.1.
(1). For any polynomial $f(X)=\sum_{i}a_{i}X^{i}\in F[X]$, we denote $f^{(p^{h})}(X)=\sum_{i}a_{i}^{p^{h}}X^{i}$. The Galois automorphism $\sigma_{p^{h}}:F\to F$, $\alpha\mapsto\alpha^{p^{h}}$, induces the map
$\tilde{\sigma}_{p^{h}}\!:~{}F[X]\,\longrightarrow\,F[X],~{}~{}~{}~{}f(X)\,\longmapsto\,f^{(p^{h})}(X),$ (2.2)
which is a $p^{h}$-automorphism of $F[X]$. Let ${\rm ord}(\sigma_{p^{h}})$ denote the order of $\sigma_{p^{h}}$. It is easy to check that ${\rm ord}(\tilde{\sigma}_{p^{h}})={\rm ord}(\sigma_{p^{h}})=\frac{\ell}{\gcd(h,\ell)}$.
(2). For the matrix algebra ${\rm M}_{n}(F)=\big{\{}(a_{ij})_{n\times n}\,\big{|}\,a_{ij}\in F\big{\}}$ consisting of all $n\times n$ matrices over $F$, the map
$\dot{\sigma}_{p^{h}}\!:{\rm M}_{n}(F)\to{\rm M}_{n}(F),~{}~{}~{}~{}\big{(}a_{ij}\big{)}_{n\times n}\longmapsto\big{(}a_{ij}^{p^{h}}\big{)}_{n\times n},$ (2.3)
is a $p^{h}$-automorphism of ${\rm M}_{n}(F)$.
Let $\lambda\in F^{\times}$. For a positive integer $k$, we write $E_{k}$ to denote the identity matrix of degree $k$. Let $P_{\lambda}$ denote the $\lambda$-constacyclic permutation matrix of degree $n$ as follows
$P_{\lambda}=\begin{pmatrix}&E_{n-1}\\ \lambda E_{1}\end{pmatrix}=\begin{pmatrix}0&1\\ &0&\ddots\\ &&\ddots&\ddots\\ &&&\ddots&1\\ \lambda&&&&0\end{pmatrix}.$ (2.4)
In particular, if $\lambda=1$ then $P:=P_{1}=\begin{pmatrix}&E_{n-1}\\ E_{1}\end{pmatrix}$ is the cyclic permutation matrix. By matrix multiplication, for $\mathbf{a}=(a_{0},a_{1},\dots,a_{n-1})\in{F^{n}}$ we have that
$\mathbf{a}\cdot P_{\lambda}=(a_{0},a_{1},a_{2},\dots,a_{n-1})\cdot P_{\!\lambda}=(\lambda a_{n-1},a_{0},a_{1},\dots,a_{n-2}),$ (2.5)
which is the vector obtained by $\lambda$-constacyclically permuting the items of the vector $(a_{0},a_{1},\dots,a_{n-1})$; and that for $\mathbf{a}=(a_{0},a_{1},\dots,a_{n-1})\in{F^{n}}$ and $\mathbf{a^{\prime}}=(a^{\prime}_{0},a^{\prime}_{1},\dots,a^{\prime}_{n-1})\in{F^{n}}$,
$\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}\begin{pmatrix}P_{\lambda}\!\\ &\!P_{\lambda}\end{pmatrix}$ $=(\lambda a_{n-1},a_{0},a_{1},\dots,a_{n-2},\,\lambda a^{\prime}_{n-1},a^{\prime}_{0},a^{\prime}_{1},\dots,a^{\prime}_{n-2}).$ (2.6)
Thus, we get another description of the $\lambda$-constacyclic codes (cf. Eq.(1.1)) and the $2$-quasi $\lambda$-constacyclic codes (cf. Eq.(1.4)) as follows.
Lemma 2.2.
(1) Let $C$ be a subspace of $F^{n}$. Then $C$ is a $\lambda$-constacyclic code if and only if $\mathbf{c}\cdot P_{\lambda}\in C$, for any $\mathbf{c}\in C$.
(2) Let $C$ be a subspace of $F^{n}\times F^{n}$. Then $C$ is a $2$-quasi $\lambda$-constacyclic code if and only if $\Big{(}\mathbf{c},\,\mathbf{c}^{\prime}\big{)}\begin{pmatrix}P_{\lambda}\!\\ &\!P_{\lambda}\end{pmatrix}\in C$, for any $\big{(}\mathbf{c},\,\mathbf{c}^{\prime}\big{)}\in C$. ∎
In the following we always denote ${\cal R}_{\kern-1.0pt\lambda}=F[X]\big{/}\langle X^{n}-\lambda\rangle$, which is the quotient algebra of the polynomial algebra $F[X]$ over the ideal $\langle X^{n}-\lambda\rangle$ generated by $X^{n}-\lambda$. Any residue class modulo $X^{n}-\lambda$ has a unique representative polynomial with degree less than $n$. Hence we can write
${\cal R}_{\kern-1.0pt\lambda}=F[X]\big{/}\langle X^{n}-\lambda\rangle=\big{\{}\,a_{0}+a_{1}X+\dots+a_{n-1}X^{n-1}\;\big{|}\;a_{i}\in F\,\big{\}}.$ (2.7)
Further, the Cartesian product
${\cal R}_{\kern-1.0pt\lambda}^{2}={\cal R}_{\kern-1.0pt\lambda}\times{\cal R}_{\kern-1.0pt\lambda}=\big{\{}\big{(}a(X),a^{\prime}(X)\big{)}\;\big{|}\;a(X),~{}a^{\prime}(X)\in{\cal R}_{\kern-1.0pt\lambda}\big{\}}$ (2.8)
is an ${\cal R}_{\kern-1.0pt\lambda}$-module. For ${\cal R}_{\kern-1.0pt\lambda}$ and ${\cal R}_{\kern-1.0pt\lambda}^{2}$, the following identifications and results will be quoted later in this article.
Remark 2.3.
(1). There is a canonical linear isomorphism
$\iota:~{}~{}{{\cal R}_{\kern-1.0pt\lambda}}~{}\mathop{\longrightarrow}^{\cong}~{}F^{n},\qquad a(X)~{}\longmapsto~{}\mathbf{a},$ (2.9)
where $a(X)=a_{0}+a_{1}X+\dots+a_{n-1}X^{n-1}\in{\cal R}_{\kern-1.0pt\lambda}$ and $\mathbf{a}=(a_{0},a_{1},\dots,a_{n-1})\in F^{n}$. It is easy to check that
$\iota\big{(}Xa(X)\big{)}={\bf a}\cdot P_{\lambda},\qquad\forall\;a(X)\in{\cal R}_{\kern-1.0pt\lambda}.$ (2.10)
Then any element $a(X)$ of ${\cal R}_{\kern-1.0pt\lambda}$ is identified with the word $\mathbf{a}=(a_{0},a_{1},\dots,a_{n-1})$ of $F^{n}$; and by Lemma 2.2(1), the $\lambda$-constacyclic codes of length $n$ are identified with the ideals (${\cal R}_{\kern-1.0pt\lambda}$-submodules) of ${\cal R}_{\kern-1.0pt\lambda}$.
(2). For the ${\cal R}_{\kern-1.0pt\lambda}$-module ${\cal R}_{\kern-1.0pt\lambda}^{2}={\cal R}_{\kern-1.0pt\lambda}\times{\cal R}_{\kern-1.0pt\lambda}$, we have the following canonical linear isomorphism
$\iota^{(2)}:~{}~{}{{\cal R}_{\kern-1.0pt\lambda}}\times{{\cal R}_{\kern-1.0pt\lambda}}~{}\mathop{\longrightarrow}^{\cong}~{}F^{n}\times F^{n},\qquad\big{(}a(X),\,a^{\prime}(X)\big{)}~{}\longmapsto~{}\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)},$ (2.11)
where $a(X)=a_{0}+a_{1}X+\dots+a_{n-1}X^{n-1}$, $a^{\prime}(X)=a^{\prime}_{0}+a^{\prime}_{1}X+\dots+a^{\prime}_{n-1}X^{n-1}$, and $(\mathbf{a},\mathbf{a}^{\prime})=(a_{0},a_{1},\dots,a_{n-1},\,a^{\prime}_{0},a^{\prime}_{1},\dots,a^{\prime}_{n-1})$. For $\big{(}a(X),a^{\prime}(X)\big{)}\in{\cal R}_{\kern-1.0pt\lambda}^{2}$,
$\iota^{(2)}\big{(}X\big{(}a(X),a^{\prime}(X)\big{)}\big{)}=\iota^{(2)}\big{(}Xa(X),Xa^{\prime}(X)\big{)}=\big{(}{\bf a},{\bf a}^{\prime}\big{)}\begin{pmatrix}P_{\lambda}\!\\ &\!P_{\lambda}\end{pmatrix}.$ (2.12)
Then any element $\big{(}a(X),a^{\prime}(X)\big{)}\in{{{\cal R}_{\kern-1.0pt\lambda}}\times{{\cal R}_{\kern-1.0pt\lambda}}}$ is identified with the word $\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}\in{F^{n}\times F^{n}}$, and by Lemma 2.2(2) the $2$-quasi $\lambda$-constacyclic codes of length $2n$ are identified with the ${{\cal R}_{\kern-1.0pt\lambda}}$-submodules of ${\cal R}_{\kern-1.0pt\lambda}^{2}$.
Remark 2.4.
If $\gcd(n,q)=1$, then the algebra ${\cal R}_{\kern-1.0pt\lambda}$ is semisimple, cf. [4]; and by Ring Theory (e.g., cf. [16] or [14, Remark 2.4]), we have the following two.
(1) For any ideal (${\cal R}_{\kern-1.0pt\lambda}$-submodule) $C$ of ${{\cal R}_{\kern-1.0pt\lambda}}$, there is an idempotent $e_{C}$ of ${{\cal R}_{\kern-1.0pt\lambda}}$ such that $C={\cal R}_{\kern-1.0pt\lambda}e_{C}$ and ${{\cal R}_{\kern-1.0pt\lambda}}=C\oplus D$ where $D={{\cal R}_{\kern-1.0pt\lambda}}(1-e_{C})$. Note that $C={\cal R}_{\kern-1.0pt\lambda}e_{C}$ is an algebra with identity $e_{C}$ (but not a subalgebra of ${\cal R}_{\kern-1.0pt\lambda}$ in general because $e_{C}\neq 1$ in general); in particular, $C^{\times}$ makes sense. Moreover, if $I={\cal R}_{\kern-1.0pt\lambda}f$ with $f$ being an idempotent and an ideal $C\subseteq I$, then $I=C\oplus C^{\prime}$ with $C^{\prime}={\cal R}_{\kern-1.0pt\lambda}(f-e_{C})$.
(2) If $C$ and $C^{\prime}$ are ${\cal R}_{\kern-1.0pt\lambda}$-submodules of ${\cal R}_{\kern-1.0pt\lambda}$; and $\varphi:C\to C^{\prime}$ is an ${\cal R}_{\kern-1.0pt\lambda}$-module isomorphism, then $C^{\prime}=C$, and there is a $g\in C^{\times}$ such that $\varphi(c)=cg$ for any $c\in C$.
If $\lambda=1$, there is another identification. By Eq.(2.7), we denote
${\cal R}={\cal R}_{1}=F[X]/\langle X^{n}-1\rangle.$ (2.13)
Let $G=\langle\,x\,|\,x^{n}=1\,\rangle=\{1,x,\dots,x^{n-1}\}$ be the cyclic group of order $n$. Let $FG$ be the cyclic group algebra, i.e., $FG=\big{\{}\sum_{i=0}^{n-1}a_{i}x^{i}\,\big{|}\,a_{i}\in F\big{\}}$ is an $F$-vector space with basis $G$ and equipped with the multiplication induced by the multiplication of the group $G$ as follows:
$\Big{(}\sum_{i=0}^{n-1}a_{i}x^{i}\Big{)}\Big{(}\sum_{j=0}^{n-1}b_{j}x^{j}\Big{)}=\sum_{k=0}^{n-1}\Big{(}\sum_{x^{i}x^{j}=x^{k}}a_{i}b_{j}\Big{)}x^{k}.$
There is a canonical algebra isomorphism:
${\cal R}=F[X]/\langle X^{n}-1\rangle\;\mathop{\longrightarrow}^{\cong}\;FG,~{}~{}~{}~{}\sum_{i=0}^{n-1}a_{i}X^{i}\;\longmapsto\;\sum_{i=0}^{n-1}a_{i}x^{i}.$ (2.14)
Thus, ${\cal R}$ is identified with the cyclic group algebra $FG=\{\sum_{i=0}^{n-1}a_{i}x^{i}\,|\,a_{i}\in F\}$. And by Remark 2.3(1), the cyclic codes over $F$ of length $n$ are identified with the ideals of the cyclic group algebra $FG$. Similarly, $2$-quasi-cyclic codes over $F$ of length $2n$ are identified with the $FG$-submodules of the $FG$-module $(FG)^{2}=FG\times FG$.
With the identifications Eq.(2.14), we have more algebraic preliminaries about ${\cal R}$ to introduce. Assume that $\gcd(n,q)=1$, then ${\cal R}$ is semisimple. Let
$e_{0}=\frac{1}{n}\sum_{i=0}^{n-1}x^{i},~{}e_{1},~{}\dots,~{}e_{m}$ (2.15)
be all primitive idempotents of ${\cal R}$. Correspondingly, the irreducible decomposition of $X^{n}-1$ in $F[X]$ is as follows
$X^{n}-1=\varphi_{0}(X)\varphi_{1}(X)\dots\varphi_{m}(X),\quad\mbox{where }\varphi_{0}(X)=X-1,$ (2.16)
such that
$\begin{array}[]{l}{\cal R}={\cal R}e_{0}\oplus{\cal R}e_{1}\oplus\dots\oplus{\cal R}e_{m};\\[3.0pt] {\cal R}e_{i}\cong F[X]/\varphi_{i}(X),\quad i=0,1,\dots,m.\end{array}$ (2.19)
Since $\varphi_{i}(X)$ is irreducible over $F$, each ${\cal R}e_{i}$ is an extension field over $F$ with identity $e_{i}$, and $\dim_{F}{\cal R}e_{i}=\deg\varphi_{i}(X)$. As $e_{0}=\frac{1}{n}\sum_{i=0}^{n-1}x^{i}$, $xe_{0}=e_{0}$; so
${\cal R}e_{0}=\!\big{\{}\alpha\!+\!\alpha x+\dots+\!\alpha x^{n-1}\,\big{|}\,\alpha\!\in\!F\big{\}},~{}\dim_{F}{\cal R}e_{0}=1,~{}{\rm w}({\cal R}e_{0})=n.$ (2.20)
For $1\leq i\leq m$, we denote
$\mu(n):=\min_{1\leq i\leq m}\deg\varphi_{i}(X)=\min_{1\leq i\leq m}\dim_{F}{\cal R}e_{i}.$ (2.21)
Remark 2.5.
[…]
$h_{q}(\delta)=\delta\log_{q}(q\!-\!1)-\delta\log_{q}(\delta)-(1\!-\!\delta)\log_{q}(1\!-\!\delta),~{}~{}~{}\delta\in[0,1\!-\!q^{-1}],$ (2.22)
which value strictly increases from 00 to $1$ while $\delta$ increases from 00 to $1-q^{-1}$. By [20, Theorem IV.17], for any real number $\delta$ with $0<\delta<1-q^{-1}$ and $h_{q}(\delta)<\frac{1}{4}$, there are self-dual $2$-quasi-cyclic codes $C_{1},C_{2},\dots$ over $F$ (hence ${\rm R}(C_{i})=\frac{1}{2}$) such that: (1) the relative minimum distance $\Delta(C_{i})>\delta$  for $i=1,2,\dots$; and (2) the code length $2n_{i}$ of $C_{i}$ satisfies that every $n_{i}$ is odd and coprime to $q$, and $\lim\limits_{i\to\infty}\frac{\log_{q}(n_{i})}{\mu(n_{i})}=0$ (in particular, $2n_{i}\to\infty$).
Definition 2.6.
Let $C\subseteq F^{n}=F^{I}$ with $I=\{0,1,\dots,n-1\}$ being the index set. If there are positive integers $s,k,t$ and subsets (repetition is allowed) $I_{1},\dots,I_{s}$ of $I$ with $|I_{j}|=k$ for $j=1,\dots,s$ satisfying the following two conditions: (1) for each $I_{j}=\{i_{1},\cdots,i_{k}\}$ ($1\leq j\leq s$) the projection $\rho_{I_{j}}$: $F^{I}~{}\to~{}F^{I_{j}}$, $(a_{0},a_{1},\dots,a_{n-1})\mapsto(a_{i_{1}},\dots,a_{i_{k}})$, maps $C$ bijectively onto $F^{I_{j}}$; and (2) for any $i$ ($0\leq i<n$) the number of the subsets $I_{j}$ which contains $i$ (i.e., $i\in I_{j}$) equals $t$; then we say that $C$ is a balanced code over $F$ of length $n$, and $I_{1},\dots,I_{s}$ are called information index sets of the code $C$.
An important result (see [11, Corollary 3.4]) is that: if $C\subseteq F^{n}$ is a balance code with cardinality $|C|=q^{k}$, then
$|C^{\leq\delta}|\leq q^{kh_{q}(\delta)},~{}~{}~{}\mbox{for}~{}~{}0\leq\delta\leq 1-q^{-1},$
where $h_{q}(\delta)$ is defined in Eq.(2.22) and
$C^{\leq\delta}=\{\,{\bf c}\,|\,{\bf c}\in C,\,{{\rm w}({\bf c})}\leq\delta n\,\}.$ (2.23)
Constacyclic codes are balanced, see [10, Lemma II.8]. By [11, Corollary 3.4 and Corollary 3.5], we can easily obtain the following lemma.
Lemma 2.7.
If $A$ is an ideal of ${\cal R}_{\kern-1.0pt\lambda}$, then the ${\cal R}_{\kern-1.0pt\lambda}$-submodule $A\times A$ of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ is a balanced code, hence for $0\leq\delta\leq 1-q^{-1}$,
$|(A\times A)^{\leq\delta}|\leq q^{2\dim_{F}(A)\cdot h_{q}(\delta)}.$
3 $2$-quasi constacyclic codes over finite fields
In this section, we are primarily concerned with the algebra properties of $2$-quasi $\lambda$-constacyclic codes over $F$. In the following, we always assume that
$\lambda\in F^{\times}~{}~{}~{}\mbox{and}~{}~{}~{}t={\rm ord}_{F^{\times}}(\lambda),$ (3.1)
where ${\rm ord}_{F^{\times}}(\lambda)$ denotes the order of $\lambda$ in the multiplication group $F^{\times}$.
Remark 3.1.
For $F^{n}\times F^{n}$, there are two projections $\rho_{1}$, $\rho_{2}$: $F^{n}\times F^{n}\to F^{n}$ as follows
$\rho_{1}({\bf a},{\bf a}^{\prime}\big{)}={\bf a},~{}~{}~{}\rho_{2}({\bf a},{\bf a}^{\prime})={\bf a}^{\prime},~{}~{}~{}~{}~{}~{}\forall~{}({\bf a},\,{\bf a}^{\prime})\in F^{n}\times F^{n}.$
By the linear isomorphism Eq.(2.11), the projections $\rho_{1}$ and $\rho_{2}$ are also defined on ${\cal R}_{\kern-1.0pt\lambda}^{2}={\cal R}_{\kern-1.0pt\lambda}\times{\cal R}_{\kern-1.0pt\lambda}$: for $\big{(}a(X),\,a^{\prime}(X)\big{)}\in{\cal R}_{\kern-1.0pt\lambda}^{2}$,
$\rho_{1}\big{(}a(X),\,a^{\prime}(X)\big{)}=a(X),~{}~{}~{}~{}~{}\rho_{2}\big{(}a(X),\,a^{\prime}(X)\big{)}=a^{\prime}(X).$
For any ${\cal R}_{\kern-1.0pt\lambda}$-submodule $C$ of ${{\cal R}_{\kern-1.0pt\lambda}}\times{{\cal R}_{\kern-1.0pt\lambda}}$, restricting $\rho_{1}$ to $C$, we have an ${{\cal R}_{\kern-1.0pt\lambda}}$-homomorphism $\rho_{1}|_{C}:C\to\rho_{1}(C)$. Observe that the kernel of the restricted homomorphism is
${\rm Ker}(\rho_{1}|_{C})=C\cap(0\times{{\cal R}_{\kern-1.0pt\lambda}})=\big{\{}\big{(}0,c^{\prime}(X)\big{)}\in C\,\big{|}\,c^{\prime}(X)\in{{{\cal R}_{\kern-1.0pt\lambda}}}\big{\}}.$ (3.2)
It is known that for $\lambda_{1}\neq\lambda_{2}\in F^{\times}$, if a linear code $C\subseteq F^{n}$ is both $\lambda_{1}$-constacyclic and $\lambda_{2}$-constacyclic, then either $C=0$ or $C=F^{n}$ (cf. [8]). Extending the result to $F^{n}\times F^{n}$, we get the following lemma.
Lemma 3.2.
Let $\lambda_{1},\lambda_{2}\in F^{\times}$ such that $\lambda_{1}\neq\lambda_{2}$. If a subspace $C$ of $F^{n}\times F^{n}$ is both a $2$-quasi $\lambda_{1}$-constacyclic code and a $2$-quasi $\lambda_{2}$-constacyclic code, then either $\rho_{1}(C)=0$ or $\rho_{1}(C)=F^{n}$; and it is the same for $\rho_{2}(C)$.
Proof.
Suppose $\rho_{1}(C)\neq 0$, there is a codeword
$({\bf c},\,{\bf c}^{\prime})=(c_{0},\dots,c_{n-1},\;c^{\prime}_{0},\dots,c^{\prime}_{n-1})\,\in\,C$
with some $c_{j}\neq 0$, $j=0,1,\dots,n-1$; so we can assume that $c_{n-1}\neq 0$. By the double $\lambda_{1}$-constacyclic permutation in Eq.(1.4), then $C$ contains the following word
$\frac{1}{(\lambda_{1}-\lambda_{2})c_{n-1}}\Big{(}\big{(}\mathbf{c}P_{\!\lambda_{1}},\,\mathbf{c}^{\prime}P_{\!\lambda_{1}}\big{)}-\big{(}\mathbf{c}P_{\!\lambda_{2}},\,\mathbf{c}^{\prime}P_{\!\lambda_{2}}\big{)}\Big{)}=\big{(}1,0,\dots,0,\,d^{\prime}_{0},\dots,d^{\prime}_{n-1}\big{)},$
for some $d^{\prime}_{i}\in{F}$, $i=0,1,\dots,n-1$. Using the double $\lambda_{1}$-constacyclic permutation $i$ times, we see that $C$ contains such words
$\big{(}\mathop{0,\dots,0,\,1,\,0,\dots,0,}\limits_{\mbox{\scriptsize index\,$i$}}\,b^{\prime}_{0},\dots,b^{\prime}_{n-1}\big{)},\qquad i=0,1,\dots,n-1,$
for some $b^{\prime}_{i}\in{F}$, $i=0,1,\dots,n-1$. It follows that $\rho_{1}(C)$ contains a basis of $F^{n}$, hence $\rho_{1}(C)=F^{n}$.
By the same argument, either $\rho_{2}(C)=0$ or $\rho_{2}(C)=F^{n}$. ∎
Extending [14, Theorem 3.2] which characterized the algebraic structure of $2$-quasi-cyclic codes in the semisimple case, in that case we characterize the algebraic structure of $2$-quasi constacyclic codes as follows.
Theorem 3.3.
Assume that $\gcd(n,q)=1$. If $C$ is an ${\cal R}_{\kern-1.0pt\lambda}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$, then there are ideals $C_{1},C_{2},C_{12}$ of ${{\cal R}_{\kern-1.0pt\lambda}}$ satisfying that $C_{1}\cap C_{12}=C_{2}\cap C_{12}=0$ and an element $g\in C_{12}^{\times}$ such that
$C=(C_{1}\times C_{2})\oplus\widehat{C}_{12},~{}~{}\mbox{where}~{}~{}\widehat{C}_{12}=\big{\{}\big{(}c_{12},c_{12}g\big{)}\,\big{|}\,c_{12}\in C_{12}\big{\}}\cong C_{12}.$ (3.3)
Conversely, if there are ideals $C_{1},C_{2},C_{12}$ of ${{\cal R}_{\kern-1.0pt\lambda}}$ with $C_{1}\cap C_{12}=C_{2}\cap C_{12}=0$ and an element $g\in C_{12}^{\times}$, then $C$ in Eq.(3.3) is an ${\cal R}_{\kern-1.0pt\lambda}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$.
Proof.
The “conversely” part is obviously true because both $C_{1}\times C_{2}$ and $\widehat{C}_{12}$ are ${\cal R}_{\kern-1.0pt\lambda}$-submodules of ${\cal R}_{\kern-1.0pt\lambda}^{2}$, and $(C_{1}\times C_{2})\cap\widehat{C}_{12}=0$.
Assume that $C$ is an ${\cal R}_{\kern-1.0pt\lambda}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$. We consult the module version of Goursat Lemma (see [14, Remark 3.1]). Take
$\begin{array}[]{l}\tilde{C_{1}}=\rho_{1}(C)=\big{\{}a_{1}\in{{{\cal R}_{\kern-1.0pt\lambda}}}~{}|~{}(a_{1},a_{2})\in{C}~{}\mbox{for some}~{}a_{2}\in{{{\cal R}_{\kern-1.0pt\lambda}}}\big{\}},\\[2.0pt] {C_{1}}=\rho_{1}(C\cap({{\cal R}_{\kern-1.0pt\lambda}}\times 0))=\big{\{}a_{1}\in{{{\cal R}_{\kern-1.0pt\lambda}}}~{}|~{}(a_{1},0)\in{C}\big{\}},\\[2.0pt] \tilde{C_{2}}=\rho_{2}(C)=\big{\{}a_{2}\in{{{\cal R}_{\kern-1.0pt\lambda}}}~{}|~{}(a_{1},a_{2})\in{C}~{}\mbox{for some}~{}a_{1}\in{{{\cal R}_{\kern-1.0pt\lambda}}}\big{\}},\\[2.0pt] {C_{2}}=\rho_{2}(C\cap(0\times{{\cal R}_{\kern-1.0pt\lambda}}))=\big{\{}a_{2}\in{{{\cal R}_{\kern-1.0pt\lambda}}}~{}|~{}(0,a_{2})\in{C}\big{\}}.\end{array}$ (3.8)
Then $C_{i}\subseteq\tilde{C_{i}}$ are ideals of ${{\cal R}_{\kern-1.0pt\lambda}}$, $i=1,2$. For any $c_{1}+C_{1}\in{\tilde{C_{1}}/C_{1}}$, there is a unique $c_{2}+C_{2}\in{\tilde{C_{2}}/C_{2}}$ such that $(c_{1},c_{2})\in{C}$, hence we have the map
$\varphi:\tilde{C}_{1}/C_{1}\to\tilde{C}_{2}/C_{2},~{}c_{1}+C_{1}\mapsto c_{2}+C_{2},$ (3.9)
which is an ${{\cal R}_{\kern-1.0pt\lambda}}$-isomorphism, and
$C=\big{\{}(c_{1},c_{2})\;\big{|}\;c_{i}\in\tilde{C}_{i}~{}{\rm for}~{}i=1,2;~{}\varphi(c_{1}+C_{1})=c_{2}+C_{2}\big{\}}.$ (3.10)
Since ${{\cal R}_{\kern-1.0pt\lambda}}$ is semisimple and ${C}_{1}\subseteq\tilde{C}_{1}$ are ideals of ${{\cal R}_{\kern-1.0pt\lambda}}$, there is an ideal $C_{12}$ of ${\cal R}_{\kern-1.0pt\lambda}$ such that $\tilde{C}_{1}=C_{1}\oplus C_{12}$; see Remark 2.4(1). Similarly, we have an ideal $C_{12}^{\prime}$ of ${{\cal R}_{\kern-1.0pt\lambda}}$ such that $\tilde{C}_{2}=C_{2}\oplus C_{12}^{\prime}$. Then $\tilde{C}_{1}/C_{1}\cong C_{12}$ and $\tilde{C}_{2}/C_{2}\cong C^{\prime}_{12}$. The ${{\cal R}_{\kern-1.0pt\lambda}}$-isomorphism $\varphi$ in Eq.(3.9) induces an ${{\cal R}_{\kern-1.0pt\lambda}}$-isomorphism $\varphi^{\prime}:C_{12}\to C_{12}^{\prime}$ such that $\varphi\big{(}c+C_{1}\big{)}=\varphi^{\prime}(c)+C_{2}$ for all $c\in C_{12}$. Thus the image $C_{12}^{\prime}=C_{12}$, and there is a $g\in C_{12}^{\times}$ such that $\varphi^{\prime}(c)=cg$ for all $c\in C_{12}$; cf. Remark 2.4(2). In conclusion, we have an ideal $C_{12}$ of ${\cal R}_{\kern-1.0pt\lambda}$ such that $\tilde{C}_{1}=C_{1}\oplus C_{12},~{}\tilde{C}_{2}=C_{2}\oplus C_{12};$ and a $g\in C_{12}^{\times}$ such that
$C=\big{\{}\,(c_{1}+c_{12},\;c_{2}+c_{12}g)\;\big{|}\;c_{1}\in C_{1},\,c_{2}\in C_{2},\,c_{12}\in C_{12}\,\big{\}}.$
Obviously, $(c_{1}\!+\!c_{12},\;c_{2}\!+\!c_{12}g)=(c_{1},c_{2})+(c_{12},c_{12}g)$. Thus Eq.(3.3) holds. ∎
Remark 3.4.
For $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}\in{{\cal R}_{\kern-1.0pt\lambda}}$, by $C_{a}$ we denote the ideal of ${{\cal R}_{\kern-1.0pt\lambda}}$ generated by $a(X)$. Similarly, for $(a(X),a^{\prime}(X))\in{\cal R}_{\kern-1.0pt\lambda}^{2}$, by $C_{a,a^{\prime}}$ we denote the ${{\cal R}_{\kern-1.0pt\lambda}}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ generated by $\big{(}a(X),\,a^{\prime}(X)\big{)}$. Note that in the semisimple case any ideal of ${\cal R}_{\kern-1.0pt\lambda}$ is generated by one element, cf. Remark 2.4(1). However, some ${\cal R}_{\kern-1.0pt\lambda}$-submodules of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ can not be generated by one element. For example, as an ${\cal R}_{\kern-1.0pt\lambda}$-submodule ${\cal R}_{\kern-1.0pt\lambda}^{2}$ can not be generated by one element (because any ${\cal R}_{\kern-1.0pt\lambda}$-module generated by one element is a quotient of the regular module).
Lemma 3.5.
Assume that $\gcd(n,q)=1$. Then for $a,b\in{{\cal R}_{\kern-1.0pt\lambda}}$, $C_{a}\cap C_{b}=0$ if and only if $ab=0$.
Proof.
If $C_{a}\cap C_{b}=0$ then $ab\in C_{a}\cap C_{b}=0$. Conversely, assume that $ab=0$. Note that $C_{b}={\cal R}_{\kern-1.0pt\lambda}\,e_{b}$ for an idempotent $e_{b}$ which is the identity of the ring $C_{b}$ (cf. Remark 2.4(1)). There is an element $f\in{\cal R}_{\kern-1.0pt\lambda}$ such that $e_{b}=fb$. For any element $c\in{C_{a}\cap C_{b}}$, we can write $c=ra=sb$ with $r,\,s\in{{\cal R}_{\kern-1.0pt\lambda}}$. Then $c=ra=rae_{b}=rafb=0$. Thus $C_{a}\cap C_{b}=0$. ∎
Corollary 3.6.
Keep the notation in Theorem 3.3. Any ${\cal R}_{\kern-1.0pt\lambda}$-submodule $C$ of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ can be written as
$C=(C_{a}\times C_{a^{\prime}})\oplus C_{b,bg}=C_{a,0}\oplus C_{0,a^{\prime}}\oplus C_{b,bg},$ (3.11)
where $a,a^{\prime},b\in{\cal R}_{\kern-1.0pt\lambda}$ and $g\in C_{b}^{\times}$ satisfy that $ab=a^{\prime}b=0$.
Proof.
Take $C_{1}=C_{a}$, $C_{2}=C_{a^{\prime}}$ and $C_{12}=C_{b}$, hence $\widehat{C}_{12}=C_{b,bg}$. By the above Lemma, we get Eq.(3.11) immediately. ∎
Definition 3.7.
For any $a(X)=a_{0}+a_{1}X+\cdots+a_{n-1}X^{n-1}\in{F[X]}$ with $\deg(a(X))<n$, we have an $n\times n$ matrix
$a(P_{\!\lambda})=a_{0}E+a_{1}P_{\!\lambda}+\cdots+a_{n-1}P_{\!\lambda}^{n-1}=\begin{pmatrix}a_{0}&a_{1}&\cdots&a_{n-1}\\ \lambda a_{n-1}&a_{0}&\cdots&a_{n-2}\\ \cdots&\cdots&\cdots&\cdots\\ \lambda a_{1}&\lambda a_{2}&\cdots&a_{0}\end{pmatrix},$ (3.12)
whose first row is the vector $\mathbf{a}=(a_{0},a_{1},\dots,a_{n-1})$, and each next row is obtained by $\lambda$-constacyclically permuting the present row (cf. Eq.(2.5)). We call $a(P_{\!\lambda})$ the $\lambda$-consta circulant matrix associated with the polynomial $a(X)$.
Lemma 3.8.
Let $a(X),a^{\prime}(X)\in{\cal R}_{\kern-1.0pt\lambda}$. Then we have:
(1)  $C_{a}$ is linearly generated by the rows of the $n\times n$ matrix $a(P_{\!\lambda})$.
(2)  $C_{a,a^{\prime}}$ is linearly generated by the rows of the $n\times 2n$ matrix $\big{(}a(P_{\!\lambda}),a^{\prime}(P_{\!\lambda})\big{)}$.
Proof.
(1). Let $C_{a}$ be the ideal of ${\cal R}_{\kern-1.0pt\lambda}$ generated by $a(X)$, i.e.,
$C_{a}=\Big{\{}f(X)a(X)~{}\Big{|}~{}f(X)=\sum_{i=0}^{n-1}f_{i}X^{i}\in{\cal R}_{\kern-1.0pt\lambda},f_{i}\in F\Big{\}}.$
Obviously, $f(X)a(X)=f_{0}\cdot a(X)+f_{1}\cdot Xa(X)+\dots+f_{n-1}\cdot X^{n-1}a(X)$. So $C_{a}$ is the subspace of ${\cal R}_{\kern-1.0pt\lambda}$ linearly generated by $a(X),Xa(X),\dots,X^{n-1}a(X)$. Eq.(2.10) and Eq.(3.12) imply that $a(X)$ is identified with the row vector ${\bf a}$ which is just the first row of the matrix $a(P_{\lambda})$; for $i=1,\dots,n-1$, $X^{i}a(X)$ is identified with the row vector ${\bf a}\cdot P_{\lambda}^{\,i}$, which is just the $i$’th row of the matrix $a(P_{\lambda})$. Therefore, $C_{a}$ is linearly generated by the rows of the matrix $a(P_{\lambda})$.
Obviously, (2) is proved in a similar way. ∎
Corollary 3.9.
For $a(X)\in{{\cal R}_{\kern-1.0pt\lambda}}$, the $2$-quasi $\lambda$-constacyclic code $C_{1,a}$ has a generating matrix $\big{(}E_{n},\,a(P_{\!\lambda})\big{)}$.
Proof.
By Lemma 3.8, the rows of the matrix $\big{(}E_{n},\,a(P_{\!\lambda})\big{)}$ linearly generate the code $C_{1,a}$. And the rows of the matrix $\big{(}E_{n},\,a(P_{\!\lambda})\big{)}$ are linearly independent. ∎
For any matrix $A=\big{(}a_{ij}\big{)}_{m\times n}$ with $a_{ij}\in F$, we denote $A^{(p^{h})}=\big{(}a_{ij}^{p^{h}}\big{)}_{m\times n}$ (cf. Example 2.1(2)). By $A^{T}$ we denote the transpose of $A$. Let
$A^{*h}=(A^{(p^{h})})^{T}=\big{(}a_{ji}^{p^{h}}\big{)}_{n\times m}=\begin{pmatrix}a_{11}^{p^{h}}&\cdots&a_{m1}^{p^{h}}\\ \dots&\dots&\dots\\ a_{1n}^{p^{h}}&\cdots&a_{mn}^{p^{h}}\end{pmatrix}$ (3.13)
be the transpose of the matrix $A^{(p^{h})}$. We call $A^{*h}$ the Galois $p^{h}$-transpose of $A$. If $h=0$, $A^{*0}=A^{T}$ is just the transpose matrix of $A$. If $\ell$ is even and $h=\frac{\ell}{2}$, $A^{*\frac{\ell}{2}}$ is the Hermitian transpose of $A$.
For $B=(b_{ij})_{m\times n}$, $C=(c_{ij})_{n\times k}$ and $\alpha\in F$, it is easy to check that
$(AC)^{*h}=C^{*h}A^{*h},~{}~{}(A+B)^{*h}=A^{*h}+B^{*h},~{}~{}(\alpha A)^{*h}=\alpha^{p^{h}}A^{*h}.$ (3.14)
[…]
With the identification Eq.(2.9) and the operator “$*h$”, we can compute the Galois $p^{h}$-inner product on ${{\cal R}_{\kern-1.0pt\lambda}}$ (see Eq.(1.5)) in a matrix version:
$\big{\langle}a,\,b\big{\rangle}_{h}=\mathbf{a}\cdot\mathbf{b}^{*h},\qquad\forall~{}a,b\in{{\cal R}_{\kern-1.0pt\lambda}}.$ (3.15)
And for any $\lambda$-constacyclic code $C$ (i.e., any idea of ${\cal R}_{\kern-1.0pt\lambda}$), the Galois $p^{h}$-dual code of $C$ is as follows:
$C^{\bot h}=\big{\{}\,a\in{{{\cal R}_{\kern-1.0pt\lambda}}}~{}\big{|}~{}\big{\langle}c,\,a\big{\rangle}_{h}=0,~{}\forall\,c\in{C}\,\big{\}}.$
Similarly, with the identification Eq.(2.11) the Galois $p^{h}$-inner product on ${{\cal R}_{\kern-1.0pt\lambda}}^{2}$ is computed in a matrix version:
$\big{\langle}(a,a^{\prime}),\,(b,b^{\prime})\big{\rangle}_{h}=\big{(}\mathbf{a},\mathbf{a}^{\prime}\big{)}\cdot\big{(}\mathbf{b},\mathbf{b}^{\prime}\big{)}^{*h},\qquad\forall~{}\big{(}a,a^{\prime}\big{)},\,\big{(}b,b^{\prime}\big{)}\in{\cal R}_{\kern-1.0pt\lambda}^{2}.$ (3.16)
And for any $2$-quasi $\lambda$-constacyclic code $C$ (${\cal R}_{\kern-1.0pt\lambda}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$), the Galois $p^{h}$-dual code of $C$ is
$C^{\bot h}=\big{\{}\,(a,a^{\prime})\in{{\cal R}_{\kern-1.0pt\lambda}^{2}}~{}\big{|}~{}\big{\langle}(c,c^{\prime}),\,(a,a^{\prime})\big{\rangle}_{h}=0,~{}\forall\,(c,c^{\prime})\in{C}\,\big{\}}.$
If $C=C^{\bot h}$, then we say that $C$ is Galois $p^{h}$-self-dual, or Galois self-dual.
Lemma 3.10.
If $C\subseteq F^{n}\times F^{n}$ is a $2$-quasi $\lambda$-constacyclic code, then $C^{\bot h}$ is a $2$-quasi $\lambda^{\!-\!p^{\ell-h}}\!$-constacyclic code. In particular, if $h=0$, $C^{\bot 0}=C^{\bot}$ is a $2$-quasi $\lambda^{\!-\!1}$-constacyclic code.
Proof.
Assume that $\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}\in C^{\bot h}$. By Lemma 2.2(2), it is enough to prove that for any $\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)}\in C$ we have
$\left\langle\big{(}\mathbf{c},\,\mathbf{c}\big{)},\,\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}\begin{pmatrix}P_{\lambda^{\!-\!p^{\ell-h}}}\!\\ &\!P_{\lambda^{\!-\!p^{\ell-h}}}\end{pmatrix}\right\rangle_{\!h}=0.$
Applying Eq.(3.14) and Eq.(3.16) yields
$\left\langle\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)},\,\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}\begin{pmatrix}P_{{\lambda^{\!-\!p^{\ell-h}}}}\!\\ &\!P_{{\lambda^{\!-\!p^{\ell-h}}}}\end{pmatrix}\right\rangle_{\!h}$
$=\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)}\cdot\left(\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}\begin{pmatrix}P_{{\lambda^{\!-\!p^{\ell-h}}}}\!\\ &\!P_{{\lambda^{\!-\!p^{\ell-h}}}}\end{pmatrix}\right)^{*h}$
$=\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)}\begin{pmatrix}P_{\lambda^{\!-\!p^{\ell-h}}}\!\\ &\!P_{\lambda^{\!-\!p^{\ell-h}}}\end{pmatrix}^{*h}\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}^{*h}$
$=\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)}\begin{pmatrix}\big{(}P_{\lambda^{\!-\!p^{\ell-h}}}\big{)}^{*h}\!\\ &\!\big{(}P_{\lambda^{\!-\!p^{\ell-h}}}\big{)}^{*h}\end{pmatrix}\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}^{*h}.$
Because $(\lambda^{\!-\!p^{\ell-h}})^{p^{h}}=(\lambda^{p^{\ell}})^{-1}=\lambda^{-1}$, it is easy to check that
$P_{\!\lambda}\big{(}P_{\lambda^{\!-\!p^{\ell-h}}}\big{)}^{*h}=\begin{pmatrix}&E_{n-1}\\ \lambda E_{1}\end{pmatrix}\begin{pmatrix}&\!(\lambda^{\!-\!p^{\ell-h}})^{p^{h}}E_{1}\\ E_{n-1}\!\end{pmatrix}=E_{n}.$
Since $P_{\!\lambda}^{n}=\lambda E_{n}$ and ${\rm ord}_{F^{\times}}(\lambda)=t$ (see Eq(3.1)), we have $P_{\!\lambda}^{nt}=E_{n}$, hence $P_{\!\lambda}^{-1}=P_{\!\lambda}^{nt-1}$. So
$\big{(}P_{\lambda^{\!-\!p^{\ell-h}}}\big{)}^{*h}=P_{\!\lambda}^{-1}=P_{\!\lambda}^{nt-1}.$ (3.17)
By Lemma 2.2(2), $\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)}\begin{pmatrix}P_{\!\lambda}\!\\ &\!P_{\!\lambda}\end{pmatrix}^{nt-1}\in C$. We get that
$\bigg{\langle}\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)},\,\big{(}\mathbf{a},\mathbf{a}^{\prime}\big{)}\begin{pmatrix}P_{\!\lambda^{\!-\!p^{\ell-h}}}\!\!\\ &\!\!P_{\!\lambda^{\!-\!p^{\ell-h}}}\end{pmatrix}\bigg{\rangle}_{\!h}=\bigg{\langle}\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)}\begin{pmatrix}P_{\!\lambda}\!\\ &\!P_{\!\lambda}\end{pmatrix}^{nt-1}\!,\,\big{(}\mathbf{a},\mathbf{a}^{\prime}\big{)}\bigg{\rangle}_{\!h}=0.$
We are done. ∎
Remark 3.11.
We note that $\lambda=\lambda^{-p^{\ell-h}}$ if and only if $\lambda^{1+p^{h}}=1$, since
$\lambda=\lambda^{-p^{\ell-h}}\;\iff\;\lambda^{p^{h}}=(\lambda^{-p^{\ell-h}})^{p^{h}}=\lambda^{-p^{\ell}}=\lambda^{-1}\;\iff\;\lambda^{1+p^{h}}=1.$
If $\lambda=\lambda^{-p^{\ell-h}}$, then for any $2$-quasi $\lambda$-constacyclic code $C$, its Galois $p^{h}$-dual code $C^{\bot h}$ is still a $2$-quasi $\lambda$-constacyclic code. Otherwise (i.e., $\lambda\neq\lambda^{-p^{\ell-h}})$, Lemma 3.2 implies that, for many $2$-quasi $\lambda$-constacyclic codes, their Galois $p^{h}$-dual codes are no longer $2$-quasi $\lambda$-constacyclic codes.
4 Galois self-dual $2$-quasi constacyclic codes
In this section we investigate the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$. By Lemma 3.10 and Remark 3.11, any Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic code (i.e., $C=C^{\bot h}$) is also $\lambda^{-p^{\ell-h}}$-constacyclic. We study them in two cases: $\lambda\neq\lambda^{-p^{\ell-h}}$ (i.e., $\lambda^{1+p^{h}}\neq 1$), or $\lambda=\lambda^{-p^{\ell-h}}$ (i.e., $\lambda^{1+p^{h}}=1$).
4.1 The case that $\lambda^{1+p^{h}}\neq 1$
Our concern in this subsection is the Galois $p^{h}$-self-dual 2-quasi $\lambda-$constacyclic codes over $F$ under the assumption that $\lambda^{1+p^{h}}\neq 1$.
Theorem 4.1.
Assume that $\lambda^{1+p^{h}}\neq 1$ (i.e., $\lambda\neq\lambda^{-p^{\ell-h}}$). The following three are equivalent to each other:
(1) $C$ is a Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic code over $F$ of length $2n$.
(2) $C=C_{1,\alpha}$ is a ${{\cal R}_{\kern-1.0pt\lambda}}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ generated by $(1,\alpha)$, where $\alpha\in F^{\times}$ with $\alpha^{1+p^{h}}=-1$, and $1,\alpha$ are viewed as the constant polynomials of ${{\cal R}_{\kern-1.0pt\lambda}}$.
(3) $C$ is an $F$-linear code of length $2n$ with a generating matrix $\big{(}E_{n},\alpha E_{n}\big{)}$, where $\alpha\in F^{\times}$ with $\alpha^{1+p^{h}}=-1$.
Proof.
Observe that by Corollary 3.9, the statements (2) and (3) are equivalent. Suppose that (3) holds. Then $\big{(}E_{n},\,\alpha E_{n}\big{)}\big{(}E_{n},\,\alpha E_{n}\big{)}^{*h}=(1+\alpha^{p^{h}+1})E_{n}=0,$ which implies $\langle C,C\rangle_{h}=0$, cf. Eq.(3.16). Since the rank of the generating matrix $\big{(}E_{n},\,\alpha E_{n}\big{)}$ is $n$, we have $\dim_{F}C=n$, and so the statement (1) follows. Therefore, it suffices to show that (1) implies (3).
Assume that (1) holds, i.e., $C=C^{\bot h}$. By Lemma 3.10, $C$ is both a $2$-quasi $\lambda$-constacyclic code and a $2$-quasi $\lambda^{\!-\!p^{\ell-h}}$-constacyclic code. Since $\lambda\neq\lambda^{-p^{\ell-h}}$, we deduce from Lemma 3.2 that either $\rho_{1}(C)=0$ or $\rho_{1}(C)={{\cal R}_{\kern-1.0pt\lambda}}$. Suppose that $\rho_{1}(C)=0$. Since $\dim_{F}C=n$, by Eq.(3.2), we have that $C=0\times{{\cal R}_{\kern-1.0pt\lambda}}$, which is impossible because $0\times{{\cal R}_{\kern-1.0pt\lambda}}$ is not Galois $p^{h}$-self-dual. So, it must be the case that $\rho_{1}(C)={{\cal R}_{\kern-1.0pt\lambda}}$. By the same argument, we have $\rho_{2}(C)={{\cal R}_{\kern-1.0pt\lambda}}$.
As $\rho_{1}(C)={{\cal R}_{\kern-1.0pt\lambda}}$, we can take $(1,a)\in C$ with $a=a(X){=\sum_{i=0}^{n-1}a_{i}X^{i}\in{{{\cal R}_{\kern-1.0pt\lambda}}}}$. Let $C_{1,a}$ be the ${{\cal R}_{\kern-1.0pt\lambda}}-$submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ generated by $(1,a)$. Obviously, $C_{1,a}\subseteq C$. By Corollary 3.9, $C_{1,a}$ has the generating matrix
$\big{(}E_{n},\;a(P_{\lambda})\big{)}=\begin{pmatrix}1&&&&&a_{0}&a_{1}&\cdots&a_{n-1}\\ &1&&&&\lambda a_{n-1}&a_{0}&\cdots&a_{n-2}\\ &&\ddots&&&\cdots&\cdots&\cdots&\cdots\\ &&&1&&\lambda a_{1}&\lambda a_{2}&\cdots&a_{0}\end{pmatrix}_{n\times 2n}.$ (4.1)
The rank of the matrix is $n$, hence $\dim_{F}C_{1,a}=n$. In a word, $C=C_{1,a}$ is linearly generated by the $n$ rows of the matrix in Eq.(4.1).
Because $C$ is also a $2$-quasi $\lambda^{\!-\!p^{\ell-h}}\!$-constacyclic code, by the same way we can also get that $C$ is linearly generated by the $n$ rows of the matrix $\big{(}E_{n},a(P_{\!\lambda^{-p^{\ell-h}}})\big{)}$. Thus any row of $\big{(}E_{n},a(P_{\!\lambda^{-p^{\ell-h}}})\big{)}$ is a linear combination of the rows of the matrix in Eq.(4.1). So there is an $n\times n$ matrix $A$ such that
$\big{(}E_{n},\;a(P_{\!\lambda^{-p^{\ell-h}}})\big{)}=A\Big{(}E_{n},\;a(P_{\lambda})\Big{)}=\Big{(}AE_{n},\;A\cdot a(P_{\lambda})\Big{)}.$
It follows that $A=E_{n}$, and so $a(P_{\!\lambda^{\!-\!p^{\ell-h}}})=A\cdot a(P_{\!\lambda})=a(P_{\!\lambda})$. The latter equality is as follows:
$\begin{pmatrix}a_{0}&a_{1}&\cdots&a_{n-1}\\ \lambda^{\!-\!p^{\ell-h}}a_{n-1}&a_{0}&\cdots&a_{n-2}\\ \cdots&\cdots&\cdots&\cdots\\ \lambda^{\!-\!p^{\ell-h}}a_{1}&\lambda^{\!-\!p^{\ell-h}}a_{2}&\cdots&a_{0}\end{pmatrix}=\begin{pmatrix}a_{0}&a_{1}&\cdots&a_{n-1}\\ \lambda a_{n-1}&a_{0}&\cdots&a_{n-2}\\ \cdots&\cdots&\cdots&\cdots\\ \lambda a_{1}&\lambda a_{2}&\cdots&a_{0}\end{pmatrix}.$
Since $\lambda^{-p^{\ell-h}}\neq\lambda$, it follows that $a_{i}=0$ for $i=1,\dots,n-1$; hence the polynomial $a(X)$ is a constant polynomial: $a(X)=\alpha$ for some $\alpha\in{F}$, and $C$ has a generating matrix $\big{(}E_{n},\,\alpha E_{n}\big{)}$. Because $C$ is Galois $p^{h}$-self-dual,
$0=\big{(}E_{n},\,\alpha E_{n}\big{)}\cdot\big{(}E_{n},\,\alpha E_{n}\big{)}^{*h}=E_{n}+\alpha^{1+\alpha^{p^{h}}}E_{n}=(1+\alpha^{1+p^{h}})E_{n},$
hence $01+\alpha^{1+p^{h}}=$ 0. In conclusion, (3) holds. ∎
Corollary 4.2.
Assume that $\lambda^{1+p^{h}}\neq 1$. The Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes of length $2n$ exist if and only if the polynomial $X^{1+p^{h}}+1$ has roots in $F$; and in that case, the ${{\cal R}_{\kern-1.0pt\lambda}}$-submodules $C_{1,\alpha}$ of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ with $\alpha\in F$ being a root of $X^{1+p^{h}}+1$ are all the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes of length $2n$. ∎
Note that in the Euclidean case “$h=0$”, $\lambda^{1+p^{h}}\neq 1$ if and only if $\lambda\neq\pm 1$.
Corollary 4.3.
[…]
Proof.
[…]
As a comparison, the Hermitian self-dual ones always exist.
Corollary 4.4.
Assume that $\ell$ is even and $\lambda^{1+p^{\ell/2}}\neq 1$. Then the Hermitian self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ of length $2n$ always exist; and in that case, the ${{\cal R}_{\kern-1.0pt\lambda}}$-submodules $C_{1,\alpha}$ of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ with $\alpha\in F$ satisfying $\alpha^{1+p^{\ell/2}}=-1$ are the all Hermitian self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ of length $2n$.
Proof.
If $q=p^{\ell}$ is even, then the polynomial $X^{1+p^{\ell/2}}+1=X^{1+p^{\ell/2}}-1$ always has roots in $F$. Assume that $p$ is odd. The order of the multiplication group
$|F^{\times}|=p^{\ell}-1=(p^{\ell/2}+1)(p^{\ell/2}-1),$
and $2\,\big{|}\,(p^{\ell/2}-1)$. So, there is a subgroup $H$ of the multiplication group $F^{\times}$ with order $|H|=2(p^{\ell/2}+1)$. Then any generator of the group $H$ is a root of the polynomial $X^{1+p^{\ell/2}}+1$. ∎
As a consequence, we get the following.
Theorem 4.5.
Assume that $\lambda^{1+p^{h}}\neq 1$ (i.e., $\lambda\neq\lambda^{-p^{\ell-h}}$). Then the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ are asymptotically bad.
Proof.
If the polynomial $X^{1+p^{h}}+1$ has no root in $F$, then Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ do not exist, hence Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ are asymptotically bad.
Otherwise, any Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic code $C_{1,\alpha}$ has minimum weight ${\rm w}(C_{1,\alpha})=2$, because any row of the generating matrix $\big{(}E_{n},\alpha E_{n}\big{)}$ has weight $2$. The relative minimum distance $\Delta(C_{1,\alpha})=\frac{2}{2n}\to 0$ while ${n\to\infty}$. So Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ are asymptotically bad. ∎
4.2 The case that $\lambda^{1+p^{h}}=1$
We start with a general result about the $\lambda$-consta circulant matrices. As in Example 2.1(2), ${\rm M}_{n}(F)$ denotes the $n\times n$ matrix algebra over $F$. We consider its subset consisting of all the $\lambda$-consta circulant matrices of degree $n$:
${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)=\big{\{}\,a(P_{\lambda})\;\big{|}\;a(X)\in{\cal R}_{\kern-1.0pt\lambda}\,\big{\}},$
where $P_{\lambda}$ is defined in Eq.(2.4) and $a(P_{\lambda})$ is defined in Eq.(3.12).
Lemma 4.6.
With the notation as above. ${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$ is a subalgebra of ${\rm M}_{n}(F)$ and the following is an algebra isomorphism:
$\nu:\;{{\cal R}_{\kern-1.0pt\lambda}}\;\mathop{\longrightarrow}\;{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F),\quad a(X)\,\longmapsto\,a(P_{\!\lambda}).$ (4.2)
Proof.
The following is obviously an $F$-algebra homomorphism:
$\delta:~{}~{}F[X]\to{\rm M}_{n}(F),~{}~{}~{}f(X)\mapsto f(P_{\!\lambda}).$
Since $P_{\!\lambda}^{n}=\lambda E_{n}$, the kernel of the homomorphism $\delta$ is the ideal $\langle X^{n}-\lambda\rangle$ of $F[X]$ generated by $X^{n}-\lambda$. Note that the quotient algebra $F[X]/\langle X^{n}-\lambda\rangle={{\cal R}_{\kern-1.0pt\lambda}}$, see Eq.(2.7). By Homomorphism Theorem, the homomorphism $\delta$ induces an injective homomorphism:
$\hat{\delta}:~{}~{}{\cal R}_{\kern-1.0pt\lambda}=F[X]/\langle X^{n}-\lambda\rangle~{}\to~{}{\rm M}_{n}(F),~{}~{}~{}a(X)~{}\mapsto~{}a(P_{\!\lambda}).$
The image of this homomorphism $\hat{\delta}$ is exactly ${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$, cf. Definition 3.7. Thus ${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$ is a subalgebra of ${\rm M}_{n}(F)$ and the homomorphism $\hat{\delta}$ induces the algebra isomorphism $\nu$ in Eq.(4.2). ∎
In the rest of this subsection we assume that $\lambda^{1+p^{h}}=1$ i.e., $\lambda=\lambda^{-p^{\ell-h}}$, cf. Remark 3.11. By Eq.(3.13), we have the following map
$\tilde{\tau}:~{}{\rm M}_{n}(F)\,\longrightarrow\,{\rm M}_{n}(F),~{}~{}~{}A\,\longmapsto\,A^{*h};$
and by Eq.(3.14), for any $A,B\in{\rm M}_{n}(F)$ and $\alpha\in F$,
$(AB)^{*h}=B^{*h}A^{*h},~{}~{}(A+B)^{*h}=A^{*h}+B^{*h},~{}~{}(\alpha A)^{*h}=\alpha^{p^{h}}\!A^{*h}.$
So $\tilde{\tau}$ is a $p^{h}$-anti-automorphism of the matrix algebra ${\rm M}_{n}(F)$.
Lemma 4.7.
Keep the notation as above. Assume that $\lambda^{1+p^{h}}=1$. Then $\tilde{\tau}\big{(}{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)\big{)}\subseteq{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$, and the restricted map $\tau=\tilde{\tau}|_{{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)}$ as follows is a $p^{h}$-automorphism of the $F$-algebra ${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$:
$\tau:\;{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)\;\mathop{\longrightarrow}\;{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F),\quad a(P_{\lambda})\,\longmapsto\,\big{(}a(P_{\lambda})\big{)}^{*h}.$ (4.3)
Proof.
Since $\lambda=\lambda^{-p^{\ell-h}}$, by Eq.(3.17) we deduce that
$P_{\lambda}^{\,*h}=P_{\!\lambda}^{-1}=P_{\!\lambda}^{nt-1}\in{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F).$ (4.4)
By Eq.(4.2), any $a(P_{\lambda})\in{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$ is associated with $a(X)\in{{\cal R}_{\kern-1.0pt\lambda}}$, where $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}\in{\cal R}_{\kern-1.0pt\lambda}$, so
$\big{(}a(P_{\lambda})\big{)}^{*h}=\Big{(}\sum_{i=0}^{n-1}a_{i}P_{\lambda}^{\,i}\Big{)}^{*h}=\sum_{i=0}^{n-1}a_{i}^{p^{h}}(P_{\lambda}^{\,*h})^{i}\in{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F).$
Thus $\tilde{\tau}\big{(}{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)\big{)}\subseteq{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$. Restricting the $p^{h}$-anti-automorphism $\tilde{\tau}$ to ${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$, we get the $p^{h}$-anti-automorphism Eq.(4.3), which is in fact a $p^{h}$-automorphism because ${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)\cong{\cal R}_{\kern-1.0pt\lambda}$ is a commutative algebra. ∎
Next, we introduce an operator “*” on ${\cal R}_{\kern-1.0pt\lambda}$, which is the key to obtaining the necessary and sufficient conditions for $2$-quasi $\lambda$-constacyclic codes being Galois $p^{h}$-self-dual. With the isomorphism $\nu$ in Eq.(4.2), inspiring by Eq.(4.3) and Eq.(4.4), for $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}\in{\cal R}_{\kern-1.0pt\lambda}$ we define
$a^{*}(X)=a^{(p^{h})}(X^{nt-1})~{}({\rm mod}~{}X^{n}-\lambda),$ (4.5)
where $a^{(p^{h})}(X)=\sum_{i=0}^{n-1}a_{i}^{p^{h}}X^{i}$, cf. Example 2.1(1).
Lemma 4.8.
Assume that $\lambda^{1+p^{h}}=1$. Let $a^{*}(X)$, $\nu$ and $\tau$ be as in Eq.(4.5), Eq.(4.2) and Eq.(4.3) respectively. Then
$a^{*}(X)=\nu^{-1}\tau\nu\big{(}a(X)\big{)},~{}~{}~{}~{}\forall\;a(X)\in{\cal R}_{\kern-1.0pt\lambda};$ (4.6)
and the following map is a $p^{h}$-automorphism of the algebra ${\cal R}_{\kern-1.0pt\lambda}$:
$*:~{}{\cal R}_{\kern-1.0pt\lambda}\;\mathop{\longrightarrow}\;{\cal R}_{\kern-1.0pt\lambda},~{}~{}~{}a(X)\,\longmapsto\,a^{*}(X)\,.$ (4.7)
Proof.
For $a(X)\in{\cal R}_{\kern-1.0pt\lambda}$, by the definition of $a^{*}(X)$ in Eq.(4.5) and by Eq.(4.4),
$a^{*}(P_{\!\lambda})$ $=\sum_{i=0}^{n-1}a_{i}^{p^{h}}(P_{\!\lambda}^{nt-1})^{i}=\sum_{i=0}^{n-1}a_{i}^{p^{h}}(P_{\!\lambda}^{*h})^{i}=\sum_{i=0}^{n-1}a_{i}^{p^{h}}(P_{\!\lambda}^{i})^{*h}$
$=\sum_{i=0}^{n-1}(a_{i}P_{\!\lambda}^{i})^{*h}=\Big{(}\sum_{i=0}^{n-1}a_{i}P_{\!\lambda}^{i}\Big{)}^{*h}=\big{(}a(P_{\!\lambda})\big{)}^{*h}.$
That is, $\nu\big{(}a^{*}(X)\big{)}=a^{*}(P_{\!\lambda})=\big{(}a(P_{\!\lambda})\big{)}^{*h}=\tau\nu\big{(}a(X)\big{)}$. So Eq.(4.6) holds; equivalently, the following diagram is commutative:
$\begin{array}[]{ccc}{\cal R}_{\kern-1.0pt\lambda}&\mathop{\longrightarrow}\limits^{\nu}&{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)\\[5.0pt] *\,\big{\downarrow}{}&&\big{\downarrow}\,\tau\\[5.0pt] {\cal R}_{\kern-1.0pt\lambda}&\mathop{\longrightarrow}\limits_{\nu}&{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)\end{array}$ (4.11)
Because $\tau$ is a $p^{h}$-automorphism and both $\nu$ and $\nu^{-1}$ are algebra isomorphism, by Eq.(4.6) we see that Eq.(4.7) is a $p^{h}$-automorphism of ${\cal R}_{\kern-1.0pt\lambda}$. ∎
Recall that the $2$-quasi $\lambda$-constacyclic code $C_{a,a^{\prime}}$ generated by $(a,a^{\prime})\in{\cal R}_{\kern-1.0pt\lambda}^{2}$ has been defined in Remark 3.4. For the ${\cal R}_{\kern-1.0pt\lambda}$-submodules of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ generated by one element, we have the following Galois self-duality criteria.
Lemma 4.9.
Assume that $\lambda^{1+p^{h}}=1$.
(1) The $2$-quasi $\lambda$-constacyclic code $C_{a,a^{\prime}}$ is Galois $p^{h}$-self-dual if and only if $aa^{*}+a^{\prime}a^{\prime*}\!=\!0$ and the rate ${\rm R}(C_{a,a^{\prime}})=1/2$.
(2) The $2$-quasi $\lambda$-constacyclic code $C_{1,g}$ is Galois $p^{h}$-self-dual if and only if $gg^{*}=-1$.
Proof.
(1) The $C_{a,a^{\prime}}$ is Galois $p^{h}$-self-dual if and only if $\langle C_{a,a^{\prime}},C_{a,a^{\prime}}\rangle_{h}=0$ and $\dim_{F}C_{a,a^{\prime}}=n$. What remains is to show that
$\langle C_{a,a^{\prime}},C_{a,a^{\prime}}\rangle_{h}=0\;\iff\;aa^{*}+a^{\prime}{a^{\prime}}^{*}=0.$ (4.12)
Observe that by Lemma 3.8, $C_{a,a^{\prime}}$ is linearly generated by the rows of the matrix $\big{(}a(P_{\!\lambda}),\,a^{\prime}(P_{\!\lambda})\big{)}$. Thus $\langle C_{a,a^{\prime}},C_{a,a^{\prime}}\rangle_{h}=0$ if and only if
$\big{(}a(P_{\!\lambda}),\,a^{\prime}(P_{\!\lambda})\big{)}\cdot\big{(}a(P_{\!\lambda}),\,a^{\prime}(P_{\!\lambda})\big{)}^{*h}=a(P_{\!\lambda})\cdot a(P_{\!\lambda})^{*h}+a^{\prime}(P_{\!\lambda})\cdot a^{\prime}(P_{\!\lambda})^{*h}=0.$
Thus, Eq.(4.12) follows from Lemma 4.8 (cf. Eq.(4.11)) immediately.
(2)  By Corollary 3.9, $C_{1,g}$ has the generating matrix $\big{(}E_{n},\,g(P_{\!\lambda})\big{)}$. In particular, $\dim_{F}C_{1,g}=n$, i.e., ${\rm R}(C_{1,g})=1/2$. Similarly to the proof of $(1)$, we have $\langle C_{1,g},C_{1,g}\rangle_{h}=0$ if and only if $01+gg^{*}=$ 0. ∎
In the semisimple case, extending [14, Theorem 4.2], we have the following theorem which characterize the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes. By Corollary 3.6, any $2$-quasi $\lambda$-constacyclic code $C$ can be written as
$C=C_{a,0}\oplus C_{0,a^{\prime}}\oplus C_{b,bg},$ (4.13)
where $a,a^{\prime},b\in{{\cal R}_{\kern-1.0pt\lambda}}$ with $ab=a^{\prime}b=0$ and $g\in{C_{b}^{\times}}$.
Theorem 4.10.
Assume that $\lambda^{1+p^{h}}=1$ and $\gcd(n,q)=1$. Let $C$ in Eq.(4.13) be any $2$-quasi $\lambda$-constacyclic code. Then $C$ is Galois $p^{h}$-self-dual if and only if the following two hold:
(1)  $aa^{*}=a^{\prime}{a^{\prime}}^{*}=ab^{*}=a^{*}b=a^{\prime}b^{*}=a^{\prime*}b=bb^{*}(1+gg^{*})=0;$
(2)  $\dim_{F}C=n$.
Proof.
The $C$ is Galois $p^{h}$-self-dual if and only if $\langle C,C\rangle_{h}=0$ and $\dim_{F}C=n$. Note that the inner product $\langle-,-\rangle_{h}$ is linear for the first variable, and it is $p^{h}$-linear for the second variable, but it is not symmetric in general. So $\langle C,C\rangle_{h}=0$ is equivalent to the following
$\langle C_{a,0},C_{0,a^{\prime}}\rangle_{h}=\langle C_{0,a^{\prime}},C_{a,0}\rangle_{h}=0,$
$\langle C_{a,0},C_{a,0}\rangle_{h}=\langle C_{0,a^{\prime}},C_{0,a^{\prime}}\rangle_{h}=0,$
$\langle C_{a,0},C_{b,bg}\rangle_{h}=\langle C_{b,bg},C_{a,0}\rangle_{h}=0,$
$\langle C_{0,a^{\prime}},C_{b,bg}\rangle_{h}=\langle C_{b,bg},C_{0,a^{\prime}}\rangle_{h}=0,$
$\langle C_{b,bg},C_{b,bg}\rangle_{h}=0.$
The first line holds obviously. By Eq.(4.12), the second and the third lines are equivalent to that $aa^{*}=a^{\prime}a^{\prime*}=0$ and $ab^{*}=a^{*}b=0$, respectively. The last line is equivalent to that $bb^{*}(1+gg^{*})=0$.
Turn to the forth line which is equivalent to $a^{\prime}b^{*}g^{*}=a^{\prime*}bg=0$. Note that $C_{b}={\cal R}_{\kern-1.0pt\lambda}\,e_{b}$ is a ring with identity $e_{b}$ which is an idempotent, cf. Remark 2.4(1), hence “$g\in C_{b}^{\times}$” implies that $gg^{\prime}=e_{b}$ for a $g^{\prime}\in C_{b}$. So $a^{\prime*}b=a^{\prime*}be_{b}=a^{\prime*}bgg^{\prime}=0g^{\prime}=0$. Similarly, $a^{\prime}b^{*}=a^{\prime}b^{*}e_{b}^{*}=a^{\prime}b^{*}g^{*}g^{\prime*}=0g^{\prime*}=0$.
The theorem is proved. ∎
5 Hermitian self-dual $2$-quasi-cyclic codes
In this section, we always assume that $\ell$ is even and $h=\ell/2$, and ${\gcd(n,q)=1}$. The map $\sigma_{\ell/2}:F\to F$, $\alpha\mapsto\alpha^{p^{\ell/2}}$, is a Galois automorphism of order $2$, and
$\langle{\bf a},{\bf a^{\prime}}\rangle_{\ell/2}=\sum_{i=0}^{n-1}a_{i}{a^{\prime}_{i}}^{p^{\ell/2}},~{}~{}~{}~{}\forall\,{\bf a}=(a_{0},\cdots,a_{n-1}),\,{\bf a^{\prime}}=(a^{\prime}_{0},\cdots,a^{\prime}_{n-1})\in F^{n},$
is the Hermitian inner product on $F^{n}$. In this section we consider ${\cal R}={\cal R}_{1}=F[X]/\langle X^{n}-1\rangle$ and ${\cal R}^{2}={\cal R}\times{\cal R}$, cf. Eq.(2.13); and prove that the Hermitian self-dual $2$-quasi-cyclic codes are asymptotically good.
5.1 The operator “*” on ${\cal R}=F[X]/\langle X^{n}-1\rangle$
Since $11^{1+p^{\ell/2}}=$ 1, the results in Subsection 4.2 can be quoted freely for ${\cal R}$ and ${\cal R}^{2}$. In particular, the operator “*” in Lemma 4.8 is a $p^{\ell/2}$-automorphism of ${\cal R}$:
$*:~{}~{}{\cal R}\;\longrightarrow\;{\cal R},~{}~{}~{}a(X)\;\longmapsto\;a^{*}(X),$ (5.1)
where $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}$ and $a^{*}(X)=a^{(p^{\ell/2})}(X^{n-1})~{}({\rm mod}~{}X^{n}-1)$, i.e., $a^{*}(X)=\sum_{i=0}^{n-1}a_{i}^{p^{\ell/2}}(X^{n-1})^{i}~{}({\rm mod}~{}X^{n}-1)$, cf. Eq.(4.5). By Eq.(2.14), we have the identification:
${\cal R}\cong FG,~{}~{}~{}~{}\sum_{i=0}^{n-1}a_{i}X^{i}~{}\mapsto~{}\sum_{i=0}^{n-1}a_{i}x^{i},$
where $G=\langle\,x\,|\,x^{n}=1\rangle$ is the cyclic group of order $n$ and $FG$ is the cyclic group algebra. So the symbol “$X$” can be identified with the element $x$ of the group $G$, and we can write $a(x)=\sum_{i=0}^{n-1}a_{i}x^{i}\in{\cal R}$, and the expressions $x^{-1}$, $a(x^{-1})$ etc. make sense. Hence $x^{n-1}=x^{-1}$ and
$a^{*}(x)=\sum_{i=0}^{n-1}a_{i}^{p^{\ell/2}}(x^{-1})^{i}={a}^{(p^{\ell/2})}(x^{-1}),~{}~{}~{}~{}\forall\;a(x)\in{\cal R}.$ (5.2)
Lemma 5.1.
The $p^{\ell/2}$-automorphism “*” of ${\cal R}$ in Eq.(5.1) is of order 2.
Proof.
Since $x^{n-1}=x^{-1}$, for any $a(x)\in{\cal R}$ by Eq.(5.2) we have
$\Big{(}\sum_{i=0}^{n-1}a_{i}x^{i}\Big{)}^{**}=\Big{(}\sum_{i=0}^{n-1}a_{i}^{p^{\ell/2}}(x^{-1})^{i}\Big{)}^{*}=\sum_{i=0}^{n-1}(a_{i}^{p^{\ell/2}})^{p^{\ell/2}}((x^{-1})^{-1})^{i}=\sum_{i=0}^{n-1}a_{i}x^{i}.$
Thus the order of the operator “*” equals $1$ or $2$. There is an $\alpha\in F$ such that $\alpha^{p^{\ell/2}}\!\neq\alpha$. Then in ${\cal R}$ we have $(\alpha 1)^{*}\neq\alpha 1$. So the order of the operator “*” equals $2$. ∎
Corollary 5.2.
If $C$ is a non-zero ideal of ${\cal R}$ which is invariant by the operator “*” (i.e., $C^{*}=C$), then the restriction of the operator “*” to $C$ induces a $p^{\ell/2}$-automorphism of $C$ of order $2$.
Proof.
By Lemma 5.1, the restriction of “*” to $C$ is of order $1$ or $2$. Let $c\in C$ with $c\neq 0$. If $c^{*}\neq c$, then the restriction of “*” to $C$ is not the identity. Otherwise, $c^{*}=c$; there is an $\alpha\in F$ with $\alpha^{p^{\ell/2}}\!\neq\alpha$; so $(\alpha c)^{*}=\alpha^{p^{\ell/2}}\!c^{*}=\alpha^{p^{\ell/2}}c\neq\alpha c$. In conclusion, the restriction of “*” to $C$ is of order $2$. ∎
Note that “$\gcd(n,q)=1$” is assumed in this section. Recall from Eq.(2.15) that $e_{0}=\frac{1}{n}\sum_{i=0}^{n-1}x^{i}$, $e_{1},\dots,e_{m}$ are all primitive idempotents of ${\cal R}$. Then the $p^{h}$-automorphism “*” permutes the primitive idempotents, i.e., every $e_{i}^{*}$ is still a primitive idempotent. Note that $e_{0}^{*}=e_{0}$. We can reorder the other primitive idempotents
$e_{1},\dots,e_{r},\;e_{r+1},e_{r+1}^{*},\dots,e_{r+s},e_{r+s}^{*},$ (5.3)
such that $e_{i}^{*}=e_{i}$ for $i=1,\dots,r$, $e_{r+j}^{*}\neq e_{r+j}$ but $e_{r+j}^{**}=e_{r+j}$ for $j=1,\dots,s$, and $1+r+s=m$.
Remark 5.3.
(1) For $i=0,1,\dots,r$, we get $({\cal R}e_{i})^{*}={\cal R}e_{i}^{*}={\cal R}e_{i}$. The restriction of the map “*” in Eq.(5.1) to ${\cal R}e_{i}$ induces a $p^{\ell/2}$-automorphism of ${\cal R}e_{i}$ of order 2 (cf. Corollary 5.2) as follows
$*\!|_{{\cal R}e_{i}}:~{}{\cal R}e_{i}\rightarrow{\cal R}e_{i},~{}~{}~{}a\mapsto a^{*}.$
Let $d_{i}=\dim_{F}{\cal R}e_{i}$. By Eq.(2.19), the ideal ${\cal R}e_{i}$ is a field extension over $F$ with cardinality $|{\cal R}e_{i}|=p^{d_{i}\ell}$. So $a^{*}=a^{p^{d_{i}\ell/2}}$ for $a\in{\cal R}e_{i}$.
(2) For $j=1,\dots,s$, $({\cal R}e_{r+j})^{*}={\cal R}e_{r+j}^{*}\neq{\cal R}e_{r+j}$, and the restriction of the map “*” in Eq.(5.1) to ${\cal R}e_{r+j}$ induces a $p^{\ell/2}$-isomorphism as follows
$*\!|_{{\cal R}e_{r+j}}:~{}{\cal R}e_{r+j}\rightarrow{\cal R}e_{r+j}^{*},~{}~{}~{}a\mapsto a^{*}.$
Let $d_{r+j}=\dim_{F}{\cal R}e_{r+j}$, then $\dim_{F}{\cal R}e_{r+j}^{*}=\dim_{F}{\cal R}e_{r+j}=d_{r+j}$. Denote $\widehat{e}_{r+j}=e_{r+j}+e_{r+j}^{*}$, by Eq.(2.19) we have that
${\cal R}\widehat{e}_{r+j}={\cal R}e_{r+j}\oplus{\cal R}e_{r+j}^{*}=\{a^{\prime}+a^{\prime\prime}\,|\,a^{\prime}\in{\cal R}e_{r+j},\,a^{\prime\prime}\in{\cal R}e_{r+j}^{*}\},$
$\dim_{F}{\cal R}\widehat{e}_{r+j}=2d_{r+j}.$
It is easy to check that $\widehat{e}_{r+j}^{\,*}=(e_{r+j}+e_{r+j}^{*})^{*}=e_{r+j}^{*}+e_{r+j}=\widehat{e}_{r+j}$. So $({\cal R}\widehat{e}_{r+j})^{*}={\cal R}\widehat{e}_{r+j}^{\;*}={\cal R}\widehat{e}_{r+j}$. The restriction of the map “*” in Eq.(5.1) to ${\cal R}\widehat{e}_{r+j}$ induces a $p^{\ell/2}$-automorphism of ${\cal R}\widehat{e}_{r+j}$ of order 2 (cf. Corollary 5.2) as follows
$*\!|_{{\cal R}\widehat{e}_{r+j}}:~{}{\cal R}\widehat{e}_{r+j}\rightarrow{\cal R}\widehat{e}_{r+j},~{}~{}~{}a^{\prime}+a^{\prime\prime}\,\mapsto\,a^{\prime\prime*}+a^{\prime*}.$ (5.4)
For convenience, in the following we denote $\widehat{e}_{i}=e_{i}$ for $i=0,1,\dots,r$. Then $\widehat{e}_{i}^{\,*}=\widehat{e}_{i}$ for $i=0,1,\dots,r+s$, and
$1=\widehat{e}_{0}+\widehat{e}_{1}+\dots+\widehat{e}_{r+s},~{}~{}~{}~{}~{}\widehat{e}_{i}\widehat{e}_{j}=\begin{cases}\widehat{e}_{i},&i=j;\\ 0,&i\neq j.\end{cases}$ (5.5)
Thus, ${\cal R}$ can be rewritten as
${\cal R}={\cal R}\widehat{e}_{0}\oplus{\cal R}\widehat{e}_{1}\oplus\dots\oplus{\cal R}\widehat{e}_{r}\oplus{\cal R}\widehat{e}_{r+1}\oplus\dots\oplus{\cal R}\widehat{e}_{r+s}.$ (5.6)
Any $a\in{\cal R}$ is decomposed into
$a=a_{0}+a_{1}+\cdots+a_{r+s},~{}~{}~{}~{}~{}\mbox{where}~{}~{}a_{i}=a\widehat{e}_{i},~{}i=0,1,\dots,r+s.$ (5.7)
The $a_{i}=a\widehat{e}_{i}$ is called the $\widehat{e}_{i}$-component of $a$. For $a=\sum_{i=0}^{r+s}a_{i},b=\sum_{i=0}^{r+s}b_{i}\in{\cal R}$, by Eq.(5.5) it is trivial to check that
$ab=a_{0}b_{0}+a_{1}b_{1}+\cdots+a_{r+s}b_{r+s}.$ (5.8)
Keep the notation in Remark 5.3, we have $d_{i}=\dim_{F}{\cal R}\widehat{e}_{i}$ for $i=0,1,\dots,r$; and $2d_{i}=\dim_{F}{\cal R}\widehat{e}_{i}$ for $i=r+1,\dots,r+s$. Thus,
$\dim_{F}{{\cal R}}=d_{0}+d_{1}+\dots+d_{r}+2d_{r+1}+\dots+2d_{r+s}=n,$ (5.9)
where $d_{0}=1$ and $d_{i}\geq\mu(n)$ for $i=1,\dots,r+s$; cf. Eq.(2.21).
5.2 A class of Hermitian self-dual $2$-quasi-cyclic codes
Recall that the $2$-quasi-cyclic code $C_{1,g}$ of ${\cal R}^{2}$ (defined in Remark 3.4) is Hermitian self-dual if and only if $gg^{*}=-1$, see Lemma 4.9. So we denote
${\cal D}=\big{\{}\,g\;\big{|}\;g\in{\cal R},\;gg^{*}=-1\,\big{\}}.$ (5.10)
Any $g\in{\cal D}$ corresponds to a Hermitian self-dual $2$-quasi-cyclic code $C_{1,g}$. Set
${\cal D}_{i}=\{\,z\;|\;z\in{\cal R}\widehat{e}_{i},\;zz^{*}=-\widehat{e}_{i}\},~{}~{}~{}~{}~{}i=0,1,\dots,r+s.$ (5.11)
By Eq.(5.5), Eq.(5.7) and Eq.(5.8),
$g=g_{0}+g_{1}+\dots+g_{r+s}\in{\cal D}~{}\iff~{}g_{i}\in{\cal D}_{i},~{}\forall\;i=0,1,\dots,r+s.$ (5.12)
Lemma 5.4.
(1) If $0\leq i\leq r$ (i.e. $\widehat{e}_{i}=e_{i}=e_{i}^{*}$), then $|{\cal D}_{i}|=p^{d_{i}\ell/2}+1$.
(2) If $r<i\leq r+s$ (i.e. $e_{i}\neq e_{i}^{*}$ and $\widehat{e}_{i}=e_{i}+e_{i}^{*}$), then $|{\cal D}_{i}|=p^{d_{i}\ell}-1$.
Proof.
(1). By Remark 5.3(1), ${\cal R}\widehat{e}_{i}$ is the field with $|{\cal R}\widehat{e}_{i}|=p^{d_{i}\ell}$. For $z\in{\cal R}\widehat{e}_{i}$, $zz^{*}=zz^{p^{d_{i}\ell/2}}\!=z^{p^{d_{i}\ell/2}+1}$. So, $zz^{*}=-\widehat{e}_{i}$ if and only if ${z^{p^{d_{i}\ell/2}+1}+\widehat{e}_{i}=0}$. Hence $|{\cal D}_{i}|$ equals the number of the roots in $F_{p^{d_{i}\ell}}$ of the $F_{p^{d_{i}\ell}}$-polynomial $X^{p^{d_{i}\ell/2}+1}+1$, where $F_{p^{d_{i}\ell}}$ denotes the finite field with cardinality $p^{d_{i}\ell}$.
If $p=2$, then $1=-1$. The order of the multiplicative group $F_{p^{d_{i}\ell}}^{\times}$ is
$\big{|}F_{p^{d_{i}\ell}}^{\times}\big{|}=p^{d_{i}\ell}-1=\big{(}p^{d_{i}\ell/2}-1\big{)}\big{(}p^{d_{i}\ell/2}+1\big{)}.$ (5.13)
Thus the multiplication group $F_{p^{d_{i}\ell}}^{\times}$ has a subgroup $H$ of order $p^{d_{i}\ell/2}+1$, and all elements of $H$ are roots of the polynomial $X^{p^{d_{i}\ell/2}+1}+1$. Hence $|{\cal D}_{i}|=p^{d_{i}\ell/2}+1$.
Otherwise $p$ is odd, then $p^{d_{i}\ell/2}-1$ is even. By Eq.(5.13), we get that $2(p^{d_{i}\ell/2}+1)\,\big{|}\,|F_{p^{d_{i}\ell}}^{\times}|$. So $F_{p^{d_{i}\ell}}^{\times}$ has a subgroup $H$ of order $2(p^{d_{i}\ell/2}+1)$. The elements of $H$ are just all roots of the polynomial $X^{2(p^{d_{i}\ell/2}+1)}-1$. Since
$X^{2(p^{d_{i}\ell/2}+1)}-1=(X^{p^{d_{i}\ell/2}+1}-1)(X^{p^{d_{i}\ell/2}+1}+1),$
all roots of the polynomial $X^{p^{d_{i}\ell/2}+1}+1$ are inside $F_{p^{d_{i}\ell}}$. Thus, $|{\cal D}_{i}|=p^{d_{i}\ell/2}+1$.
(2). For $r<i\leq r+s$, $e_{i}\neq e_{i}^{*}$ and $\widehat{e}_{i}=e_{i}+e_{i}^{*}$. By Remark 5.3(2), ${\cal R}\widehat{e}_{i}={\cal R}e_{i}\oplus{\cal R}e_{i}^{*}$. For $z=z^{\prime}+z^{\prime\prime}\in{\cal R}e_{i}\oplus{\cal R}e_{i}^{*}$ with $z^{\prime}\in{\cal R}e_{i}$ and $z^{\prime\prime}\in{\cal R}e_{i}^{*}$, by Eq.(5.4) we have $z^{*}=z^{\prime\prime*}+z^{\prime*}$, and so $zz^{*}=(z^{\prime}+z^{\prime\prime})(z^{\prime\prime*}+z^{\prime*})=z^{\prime}z^{\prime\prime*}+z^{\prime*}z^{\prime\prime}$. It follows that: $zz^{*}=-\widehat{e}_{i}=-e_{i}-e_{i}^{*}$ if and only if $z^{\prime}z^{\prime\prime*}=-e_{i}$ and $z^{\prime*}z^{\prime\prime}=-e_{i}^{*}$. We take $z^{\prime}\in({\cal R}e_{i})^{\times}$, then $z^{\prime\prime*}=-z^{\prime-1}$ (where $z^{\prime-1}$ is the inverse of $z^{\prime}$ in ${\cal R}e_{i}$, not in ${\cal R}$), hence $z^{\prime\prime}=z^{\prime\prime**}=-(z^{\prime*})^{-1}$ is uniquely determined. Thus,
$zz^{*}=-\widehat{e}_{i}~{}\iff~{}z=z^{\prime}-(z^{\prime*})^{-1}~{}~{}\mbox{for a}~{}z^{\prime}\in({\cal R}e_{i})^{\times}.$ (5.14)
Since ${\cal R}e_{i}$ is a field with cardinality $p^{d_{i}\ell}$, $|{\cal D}_{i}|=|({\cal R}e_{i})^{\times}|=p^{d_{i}\ell}-1$. ∎
For any subset $\omega\subseteq\{0,1,\dots,r\!+\!s\}$, refining the notation in Eq.(5.6) and in Remark 5.3, we define an ideal $A_{\omega}$ of ${\cal R}$ and an integer $d_{\omega}$ as follows
$\textstyle A_{\omega}=\bigoplus\limits_{i\in\omega}{\cal R}\widehat{e}_{i},~{}~{}~{}~{}d_{\omega}=\dim_{F}A_{\omega}.$
Obviously, $\omega$ can be written as a disjoint union:
$\omega=\omega^{\prime}\cup\omega^{\prime\prime},~{}~{}~{}\mbox{where }\omega^{\prime}\!=\!\omega\cap\{0,1,\dots,r\},~{}\omega^{\prime\prime}\!=\!\omega\cap\{r\!+\!1,\dots,r\!+\!s\}.$
Similarly to Eq.(5.9), we have
$A_{\omega}=A_{\omega^{\prime}}\oplus A_{\omega^{\prime\prime}},~{}~{}~{}~{}d_{\omega}=\sum_{i\in\omega^{\prime}}d_{i}+2\sum_{i\in\omega^{\prime\prime}}d_{i}.$ (5.15)
It is known that ([20, Lemma 4.7]) for integers $k_{1},\dots,k_{v}$, if $k_{i}\geq\log_{p}(v)$ for $i=1,\dots,v$, then
$\begin{array}[]{l}(p^{k_{1}}-1)\dots(p^{k_{v}}-1)\geq p^{k_{1}+\dots+k_{v}-2};\\[3.0pt] (p^{k_{1}}+1)\dots(p^{k_{v}}+1)\leq p^{k_{1}+\dots+k_{v}+2}.\end{array}$ (5.18)
Lemma 5.5.
For a subset $\omega\subseteq\{0,1,\dots,r+s\}$, if $\mu(n)\geq\log_{p}(n)$ (where $\mu(n)$ is defined in Eq.(2.21)), then
$p^{-2}p^{d_{\omega}\ell/2}\leq\prod_{i\in\omega}|{\cal D}_{i}|\leq p^{3}p^{d_{\omega}\ell/2}.$
Proof.
For $1\leq i\leq r+s$, we deduce that $d_{i}\geq\mu(n)\geq\log_{p}(r+s)$ since $n\geq r+s$. By Lemma 5.4 we have that
$\prod_{i\in\omega}|{\cal D}_{i}|$ $=\prod_{i\in\omega^{\prime}}|{\cal D}_{i}|\cdot\prod_{i\in\omega^{\prime\prime}}|{\cal D}_{i}|=\prod\limits_{i\in\omega^{\prime}}(p^{d_{i}\ell/2}+1)\cdot\prod\limits_{i\in\omega^{\prime\prime}}(p^{d_{i}\ell}-1).$
If $0\notin\omega$, using Eq.(5.18) we get
$\prod_{i\in\omega}|{\cal D}_{i}|$ $\geq\prod\limits_{i\in\omega^{\prime}}(p^{d_{i}\ell/2}-1)\cdot\prod\limits_{i\in\omega^{\prime\prime}}(p^{d_{i}\ell}-1)\geq p^{\big{(}\sum_{i\in\omega^{\prime}}d_{i}\ell/2+\sum_{i\in\omega^{\prime\prime}}d_{i}\ell\big{)}-2}$
$=p^{-2}p^{\big{(}\sum_{i\in\omega^{\prime}}d_{i}+\sum_{i\in\omega^{\prime\prime}}2d_{i}\big{)}\ell/2}=p^{-2}p^{d_{\omega}\ell/2},$
where the last equality follows by Eq.(5.15); and
$\prod_{i\in\omega}|{\cal D}_{i}|$ $\leq\prod\limits_{i\in\omega^{\prime}}(p^{d_{i}\ell/2}+1)\cdot\prod\limits_{i\in\omega^{\prime\prime}}(p^{d_{i}\ell}+1)\leq p^{\big{(}\sum_{i\in\omega^{\prime}}d_{i}\ell/2+\sum_{i\in\omega^{\prime\prime}}d_{i}\ell\big{)}+2}$
$=p^{2}p^{\big{(}\sum_{i\in\omega^{\prime}}d_{i}+\sum_{i\in\omega^{\prime\prime}}2d_{i}\big{)}\ell/2}=p^{2}p^{d_{\omega}\ell/2}.$
That is, if $0\notin\omega$ then
$p^{-2}p^{d_{\omega}\ell/2}\leq\prod_{i\in\omega}|{\cal D}_{i}|\leq p^{2}p^{d_{\omega}\ell/2}.$ (5.19)
If $0\in\omega$, we set $\tilde{\omega}=\omega\setminus\{0\}$, and so $d_{\omega}=d_{0}+d_{\tilde{\omega}}$, where $d_{0}=1$. By Lemma 5.4(1), we see that $|{\cal D}_{0}|=p^{d_{0}\ell/2}+1$, and
$\prod_{i\in\omega}|{\cal D}_{i}|$ $=(p^{d_{0}\ell/2}+1)\prod\limits_{i\in\tilde{\omega}}|{\cal D}_{i}|.$
It follows by Eq.(5.19) that
$\prod_{i\in\omega}|{\cal D}_{i}|\geq p^{d_{0}\ell/2}\cdot p^{-2}p^{d_{\tilde{\omega}}\ell/2}$ $=p^{-2}p^{\big{(}d_{0}+d_{\tilde{\omega}}\big{)}\ell/2}=p^{-2}p^{d_{\omega}\ell/2};$
and that
$\prod_{i\in\omega}|{\cal D}_{i}|\leq p^{1+d_{0}\ell/2}\cdot p^{2}p^{d_{\tilde{\omega}}\ell/2}=p^{3}p^{\big{(}d_{0}+d_{\tilde{\omega}}\big{)}\ell/2}=p^{3}p^{d_{\omega}\ell/2}.$
Thus the lemma holds. ∎
By Eq.(5.12), $|{\cal D}|=\prod_{i=0}^{r+s}|{\cal D}_{i}|$. We have the following at once.
Corollary 5.6.
If $\mu(n)\geq\log_{p}(n)$ (where $\mu(n)$ is defined in Eq.(2.21)), then
$p^{-2}p^{n\ell/2}\leq|{\cal D}|\leq p^{3}p^{n\ell/2}.$
Any element $a\in{\cal R}$ can be written as $a=a_{0}\hat{e}_{0}+a_{1}\hat{e}_{1}+\dots+a_{r+s}\hat{e}_{r+s}$ (cf. Eq.(5.7)). We denote
$\omega_{a}=\big{\{}i\,\big{|}\,a_{i}=a\widehat{e}_{i}\neq 0\big{\}}\subseteq\{0,1,\dots,r+s\};~{}~{}d_{a}=d_{\omega_{a}}=\dim_{F}A_{\omega_{a}}.$
Obviously ${\cal R}a\subseteq A_{\omega_{a}}$. For $a,b\in{\cal R}$ it is easy to check that
${\cal R}a={\cal R}b~{}~{}\implies~{}~{}A_{\omega_{a}}=A_{\omega_{b}}~{}(\mbox{equivalently, }\omega_{a}=\omega_{b});$ (5.20)
but the converse is not true in general.
Lemma 5.7.
For $(a,b)\in{\cal R}^{2}$, we denote
${\cal D}_{(a,b)}=\big{\{}g\>\big{|}\;g\in{\cal D},\;(a,b)\in C_{1,g}\big{\}},$ (5.21)
where $C_{1,g}$ defined in Remark 3.4 and ${\cal D}$ defined in Eq.(5.10).
(1) If ${\cal D}_{(a,b)}\neq\emptyset$, then ${\cal R}a={\cal R}b$, hence $\omega_{a}=\omega_{b}$.
(2) If $\mu(n)\geq\log_{p}(n)$, then $|{\cal D}_{(a,b)}|\leq p^{5}p^{(n-d_{a})\ell/2}.$
Proof.
(1). Assume that $g\in{\cal D}_{(a,b)}$, then $gg^{*}=-1$ (cf. Eq.(5.10)) and $u(1,g)=(a,b)$ for an element $u\in{\cal R}$. The former equality implies that $g$ is invertible with $g^{-1}=-g^{*}$. The latter equality implies that $a=u$ and $b=ug=ag$. We deduce that ${\cal R}b={\cal R}ag={\cal R}ga={\cal R}a$ since $g$ is invertible.
(2). By Eq.(5.7), we write $a=a_{0}+\dots+a_{r+s}$, $b=b_{0}+\dots+b_{r+s}$ and $g=g_{0}+\dots+g_{r+s}$ with $g_{i}\in{\cal D}_{i}$. By Eq.(5.8),
$ag=b~{}~{}\iff a_{i}g_{i}=b_{i}~{}\mbox{for}~{}~{}i=0,1,\dots,r+s.$
We count the number of such $g_{i}$ in two cases.
Case 1: $a_{i}=0$, i.e., $i\notin\omega_{a}$. Since ${\cal R}a={\cal R}b$, we have that ${\cal R}b_{i}={\cal R}a_{i}=0$, hence $b_{i}=0$. Then any $g_{i}\in{\cal D}_{i}$ satisfies that $a_{i}g_{i}=b_{i}$.
Case 2: $a_{i}\neq 0$, i.e., $i\in\omega_{a}$. There are two subcases.
Subcase 2.1: $0\leq i\leq r$. Since ${\cal R}\widehat{e}_{i}$ is a field, $a_{i}$ is invertible in ${\cal R}\widehat{e}_{i}$, and so there is a unique $z\in{\cal R}\widehat{e}_{i}$ such that $a_{i}z=b_{i}$. We see that there is at most one $g_{i}\in{\cal D}_{i}$ such that $a_{i}g_{i}=b_{i}$.
Subcase 2.2: $r<i\leq r+s$. By Remark 5.3(2), ${\cal R}\widehat{e}_{i}={\cal R}e_{i}\oplus{\cal R}e_{i}^{*}$. We can write $a_{i}=\alpha^{\prime}+\alpha^{\prime\prime}$ and $b_{i}=\beta^{\prime}+\beta^{\prime\prime}$, where $\alpha^{\prime},\beta^{\prime}\in{\cal R}e_{i}$ and $\alpha^{\prime\prime},\beta^{\prime\prime}\in{\cal R}e_{i}^{*}$. Since $a_{i}\neq 0$, at least one of $\alpha^{\prime}$ and $\alpha^{\prime\prime}$ is nonzero. We may assume that $\alpha^{\prime}\neq 0$. Take $g_{i}=z^{\prime}-(z^{\prime*})^{-1}\in{\cal D}_{i}$ as in Eq.(5.14), then $a_{i}g_{i}=b_{i}$ implies that $\alpha^{\prime}z^{\prime}=\beta^{\prime}$ in ${\cal R}e_{i}$. Hence $z^{\prime}=\alpha^{\prime-1}\beta^{\prime}$ is uniquely determined.
In a word, if $a_{i}\neq 0$ (i.e., $i\in\omega_{a}$), then there is at most one $g_{i}\in{\cal D}_{i}$ such that $a_{i}g_{i}=b_{i}$. Thus
$|{\cal D}_{(a,b)}|\leq\prod_{i\notin\omega_{a}}|{\cal D}_{i}|=\prod_{i=0}^{r+s}|{\cal D}_{i}|\Big{/}\prod_{i\in\omega_{a}}|{\cal D}_{i}|=|{\cal D}|\Big{/}\prod_{i\in\omega_{a}}|{\cal D}_{i}|.$
By Lemma 5.5 and Corollary 5.6,
$|{\cal D}_{(a,b)}|\leq p^{3}p^{n\ell/2}\Big{/}p^{-2}p^{d_{\omega_{a}}\ell/2}=p^{5}p^{(n-d_{\omega_{a}})\ell/2}.$
We are done. ∎
5.3 Hermitian self-dual $2$-quasi-cyclic codes are good
Keep the notation in Subsection 5.2. From now on, let $\delta$ be a real number satisfying that (where $h_{q}(\delta)$ is the $q$-entropy function defined in Eq.(2.22))
$\delta\in(0,1-q^{-1})\quad\mbox{and}\quad h_{q}(\delta)<1/4.$ (5.22)
And set
$\begin{array}[]{l}{\cal D}^{\leq\delta}=\big{\{}\,g\;\big{|}\;g\in{\cal D},\;\Delta(C_{1,g})\leq\delta\,\big{\}};\\[3.0pt] \Omega_{d}=\big{\{}A~{}|~{}A~{}{\rm is~{}an~{}ideal~{}of~{}}{\cal R}{\rm~{}with~{}}\dim_{F}A=d\big{\}}.\end{array}$ (5.25)
Recall that $\mu(n)$ is defined in Eq.(2.21), $(A\times A)^{\leq\delta}$ is defined in Eq.(2.23) and ${\cal D}_{(a,b)}$ is defined in Eq.(5.21).
Lemma 5.8.
${\cal D}^{\leq\delta}\subseteq\bigcup_{d=\mu(n)}^{n}~{}\bigcup_{A\in\Omega_{d}}~{}\bigcup_{(a,b)\in(A\times A)^{\leq\delta}}{\cal D}_{(a,b)}$.
Proof.
Assume that $g\in{\cal D}^{\leq\delta}$. By Eq.(5.25), we have an $(a,b)\in{\cal R}^{2}$ such that $0<{\rm w}(a,b)\leq 2\delta n$ and $(a,b)\in C_{1,g}$, i.e., $g\in{\cal D}_{(a,b)}$. From Lemma 5.7(1) and Eq.(5.20), we deduce that $A_{\omega_{a}}=A_{\omega_{b}}$; hence $d_{a}\!=\!\dim_{F}A_{\omega_{a}}\!=\!\dim_{F}A_{\omega_{b}}=d_{b}$, and $(a,b)\in A_{\omega_{a}}\times A_{\omega_{a}}$ and $A_{\omega_{a}}\in\Omega_{d_{a}}$. If $A_{\omega_{a}}={\cal R}e_{0}$, then $0\neq a,b\in{\cal R}e_{0}$, hence ${\rm w}(a,b)=2n$ (cf. Eq.(2.20)), which contradicts that ${\rm w}(a,b)\leq 2\delta n$ and $\delta<1-q^{-1}$ (see Eq.(5.22)). Thus $A_{\omega_{a}}\neq{\cal R}e_{0}$. By the definition of $\mu(n)$ in Eq.(2.21), we obtain that $d_{a}\geq\mu(n)$. So the lemma is proved. ∎
Lemma 5.9.
If $\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{\mu(n)\ell}>0$ and $\mu(n)\geq\log_{p}(n)$, then
$|{\cal D}^{\leq\delta}|\leq p^{5}p^{\frac{n\ell}{2}-2\mu(n)\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{\mu(n)\ell}\big{)}}.$
Proof.
Note that any ideal $A$ of ${\cal R}$ is a direct sum of some of ${\cal R}e_{0},{\cal R}e_{1},\dots,{\cal R}e_{m}$ (cf. Eq.(2.19)). For $1\leq i\leq m$, the dimension $\dim_{F}{\cal R}e_{i}\geq\mu(n)$, where $\mu(n)$ is defined in Eq.(2.21). Suppose that $A\in{\Omega_{d}}$ where $\mu(n)\leq d\leq n$. If ${\cal R}e_{0}$ is not a direct summand of $A$, then the number of such $A$ is at most $m^{\frac{d}{\mu(n)}}$; otherwise ${\cal R}e_{0}$ is a direct summand of $A$, then the number of such $A$ is at most $m^{\frac{d-1}{\mu(n)}}<m^{\frac{d}{\mu(n)}}$. Hence
$|\Omega_{d}|\leq m^{\frac{d}{\mu(n)}}\leq n^{\frac{d}{\mu(n)}}.$ (5.26)
Applying Lemma 5.8 and Lemma 5.7(2), we obtain
$|{\cal D}^{\leq\delta}|$ $\leq\sum_{d=\mu(n)}^{n}~{}\sum_{A\in\Omega_{d}}~{}\sum_{(a,b)\in(A\times A)^{\leq\delta}}|{\cal D}_{(a,b)}|$
$\leq\sum_{d=\mu(n)}^{n}~{}\sum_{A\in\Omega_{d}}~{}\sum_{(a,b)\in(A\times A)^{\leq\delta}}p^{5}p^{(n-d)\ell/2},$
where the number $p^{5}p^{(n-d)\ell/2}$ is independent of the choice of $(a,b)\in(A\times A)^{\leq\delta}$. By Lemma 2.7, for $A\in\Omega_{d}$ we have $|(A\times A)^{\leq\delta}|\leq(p^{\ell})^{2dh_{q}(\delta)}$. So
$|{\cal D}^{\leq\delta}|\leq\sum_{d=\mu(n)}^{n}~{}\sum_{A\in\Omega_{d}}~{}|(A\times A)^{\leq\delta}|\cdot p^{5}p^{(n-d)\ell/2}$
$\leq\sum_{d=\mu(n)}^{n}\sum_{A\in\Omega_{d}}p^{2d\ell h_{q}(\delta)}p^{5}p^{(n-d)\ell/2}=p^{5}\!\sum_{d=\mu(n)}^{n}\sum_{A\in\Omega_{d}}p^{\frac{n\ell}{2}+2d\ell\big{(}h_{q}(\delta)-\frac{1}{4}\big{)}},$
where the number $p^{\frac{n\ell}{2}+2d\ell\big{(}h_{q}(\delta)-\frac{1}{4}\big{)}}$ is independent of the choice of $A\in\Omega_{d}$. Using Eq.(5.26) yields
$|{\cal D}^{\leq\delta}|\leq p^{5}\!\sum_{d=\mu(n)}^{n}\!|\Omega_{d}|\!\cdot\!p^{\frac{n\ell}{2}+2d\ell\big{(}h_{q}(\delta)-\frac{1}{4}\big{)}}\leq p^{5}\!\sum_{d=\mu(n)}^{n}\!n^{\frac{d}{\mu(n)}}p^{\frac{n\ell}{2}+2d\ell\big{(}h_{q}(\delta)-\frac{1}{4}\big{)}}$
$~{}=p^{5}\!\sum_{d=\mu(n)}^{n}\!p^{\frac{n\ell}{2}+2d\ell\big{(}h_{q}(\delta)-\frac{1}{4}\big{)}+\frac{d\log_{p}(n)}{\mu(n)}}=p^{5}\!\sum_{d=\mu(n)}^{n}p^{\frac{n\ell}{2}-2d\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{2\mu(n)\ell}\big{)}}.$
Note that $\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{2\mu(n)\ell}>0$ (since $\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{\mu(n)\ell}>0$) and $d\geq\mu(n)$. We further get
$|{\cal D}^{\leq\delta}|$ $\leq p^{5}\sum_{d=\mu(n)}^{n}p^{\frac{n\ell}{2}-2\mu(n)\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{2\mu(n)\ell}\big{)}}$
$\leq p^{5}\cdot n\cdot p^{\frac{n\ell}{2}-2\mu(n)\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{2\mu(n)\ell}\big{)}}.$
That is, $|{\cal D}^{\leq\delta}|\leq p^{5}p^{\frac{n\ell}{2}-2\mu(n)\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{\mu(n)\ell}\big{)}}$. ∎
By [3, Lemma 2.6] (or [12, Lemma II.6]), there are odd positive integers $n_{1},n_{2},\dots$ coprime to $q=p^{\ell}$ such that $\lim\limits_{i\to\infty}\frac{\log_{q}(n_{i})}{\mu(n_{i})}=0$, where $\mu(n)$ is defined in Eq.(2.21). Since $\log_{q}(n_{i})=\log_{p}(n_{i})/\log_{p}(q)=\log_{p}(n_{i})/\ell$, we see that there are odd positive integers $n_{1},n_{2},\dots$ coprime to $p$ such that
$\lim\limits_{i\to\infty}\frac{\log_{p}(n_{i})}{\mu(n_{i})}=0,$ (5.27)
which implies that $\mu(n_{i})\to\infty$, hence $n_{i}\to\infty$. Obviously, we can assume that $\mu(n_{i})\geq\log_{p}(n_{i})$ for $i=1,2,\dots$.
Theorem 5.10.
Assume that $0<\delta<1-q^{-1}$ and $h_{q}(\delta)<\frac{1}{4}$ as in Eq.(5.22). Then there are Hermitian self-dual $2$-quasi-cyclic codes $C_{1},C_{2},\dots$ over $F$ (hence ${\rm R}(C_{i})=\frac{1}{2}$) such that the code length $2n_{i}$ of $C_{i}$ goes to infinity and the relative minimum distance $\Delta(C_{i})>\delta$ for $i=1,2,\dots$.
Proof.
Take $n_{1},n_{2},\dots$ as in Eq.(5.27). There is a positive real number $\varepsilon>0$ such that $\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n_{i})}{\mu(n_{i})\ell}\geq\varepsilon$ for large enough index $i$. So we can further assume that
$\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n_{i})}{\mu(n_{i})\ell}\geq\varepsilon,~{}~{}~{}~{}i=1,2,\dots.$ (5.28)
Taking $n=n_{i}$ in Corollary 5.6 and Lemma 5.9, and denoting ${\cal D}$ by ${\cal D}(i)$, we get
$\frac{|{\cal D}(i)^{\leq\delta}|}{|{\cal D}(i)|}\leq\frac{p^{5}p^{\frac{n_{i}\ell}{2}-2\mu(n_{i})\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n_{i})}{\mu(n_{i})\ell}\big{)}}}{p^{-2}p^{\frac{n_{i}\ell}{2}}}=p^{7}p^{-2\mu(n_{i})\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n_{i})}{\mu(n_{i})\ell}\big{)}}.$
By Eq.(5.28) and that $\mu(n_{i})\to\infty$, we get that
$\lim_{i\to\infty}|{\cal D}(i)^{\leq\delta}|\big{/}|{\cal D}(i)|\leq\lim_{i\to\infty}p^{7}p^{-2\mu(n_{i})\ell\varepsilon}=0.$
Therefore, we can further assume that $|{\cal D}(i)^{\leq\delta}|<|{\cal D}(i)|$ for $i=1,2,\dots$. So we can take $g_{i}\in{\cal D}(i)\setminus{\cal D}(i)^{\leq\delta}$. Then $C_{i}=C_{1,g_{i}}$ is a Hermitian self-dual $2$-quasi-cyclic code of length $2n_{i}$ with $\Delta(C_{i})>\delta$. ∎
6 Hermitian (Euclidean) self-dual $2$-quasi
constacyclic codes
[…]
Lemma 6.1.
Assume that ${\lambda^{1+p^{h}}=1}$ and $\gcd(n,t)=1$, where $t={\rm ord}_{F^{\times}}(\lambda)$ as in Eq.(3.1). Then there is a $\gamma=\lambda^{s}\in F^{\times}$ such that $\gamma^{n}=\lambda^{-1}$ and the map
$\eta:~{}{\cal R}\,\longrightarrow\,{\cal R}_{\kern-1.0pt\lambda},\quad\sum_{i=0}^{n-1}a_{i}X^{i}\,\longmapsto\,\sum_{i=0}^{n-1}a_{i}(\gamma X)^{i}~{}({\rm mod}~{}X^{n}-\lambda),$ (6.1)
satisfies the following:
(1) $\eta$ is an algebra isomorphism.
(2) The weight ${\rm w}\big{(}\eta(a(X))\big{)}={\rm w}\big{(}a(X)\big{)}$,   $\forall$ $a(X)\in{\cal R}$.
(3) $\big{\langle}\eta(a(X)),\,\eta(b(X))\big{\rangle}_{h}=\big{\langle}a(X),\,b(X)\big{\rangle}_{h}$,   $\forall$ $a(X),b(X)\in{\cal R}$.
Proof.
Since $\gcd(n,t)=1$, there are integers $s,s^{\prime}$ such that $ns+ts^{\prime}=-1$. We deduce that $\lambda^{-1}=\lambda^{ns+ts^{\prime}}=\lambda^{ns}\lambda^{ts^{\prime}}=\lambda^{ns}$ since $\lambda^{t}=1$. Take $\gamma=\lambda^{s}$, then $\gamma^{n}=\lambda^{-1}$. The following is a surjective algebra homomorphism:
$\tilde{\eta}:F[X]\,\longrightarrow\,F[X]/\langle X^{n}-\lambda\rangle,\quad f(X)\,\longmapsto\,f(\gamma X)~{}({\rm mod}~{}X^{n}-\lambda),$
and
$\tilde{\eta}(X^{n}-1)=\gamma^{n}X^{n}-1=\lambda^{-1}X^{n}-1=\lambda^{-1}(X^{n}-\lambda)\equiv 0~{}~{}({\rm mod}~{}X^{n}-\lambda).$
Thus the kernel of $\tilde{\eta}$ is the ideal $\langle X^{n}-1\rangle$, and the surjective homomorphism $\tilde{\eta}$ induces the isomorphism $\eta$ in Eq.(6.1). So (1) is proved.
For any $a(X)\!=\!\sum_{i=0}^{n-1}a_{i}X^{i}\in{\cal R}$, by Eq.(6.1) we get $\eta\big{(}a(X)\big{)}\!=\!\sum_{i=0}^{n-1}a_{i}\gamma^{i}X^{i}$. So we see that $a_{i}\neq 0$ if and only if $a_{i}\gamma^{i}\neq 0$, hence ${\rm w}\big{(}\eta(a(X))\big{)}={\rm w}\big{(}a(X)\big{)}$. Thus (2) holds.
Let $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}\in{\cal R}$ and $b(X)=\sum_{i=0}^{n-1}b_{i}X^{i}\in{\cal R}$. Then $\eta\big{(}a(X)\big{)}=\sum_{i=0}^{n-1}a_{i}\gamma^{i}X^{i}$ and $\eta\big{(}b(X)\big{)}=\sum_{i=0}^{n-1}b_{i}\gamma^{i}X^{i}$. By Eq.(3.15), we have that
$\big{\langle}\eta(a(X)),\eta(b(X))\big{\rangle}_{h}=\sum_{i=0}^{n-1}(a_{i}\gamma^{i})\cdot(b_{i}\gamma^{i})^{p^{h}}=\sum_{i=0}^{n-1}a_{i}b_{i}^{p^{h}}\cdot\gamma^{(1+p^{h})i}.$
Applying the equality $\gamma=\lambda^{s}$ and the assumption $\lambda^{1+p^{h}}=1$, we get
${\gamma^{1+p^{h}}=(\lambda^{s})^{1+p^{h}}=(\lambda^{1+p^{h}})^{s}}=1.$
Thus
$\big{\langle}\eta(a(X)),\eta(b(X))\big{\rangle}_{h}=\sum_{i=0}^{n-1}a_{i}b_{i}^{p^{h}}=\big{\langle}a(X),b(X)\big{\rangle}_{h},$
which proves (3). ∎
Corollary 6.2.
Assume that ${\lambda^{1+p^{h}}=1}$ and $\gcd(n,t)=1$, where $t={\rm ord}_{F^{\times}}(\lambda)$ as in Eq.(3.1). Let
$\eta^{(2)}:~{}{\cal R}^{2}\,\longrightarrow\,{\cal R}_{\kern-1.0pt\lambda}^{2},~{}~{}\big{(}a(X),a^{\prime}(X)\big{)}\,\longmapsto\,\big{(}\eta(a(X)),\eta(a^{\prime}(X))\big{)},$ (6.2)
where $\eta$ is defined in Lemma 6.1. Then the following hold.
(1) $\eta^{(2)}$ is a module isomorphism.
(2) ${\rm w}\big{(}\eta^{(2)}(a(X),a^{\prime}(X))\big{)}={\rm w}\big{(}a(X),a^{\prime}(X)\big{)}$$\forall\;(a(X),a^{\prime}(X))\in{\cal R}^{2}$.
(3) $\!\big{\langle}\eta^{(2)}(a(X),a^{\prime}(X)),\eta^{(2)}(b(X),b^{\prime}(X))\big{\rangle}_{h}\!=\!\big{\langle}(a(X),a^{\prime}(X)),\!(b(X),b^{\prime}(X))\big{\rangle}_{h}$,  $\forall\;(a(X),a^{\prime}(X)),(b(X),b^{\prime}(X))\in{\cal R}^{2}$.
Proof.
(1)  For $\big{(}a(X),a^{\prime}(X)\big{)}\in{{\cal R}}^{2}$ and $f(X)\in{{\cal R}}$, by Lemma 6.1(1),
$\eta^{(2)}\big{(}f(X)\cdot(a(X),a^{\prime}(X))\big{)}=\big{(}\eta(f(X)a(X)),\eta(f(X)a^{\prime}(X))\big{)}$
$=\big{(}\eta(f(X))\eta(a(X)),\eta(f(X))\eta(a^{\prime}(X))\big{)}=\eta(f(X))\cdot\eta^{(2)}\big{(}a(X),a^{\prime}(X)\big{)}.$
Thus Eq.(6.2) is a module homomorphism. Observe that by Lemma 6.1(1), $\eta^{(2)}$ must be bijective, hence it is a module isomorphism.
(2)  By Lemma 6.1(2), we get
${\rm w}\big{(}\eta^{(2)}(a(X),a^{\prime}(X))\big{)}$ $={\rm w}\big{(}\eta(a(X)),\eta(a^{\prime}(X))\big{)}={\rm w}\big{(}\eta(a(X))\big{)}+{\rm w}\big{(}\eta(a^{\prime}(X))\big{)}$
$={\rm w}\big{(}a(X)\big{)}+{\rm w}\big{(}a^{\prime}(X)\big{)}={\rm w}\big{(}a(X),a^{\prime}(X)\big{)}.$
(3)  For $(a(X),a^{\prime}(X))\in{{\cal R}^{2}}$, $(b(X),b^{\prime}(X))\in{{\cal R}^{2}}$, where $a(X)\!=\sum_{i=0}^{n-1}a_{i}X^{i}$, $a^{\prime}(X)\!=\sum_{i=0}^{n-1}a^{\prime}_{i}X^{i}$ and $b(X)=\sum_{i=0}^{n-1}b_{i}X^{i}$, $b^{\prime}(X)=\sum_{i=0}^{n-1}b^{\prime}_{i}X^{i}$, it is obvious that
$\big{\langle}(a(X),a^{\prime}(X)),(b(X),b^{\prime}(X))\big{\rangle}_{h}$ $=\sum_{i=0}^{n-1}a_{i}{b_{i}}^{p^{h}}+\sum_{i=0}^{n-1}a^{\prime}_{i}{b^{\prime}_{i}}^{p^{h}}$
$=\big{\langle}a(X),b(X)\big{\rangle}_{h}+\big{\langle}a^{\prime}(X),b^{\prime}(X)\big{\rangle}_{h}.$
Then
$\big{\langle}\eta^{(2)}(a(X),a^{\prime}(X)),\;\eta^{(2)}(b(X),b^{\prime}(X))\big{\rangle}_{h}$
$=$ $\big{\langle}\big{(}\eta(a(X)),\eta(a^{\prime}(X))\big{)},\;\big{(}\eta(b(X)),\eta(b^{\prime}(X))\big{)}\big{\rangle}_{h}$
$=$ $\big{\langle}\eta(a(X)),\eta(b(X))\big{\rangle}_{h}+\big{\langle}\eta(a^{\prime}(X)),\eta(b^{\prime}(X))\big{\rangle}_{h}$
[…]
$=$ $\big{\langle}(a(X),a^{\prime}(X)),\;(b(X),b^{\prime}(X))\big{\rangle}_{h}.$
We are done. ∎
Theorem 6.3.
Assume that ${\lambda^{1+p^{h}}=1}$ and $\gcd(n,t)=1$, where $t={\rm ord}_{F^{\times}}\!(\lambda)$ as in Eq.(3.1). Let $\eta^{(2)}$ be as in Eq.(6.2). Then $C$ is an ${\cal R}$-submodule of ${\cal R}^{2}$ if and only if $\eta^{(2)}(C)$ is an ${\cal R}_{\kern-1.0pt\lambda}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$. At that case, the following hold.
(1) The rate ${\rm R}\big{(}\eta^{(2)}(C)\big{)}={\rm R}(C)$.
(2) The relative minimum distance $\Delta\big{(}\eta^{(2)}(C)\big{)}=\Delta(C)$.
(3) $\eta^{(2)}(C)$ is Galois $p^{h}$-self-dual if and only if $C$ is Galois $p^{h}$-self-dual.
Proof.
By Corollary 6.2(1), the map $\eta^{(2)}$ in Eq.(6.2) is a module isomorphism. So $C$ is an ${\cal R}$-submodule of ${\cal R}^{2}$ if and only if $\eta^{(2)}(C)$ is an ${\cal R}_{\kern-1.0pt\lambda}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$. Assume that it is this case.
(1) Since $\eta^{(2)}$ is an isomorphism, $\dim_{F}\eta^{(2)}(C)=\dim_{F}C$, and so
${\rm R}\big{(}\eta^{(2)}(C)\big{)}=\frac{\dim_{F}\eta^{(2)}(C)}{2n}=\frac{\dim_{F}C}{2n}={\rm R}(C).$
(2)  It holds by Corollary 6.2(2) obviously.
(3)  Observe that by Corollary 6.2(3), $\big{\langle}\eta^{(2)}(C),\eta^{(2)}(C)\big{\rangle}_{h}=0$ in ${\cal R}_{\kern-1.0pt\lambda}^{2}$ if and only if $\big{\langle}C,\,C\big{\rangle}_{h}=0$ in ${\cal R}^{2}$. Further, applying the above conclusion (1) yields that ${\rm R}\big{(}\eta^{(2)}(C)\big{)}=\frac{1}{2}$ if and only if ${\rm R}(C)=\frac{1}{2}$. Thus (3) holds. ∎
In the following, we consider the asymptotic property of Hermitian (Euclidean) self-dual $2$-quasi $\lambda$-constacyclic codes.
Theorem 6.4.
Assume that $\ell$ is even. The Hermitian self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ are asymptotically good if and only if $\lambda^{1+p^{\ell/2}}=1$.
Proof.
If $\lambda^{1+p^{\ell/2}}\neq 1$, then Hermitian self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ are asymptotically bad, see Theorem 4.5.
Assume that $\lambda^{1+p^{\ell/2}}=1$. Let $0<\delta<1-q^{-1}$ and $h_{q}(\delta)<\frac{1}{4}$. By Theorem 5.10, there are Hermitian self-dual $2$-quasi-cyclic codes $C_{1},C_{2},\dots$ over $F$ such that:
  • the code length $2n_{i}$ of $C_{i}$ satisfy that $n_{i}$ is odd and coprime to $q$ and $\lim\limits_{i\to\infty}\frac{\log_{p}(n_{i})}{\mu(n_{i})}=0$; in particular, $\mu(n_{i})\to\infty$, and hence $n_{i}\to\infty$;
  • the rate ${\rm R}(C_{i})=\frac{1}{2}$ and the relative minimum distance $\Delta(C_{i})>\delta$  for $i=1,2,\dots$.
Let $\tilde{t}=\max\big{\{}\mu(r)\,\big{|}\,\mbox{$r$ runs over the prime divisors of $t$}\big{\}}$, where $t={\rm ord}_{F^{\times}}\!(\lambda)$ as in Eq.(3.1). By [20, Lemma II.2],
$\mu(n_{i})=\min\big{\{}\mu(r)\;|\;\mbox{$r$ runs over the prime divisors of $n_{i}$}\big{\}}.$
If $\gcd(n_{i},t)\neq 1$, then $\mu(n_{i})\leq\tilde{t}$. Since $\mu(n_{i})\to\infty$, there are only finitely many $n_{i}$ such that $\gcd(n_{i},t)\neq 1$. Removing such $n_{i}$, we can further assume that $\gcd(n_{i},t)=1$ for $i=1,2,\dots$. Thus, applying the isomorphism Eq.(6.2) to $C_{i}$ yields $\tilde{C}_{i}=\eta^{(2)}(C_{i})$ in ${\cal R}_{\kern-1.0pt\lambda}^{2}$, $i=1,2,\dots$. We get the code sequence
$\tilde{C}_{1},~{}\tilde{C}_{2},~{}\dots.$
By Theorem 6.3, each $\tilde{C}_{i}$ is Hermitian self-dual $2$-quasi $\lambda$-constacyclic codes over $F$, and the code length $2n_{i}$ goes to infinity, the rate ${\rm R}(\tilde{C}_{i})=\frac{1}{2}$ and the relative minimum distance $\Delta(\tilde{C}_{i})=\Delta(C_{i})>\delta$ for $i=1,2,\dots$. ∎
For Euclidean case, the “Euclidean self-dual” is referred to as “self-dual”.
Theorem 6.5.
[…]
Proof.
By Theorem 4.5, if $\lambda^{2}\neq 1$ then the self-dual $2$-quasi $\lambda$-constacyclic codes are asymptotically bad. In the following we assume that $\lambda=\pm 1$
If $\lambda=1$, it has been shown in [20] (cf. Remark 2.5) that the self-dual $2$-quasi-cyclic codes over $F$ are asymptotically good.
[…]
  • the code length $2n_{i}$ of $C_{i}$ satisfy that $n_{i}$ is odd and coprime to $q$, and $\lim\limits_{i\to\infty}\frac{\log_{p}(n_{i})}{\mu(n_{i})}=0$; in particular, $\mu(n_{i})\to\infty$, and hence $n_{i}\to\infty$;
  • the rate ${\rm R}(C_{i})=\frac{1}{2}$ and the relative minimum distance $\Delta(C_{i})>\delta$  for $i=1,2,\dots$.
Similarly to the proof of Theorem 6.4, we get self-dual $2$-quasi negacyclic codes
$\tilde{C}_{1},~{}\tilde{C}_{2},~{}\dots$
over $F$ such that their code length $2n_{i}$ goes to infinity, ${\rm R}(\tilde{C}_{i})=\frac{1}{2}$ and their relative minimum distance $\Delta(\tilde{C}_{i})=\Delta(C_{i})>\delta$ for $i=1,2,\dots$. ∎
7 Conclusions
The purpose of this paper is to characterize the Galois self-dual $2$-quasi $\lambda$-constacyclic codes, and to investigate their asymptotic properties.
We first showed the algebraic structure of $2$-quasi $\lambda$-constacyclic codes. Then we found that the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes behave much differently according to whether $\lambda^{1+p^{h}}\!\neq 1$ ($\lambda\neq\lambda^{-p^{\ell-h}}$ equivalently) or ${\lambda^{1+p^{h}}\!=1}$ ($\lambda=\lambda^{-p^{\ell-h}}$ equivalently). In both the cases, we exhibited the necessary and sufficient conditions for the $2$-quasi $\lambda$-constacyclic codes being Galois $p^{h}$-self-dual, see Theorem 4.1 (for the former case), Lemma 4.9 and Theorem 4.10 (for the latter case). And in the former case we proved that the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes are asymptotically bad.
[…]