### v2

# Galois Self-dual 2-quasi Constacyclic Codes over Finite Fields

Abstract
[…]
Key words: Finite fields; $2$-quasi constacyclic codes; Galois self-dual codes; Hermitian self-dual codes; asymptotic property.
Email address: yfan@mail.ccnu.edu.cn (Yun Fan); yueleng0213@126.com (Yue Leng).
1 Introduction
Let $F$ be a finite field with cardinality $|F|=q=p^{\ell}$ where $p$ is a prime and $\ell$ is a positive integer, let $F^{\times}=F\setminus\{0\}$. Any ${\bf a}=(a_{0},a_{1},\dots,a_{n-1})\in{F}^{n}$ is called a word over $F$ of length $n$, where $n$ is a positive integer. The Hamming weight ${\rm w}({\bf a})$ is defined as the number of the indexes $i$ with $a_{i}\neq 0$. The Hamming distance between two words ${\bf a},{\bf a}^{\prime}\in F^{n}$ is defined as ${\rm d}({\bf a},{\bf a}^{\prime})={\rm w}({\bf a}-{\bf a}^{\prime})$. Any nonempty subset $C\subseteq F^{n}$ is called a code of length $n$ over $F$, and the words in the code are called codewords. The minimum distance ${\rm d}(C):=\min\limits_{{\bf c}\neq{\bf c}^{\prime}\in C}{\rm d}({\bf c},{\bf c}^{\prime})$. For any linear subspace $C$ of $F^{n}$, called a linear code, the minimum weight ${\rm w}(C):=\min\limits_{0\neq{\bf c}\in C}{\rm w}({\bf c})$; and it is known that ${\rm w}(C)={\rm d}(C)$. The fraction $\Delta(C)=\frac{{\rm d}(C)}{n}=\frac{{\rm w}(C)}{n}$ is called the relative minimum distance of $C$, and ${\rm R}(C)=\frac{\dim_{F}C}{n}$ is called the rate of $C$. A code sequence $C_{1},C_{2},\dots$ is said to be asymptotically good if the length $n_{i}$ of $C_{i}$ goes to infinity and there is a real number $\delta>0$ such that ${\rm R}(C_{i})>\delta$ and $\Delta(C_{i})>\delta$ for $i=1,2,\dots$. A class of codes is said to be asymptotically good if there is an asymptotically good sequence of codes in the class; otherwise, we say that the class of codes is asymptotically bad.
A linear code $C$ of $F^{n}$ is called a cyclic code if $C$ is invariant under the cyclic permutation on items, i.e.,
$(c_{0},c_{1},\ldots,c_{n-1})\in{C}\implies(c_{n-1},c_{0},c_{1},\dots,c_{n-2})\in{C}.$
A linear code $C$ of $F^{n}\times F^{n}$ is called a quasi-cyclic code of index $2$, abbreviated as $2$-quasi-cyclic code, if $C$ is invariant under the double cyclic permutation on items, i.e.,
$\begin{array}[]{l}(c_{0},c_{1},\dots,c_{n-1},c^{\prime}_{0},c^{\prime}_{1},\dots,c^{\prime}_{n-1})\in{C}\\[1.0pt] \implies(c_{n-1},c_{0},c_{1},\dots,c_{n-2},\,c^{\prime}_{n-1},c^{\prime}_{0},c^{\prime}_{1},\dots,c^{\prime}_{n-2})\in{C}.\end{array}$
The Euclidean inner product of words ${\bf a}=(a_{0},\dots,a_{n-1}),{\bf b}=(b_{0},\dots,b_{n-1})$ of $F^{n}$ is defined to be $\langle{\bf a},{\bf b}\rangle=\sum_{i=0}^{n-1}a_{i}b_{i}$. For a code $C\subseteq F^{n}$, the $C^{\bot}=\{\,{\bf a}\in{F^{n}\,|\,\langle{\bf c},{\bf a}\rangle=0,\,\forall{\bf c}\in{C}}\,\}$ is the dual code of $C$. A code $C$ is said to be self-dual if $C=C^{\bot}$. Obviously, the rate $R(C)=\frac{1}{2}$ if $C$ is self-dual.
[…]
Cyclic codes and $2$-quasi-cyclic codes had been extended widely. Let $\lambda\in F^{\times}=F\setminus\{0\}$. A linear code $C$ of $F^{n}$ is called a $\lambda$-constacyclic code if $C$ is invariant under the $\lambda$-constacyclic permutation on items, i.e.,
$(c_{0},c_{1},\ldots,c_{n-1})\in C\implies(\lambda c_{n-1},c_{0},c_{1},\dots,c_{n-2})\in C.$ (1.1)
If $\lambda=-1$, the $\lambda$-constacyclic codes are called negacyclic codes. Further, a linear code $C$ of $F^{n}\times F^{n}$ is called a $2$-quasi $\lambda$-constacyclic code if $C$ is invariant under the double $\lambda$-constacyclic permutation on items, i.e.,
$\begin{array}[]{l}(c_{0},c_{1},\dots,c_{n-1},c^{\prime}_{0},c^{\prime}_{1},\dots,c^{\prime}_{n-1})\in{C}\\[1.0pt] \implies(\lambda c_{n-1},c_{0},c_{1},\dots,c_{n-2},\,\lambda c^{\prime}_{n-1},c^{\prime}_{0},c^{\prime}_{1},\dots,c^{\prime}_{n-2})\in{C}.\end{array}$ (1.4)
If $q$ is odd and $q\,{\not\equiv}\pm 1~{}({\rm mod}~{}8)$, the self-dual $2$-quasi negacyclic codes over $F$ are proved asymptotically good in [2]. While in [28] for $q\,{\equiv}-1~{}({\rm mod}~{}4)$ it is shown, based on Artin’s primitive root conjecture, that the $q$-ary self-dual $2$-quasi negacyclic codes are asymptotically good. Recently, for any $q$ and any $\lambda\in F^{\times}$ the $2$-quasi $\lambda$-constacyclic codes over $F$ are proved asymptotically good, see [12, Corollary I.3].
About the self-dualities, in the semisimple case (i.e., $\gcd(n,q)=1$), the self-dual cyclic codes over $F$ does not exist. Leon et al. [21] and many references, e.g. [8, 25, 29], devoted to the study on various generalizations, e.g., duadic codes, extended self-dual cyclic codes, etc. On the other hand, Dinh and Lopez-Permouth [10], Dinh [9] studied $\lambda$-constacyclic codes, and showed that in the semisimple case the self-dual $\lambda$-constacyclic codes exist only if $\lambda=-1$.
Extending the Euclidean inner product and the Hermitian inner product, Fan and Zhang [16] introduced the so-called Galois inner products. Recall that $|F|=q=p^{\ell}$. For $0\leq h<\ell$, the map $\sigma_{p^{h}}:F\to F$, $\alpha\mapsto\alpha^{p^{h}}$, is a Galois automorphism of $F$, which induces an automorphism $F^{n}\to F^{n}$, $(a_{0},\dots,a_{n-1})\mapsto(a_{0}^{p^{h}},\dots,a_{n-1}^{p^{h}})$. The following
$\langle{\bf a},{\bf b}\rangle_{h}=\sum_{i=0}^{n-1}a_{i}b_{i}^{p^{h}},~{}~{}~{}~{}\forall\,{\bf a}=(a_{0},\cdots,a_{n-1}),{\bf b}=(b_{0},\cdots,b_{n-1})\in F^{n},$ (1.5)
is called the Galois $p^{h}$-inner product on $F^{n}$. And for any code $C\subseteq F^{n}$, the following code
$C^{\bot{h}}=\big{\{}\,{\bf a}\in F^{n}\,\big{|}\,\langle{\bf c},{\bf a}\rangle_{h}=0,\;\forall\,{\bf c}\in C\big{\}}$ (1.6)
is called the Galois $p^{h}$-dual code of $C$. The code $C$ is said to be Galois $p^{h}$-self-dual (or Galois self-dual when $h$ is known from context) if $C=C^{\bot h}$. It is also obvious that $R(C)=\frac{1}{2}$ if $C$ is Galois $p^{h}$-self-dual. When $h=0$ ($h=\ell/2$ when $\ell$ is even, respectively), $\langle{\bf a},{\bf b}\rangle_{h}$ is just the Euclidean (Hermitian, respectively) inner product, and $C^{\bot{h}}$ is the Euclidean (Hermitian, respectively) dual code, and Galois $p^{h}$-self-dual codes are just the Euclidean self-dual (Hermitian self-dual, respectively) codes. The existence and the structure of Galois $p^{h}$-self-dual $\lambda$-constacyclic codes are studied in [16].
In this paper we study the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ and their asymptotic properties. The main contributions of this paper are the following.
  • We characterize the algebraic structure of the $2$-quasi $\lambda$-constacyclic codes and their Galois $p^{h}$-dual codes. We find that the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes behave very differently depending on whether ${\lambda^{1+p^{h}}=1}$ or not. In both the cases we obtain necessary and sufficient conditions for $2$-quasi $\lambda$-constacyclic codes being Galois $p^{h}$-self-dual.
[…]
In Section 2, some preliminaries are sketched.
In Section 3 we characterize the algebraic structure of the $2$-quasi $\lambda$-consta-cyclic codes over the finite field $F$ and their Galois $p^{h}$-dual codes.
In Section 4 we study the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$. Our discussion divide into two cases: $\lambda^{1+p^{h}}\!\neq 1$ or $\lambda^{1+p^{h}}\!=1$. In both the cases, we exhibit the necessary and sufficient conditions for a $2$-quasi $\lambda$-constacyclic code being Galois $p^{h}$-self-dual. And we show that if $\lambda^{1+p^{h}}\!\neq 1$, then Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes are asymptotically bad.
In Section 5, the Hermitian self-dual $2$-quasi-cyclic codes over $F$ are proved asymptotically good.
[…]
Finally, we end this paper by a conclusion in Section 7.
2 Preliminaries
In this paper, $F$ is always a finite field of cardinality $|F|=q=p^{\ell}$ (by $|S|$ we denote the cardinality of any set $S$), where $p$ is a prime and $\ell$ is a positive integer; and $h$ is an integer such that $0\leq h<\ell$; and $n>1$ is an integer. Any ring $R$ in this paper has identity $1_{R}$ (or denoted by $1$ for short); and ring homomorphisms and subrings are identity preserving. By $R^{\times}$ we denote the multiplication group consisting of all units (invertible elements) of $R$. In particular, $F^{\times}=F\,{\setminus}\,\{0\}$. If a ring $R$ is also an $F$-vector space, then $R$ is said to be an $F$-algebra. In that case, $F\to R$, $\alpha\mapsto\alpha 1_{R}$, is an embedding, so that we write that $F\subseteq R$.
Let $R$ and $S$ be $F$-algebras. A map $\psi:R\to S$ is called an $F$-algebra homomorphism if it is both a ring homomorphism and an $F$-linear map, i.e., $\psi(1_{R})=1_{S}$ and for any $a,b\in R$ and any $\alpha\in F$,
$\psi(ab)=\psi(a)\psi(b),~{}~{}\psi(a+b)=\psi(a)+\psi(b),~{}~{}\psi(\alpha a)=\alpha\psi(a),$
where the first two mean that $\psi$ is a ring homomorphism, and the last two mean that it is a linear map. Recall that for $0\leq h<\ell$, the map $\sigma_{p^{h}}:F\to F$, $\sigma_{p^{h}}(\alpha)=\alpha^{p^{h}}$ for $\alpha\in F$, is a Galois automorphism of the field $F$. If $\psi:R\to S$ is bijective and satisfies that for any $a,b\in R$ and any $\alpha\in F$,
$\psi(ab)=\psi(a)\psi(b),~{}~{}\psi(a+b)=\psi(a)+\psi(b),~{}~{}\psi(\alpha a)=\alpha^{p^{h}}\psi(a),$ (2.1)
then $\psi$ is called a $\sigma_{p^{h}}$-algebra isomorphism, or $p^{h}$-isomorphism for short. Note that the last two equalities of Eq.(2.1) mean that $\psi$ is a $p^{h}$-linear map. In particular, if $R=S$, then $\psi$ is called a $\sigma_{p^{h}}$-algebra automorphism, or a $p^{h}$-automorphism for short. And, if $\psi:R\to S$ is a $p^{h}$-isomorphism, then for any ideal $I$ of $R$, the image $\psi(I)$ is an ideal of $S$ and $\dim_{F}\psi(I)=\dim_{F}I$. There are two typical examples as follows.
Example 2.1.
(1). For any polynomial $f(X)=\sum_{i}a_{i}X^{i}\in F[X]$, we denote $f^{(p^{h})}(X)=\sum_{i}a_{i}^{p^{h}}X^{i}$. The Galois automorphism $\sigma_{p^{h}}:F\to F$, $\alpha\mapsto\alpha^{p^{h}}$, induces the map
$\tilde{\sigma}_{p^{h}}\!:~{}F[X]\,\longrightarrow\,F[X],~{}~{}~{}~{}f(X)\,\longmapsto\,f^{(p^{h})}(X),$ (2.2)
which is a $p^{h}$-automorphism of $F[X]$. Let ${\rm ord}(\sigma_{p^{h}})$ denote the order of $\sigma_{p^{h}}$. It is easy to check that ${\rm ord}(\tilde{\sigma}_{p^{h}})={\rm ord}(\sigma_{p^{h}})=\frac{\ell}{\gcd(h,\ell)}$.
(2). For the matrix algebra ${\rm M}_{n}(F)=\big{\{}(a_{ij})_{n\times n}\,\big{|}\,a_{ij}\in F\big{\}}$ consisting of all $n\times n$ matrices over $F$, the map
$\dot{\sigma}_{p^{h}}\!:{\rm M}_{n}(F)\to{\rm M}_{n}(F),~{}~{}~{}~{}\big{(}a_{ij}\big{)}_{n\times n}\longmapsto\big{(}a_{ij}^{p^{h}}\big{)}_{n\times n},$ (2.3)
is a $p^{h}$-automorphism of ${\rm M}_{n}(F)$.
Let $\lambda\in F^{\times}$. For a positive integer $k$, we write $E_{k}$ to denote the identity matrix of degree $k$. Let $P_{\lambda}$ denote the $\lambda$-constacyclic permutation matrix of degree $n$ as follows
$P_{\lambda}=\begin{pmatrix}&E_{n-1}\\ \lambda E_{1}\end{pmatrix}=\begin{pmatrix}0&1\\ &0&\ddots\\ &&\ddots&\ddots\\ &&&\ddots&1\\ \lambda&&&&0\end{pmatrix}.$ (2.4)
In particular, if $\lambda=1$ then $P:=P_{1}=\begin{pmatrix}&E_{n-1}\\ E_{1}\end{pmatrix}$ is the cyclic permutation matrix. By matrix multiplication, for $\mathbf{a}=(a_{0},a_{1},\dots,a_{n-1})\in{F^{n}}$ we have that
$\mathbf{a}\cdot P_{\lambda}=(a_{0},a_{1},a_{2},\dots,a_{n-1})\cdot P_{\!\lambda}=(\lambda a_{n-1},a_{0},a_{1},\dots,a_{n-2}),$ (2.5)
which is the vector obtained by $\lambda$-constacyclically permuting the items of the vector $(a_{0},a_{1},\dots,a_{n-1})$; and that for $\mathbf{a}=(a_{0},a_{1},\dots,a_{n-1})\in{F^{n}}$ and $\mathbf{a^{\prime}}=(a^{\prime}_{0},a^{\prime}_{1},\dots,a^{\prime}_{n-1})\in{F^{n}}$,
$\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}\begin{pmatrix}P_{\lambda}\!\\ &\!P_{\lambda}\end{pmatrix}$ $=(\lambda a_{n-1},a_{0},a_{1},\dots,a_{n-2},\,\lambda a^{\prime}_{n-1},a^{\prime}_{0},a^{\prime}_{1},\dots,a^{\prime}_{n-2}).$ (2.6)
Thus, we get another description of the $\lambda$-constacyclic codes (cf. Eq.(1.1)) and the $2$-quasi $\lambda$-constacyclic codes (cf. Eq.(1.4)) as follows.
Lemma 2.2.
(1) Let $C$ be a subspace of $F^{n}$. Then $C$ is a $\lambda$-constacyclic code if and only if $\mathbf{c}\cdot P_{\lambda}\in C$, for any $\mathbf{c}\in C$.
(2) Let $C$ be a subspace of $F^{n}\times F^{n}$. Then $C$ is a $2$-quasi $\lambda$-constacyclic code if and only if $\Big{(}\mathbf{c},\,\mathbf{c}^{\prime}\big{)}\begin{pmatrix}P_{\lambda}\!\\ &\!P_{\lambda}\end{pmatrix}\in C$, for any $\big{(}\mathbf{c},\,\mathbf{c}^{\prime}\big{)}\in C$. ∎
In the following we always denote ${\cal R}_{\kern-1.0pt\lambda}=F[X]\big{/}\langle X^{n}-\lambda\rangle$, which is the quotient algebra of the polynomial algebra $F[X]$ over the ideal $\langle X^{n}-\lambda\rangle$ generated by $X^{n}-\lambda$. Any residue class modulo $X^{n}-\lambda$ has a unique representative polynomial with degree less than $n$. Hence we can write
${\cal R}_{\kern-1.0pt\lambda}=F[X]\big{/}\langle X^{n}-\lambda\rangle=\big{\{}\,a_{0}+a_{1}X+\dots+a_{n-1}X^{n-1}\;\big{|}\;a_{i}\in F\,\big{\}}.$ (2.7)
Further, the Cartesian product
${\cal R}_{\kern-1.0pt\lambda}^{2}={\cal R}_{\kern-1.0pt\lambda}\times{\cal R}_{\kern-1.0pt\lambda}=\big{\{}\big{(}a(X),a^{\prime}(X)\big{)}\;\big{|}\;a(X),~{}a^{\prime}(X)\in{\cal R}_{\kern-1.0pt\lambda}\big{\}}$ (2.8)
is an ${\cal R}_{\kern-1.0pt\lambda}$-module. For ${\cal R}_{\kern-1.0pt\lambda}$ and ${\cal R}_{\kern-1.0pt\lambda}^{2}$, the following identifications and results will be quoted later in this article.
Remark 2.3.
(1). There is a canonical linear isomorphism
$\iota:~{}~{}{{\cal R}_{\kern-1.0pt\lambda}}~{}\mathop{\longrightarrow}^{\cong}~{}F^{n},\qquad a(X)~{}\longmapsto~{}\mathbf{a},$ (2.9)
where $a(X)=a_{0}+a_{1}X+\dots+a_{n-1}X^{n-1}\in{\cal R}_{\kern-1.0pt\lambda}$ and $\mathbf{a}=(a_{0},a_{1},\dots,a_{n-1})\in F^{n}$. It is easy to check that
$\iota\big{(}Xa(X)\big{)}={\bf a}\cdot P_{\lambda},\qquad\forall\;a(X)\in{\cal R}_{\kern-1.0pt\lambda}.$ (2.10)
Then any element $a(X)$ of ${\cal R}_{\kern-1.0pt\lambda}$ is identified with the word $\mathbf{a}=(a_{0},a_{1},\dots,a_{n-1})$ of $F^{n}$; and by Lemma 2.2(1), the $\lambda$-constacyclic codes of length $n$ are identified with the ideals (${\cal R}_{\kern-1.0pt\lambda}$-submodules) of ${\cal R}_{\kern-1.0pt\lambda}$.
(2). For the ${\cal R}_{\kern-1.0pt\lambda}$-module ${\cal R}_{\kern-1.0pt\lambda}^{2}={\cal R}_{\kern-1.0pt\lambda}\times{\cal R}_{\kern-1.0pt\lambda}$, we have the following canonical linear isomorphism
$\iota^{(2)}:~{}~{}{{\cal R}_{\kern-1.0pt\lambda}}\times{{\cal R}_{\kern-1.0pt\lambda}}~{}\mathop{\longrightarrow}^{\cong}~{}F^{n}\times F^{n},\qquad\big{(}a(X),\,a^{\prime}(X)\big{)}~{}\longmapsto~{}\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)},$ (2.11)
where $a(X)=a_{0}+a_{1}X+\dots+a_{n-1}X^{n-1}$, $a^{\prime}(X)=a^{\prime}_{0}+a^{\prime}_{1}X+\dots+a^{\prime}_{n-1}X^{n-1}$, and $(\mathbf{a},\mathbf{a}^{\prime})=(a_{0},a_{1},\dots,a_{n-1},\,a^{\prime}_{0},a^{\prime}_{1},\dots,a^{\prime}_{n-1})$. For $\big{(}a(X),a^{\prime}(X)\big{)}\in{\cal R}_{\kern-1.0pt\lambda}^{2}$,
$\iota^{(2)}\big{(}X\big{(}a(X),a^{\prime}(X)\big{)}\big{)}=\iota^{(2)}\big{(}Xa(X),Xa^{\prime}(X)\big{)}=\big{(}{\bf a},{\bf a}^{\prime}\big{)}\begin{pmatrix}P_{\lambda}\!\\ &\!P_{\lambda}\end{pmatrix}.$ (2.12)
Then any element $\big{(}a(X),a^{\prime}(X)\big{)}\in{{{\cal R}_{\kern-1.0pt\lambda}}\times{{\cal R}_{\kern-1.0pt\lambda}}}$ is identified with the word $\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}\in{F^{n}\times F^{n}}$, and by Lemma 2.2(2) the $2$-quasi $\lambda$-constacyclic codes of length $2n$ are identified with the ${{\cal R}_{\kern-1.0pt\lambda}}$-submodules of ${\cal R}_{\kern-1.0pt\lambda}^{2}$.
Remark 2.4.
If $\gcd(n,q)=1$, then the algebra ${\cal R}_{\kern-1.0pt\lambda}$ is semisimple, cf. [4]; and by Ring Theory (e.g., cf. [19] or [17, Remark 2.4]), we have the following two.
(1) For any ideal (${\cal R}_{\kern-1.0pt\lambda}$-submodule) $C$ of ${{\cal R}_{\kern-1.0pt\lambda}}$, there is an idempotent $e_{C}$ of ${{\cal R}_{\kern-1.0pt\lambda}}$ such that $C={\cal R}_{\kern-1.0pt\lambda}e_{C}$ and ${{\cal R}_{\kern-1.0pt\lambda}}=C\oplus D$ where $D={{\cal R}_{\kern-1.0pt\lambda}}(1-e_{C})$. Note that $C={\cal R}_{\kern-1.0pt\lambda}e_{C}$ is an algebra with identity $e_{C}$ (but not a subalgebra of ${\cal R}_{\kern-1.0pt\lambda}$ in general because $e_{C}\neq 1$ in general); in particular, $C^{\times}$ makes sense. Moreover, if $I={\cal R}_{\kern-1.0pt\lambda}f$ with $f$ being an idempotent and an ideal $C\subseteq I$, then $I=C\oplus C^{\prime}$ with $C^{\prime}={\cal R}_{\kern-1.0pt\lambda}(f-e_{C})$.
(2) If $C$ and $C^{\prime}$ are ${\cal R}_{\kern-1.0pt\lambda}$-submodules of ${\cal R}_{\kern-1.0pt\lambda}$; and $\varphi:C\to C^{\prime}$ is an ${\cal R}_{\kern-1.0pt\lambda}$-module isomorphism, then $C^{\prime}=C$, and there is a $g\in C^{\times}$ such that $\varphi(c)=cg$ for any $c\in C$.
If $\lambda=1$, there is another identification. By Eq.(2.7), we denote
${\cal R}={\cal R}_{1}=F[X]/\langle X^{n}-1\rangle.$ (2.13)
Let $G=\langle\,x\,|\,x^{n}=1\,\rangle=\{1,x,\dots,x^{n-1}\}$ be the cyclic group of order $n$. Let $FG$ be the cyclic group algebra, i.e., $FG=\big{\{}\sum_{i=0}^{n-1}a_{i}x^{i}\,\big{|}\,a_{i}\in F\big{\}}$ is an $F$-vector space with basis $G$ and equipped with the multiplication induced by the multiplication of the group $G$ as follows:
$\Big{(}\sum_{i=0}^{n-1}a_{i}x^{i}\Big{)}\Big{(}\sum_{j=0}^{n-1}b_{j}x^{j}\Big{)}=\sum_{k=0}^{n-1}\Big{(}\sum_{x^{i}x^{j}=x^{k}}a_{i}b_{j}\Big{)}x^{k}.$
There is a canonical algebra isomorphism:
${\cal R}=F[X]/\langle X^{n}-1\rangle\;\mathop{\longrightarrow}^{\cong}\;FG,~{}~{}~{}~{}\sum_{i=0}^{n-1}a_{i}X^{i}\;\longmapsto\;\sum_{i=0}^{n-1}a_{i}x^{i}.$ (2.14)
Thus, ${\cal R}$ is identified with the cyclic group algebra $FG=\{\sum_{i=0}^{n-1}a_{i}x^{i}\,|\,a_{i}\in F\}$. And by Remark 2.3(1), the cyclic codes over $F$ of length $n$ are identified with the ideals of the cyclic group algebra $FG$. Similarly, $2$-quasi-cyclic codes over $F$ of length $2n$ are identified with the $FG$-submodules of the $FG$-module $(FG)^{2}=FG\times FG$.
With the identifications Eq.(2.14), we have more algebraic preliminaries about ${\cal R}$ to introduce. Assume that $\gcd(n,q)=1$, then ${\cal R}$ is semisimple. Let
$e_{0}=\frac{1}{n}\sum_{i=0}^{n-1}x^{i},~{}e_{1},~{}\dots,~{}e_{m}$ (2.15)
be all primitive idempotents of ${\cal R}$. Correspondingly, the irreducible decomposition of $X^{n}-1$ in $F[X]$ is as follows
$X^{n}-1=\varphi_{0}(X)\varphi_{1}(X)\dots\varphi_{m}(X),\quad\mbox{where }\varphi_{0}(X)=X-1,$ (2.16)
such that
$\begin{array}[]{l}{\cal R}={\cal R}e_{0}\oplus{\cal R}e_{1}\oplus\dots\oplus{\cal R}e_{m};\\[3.0pt] {\cal R}e_{i}\cong F[X]/\langle\varphi_{i}(X)\rangle,\quad i=0,1,\dots,m.\end{array}$ (2.19)
Since $\varphi_{i}(X)$ is irreducible over $F$, each ${\cal R}e_{i}$ is an extension field over $F$ with identity $e_{i}$, and $\dim_{F}{\cal R}e_{i}=\deg\varphi_{i}(X)$. As $e_{0}=\frac{1}{n}\sum_{i=0}^{n-1}x^{i}$, $xe_{0}=e_{0}$; so
${\cal R}e_{0}=\!\big{\{}\alpha\!+\!\alpha x+\dots+\!\alpha x^{n-1}\,\big{|}\,\alpha\!\in\!F\big{\}},~{}\dim_{F}{\cal R}e_{0}=1,~{}{\rm w}({\cal R}e_{0})=n.$ (2.20)
For $1\leq i\leq m$, we denote
$\mu(n):=\min_{1\leq i\leq m}\deg\varphi_{i}(X)=\min_{1\leq i\leq m}\dim_{F}{\cal R}e_{i}.$ (2.21)
Remark 2.5.
In general, in this paper we consider ${\cal R}_{\kern-1.0pt\lambda}$ for any $q$ and $n$ (not restricted to the semisimple case) unless the hypothesis “$\gcd(n,q)=1$” is explicitly assumed. Once “$\gcd(n,q)=1$” is assumed, the above preliminaries on the semisimple case can be quoted.
Remark 2.6.
[…]
$h_{q}(\delta)=\delta\log_{q}(q\!-\!1)-\delta\log_{q}(\delta)-(1\!-\!\delta)\log_{q}(1\!-\!\delta),~{}~{}~{}\delta\in[0,1\!-\!q^{-1}],$ (2.22)
which value strictly increases from 00 to $1$ while $\delta$ increases from 00 to $1-q^{-1}$. By [23, Theorem IV.17], for any real number $\delta$ with $0<\delta<1-q^{-1}$ and $h_{q}(\delta)<\frac{1}{4}$, there are self-dual $2$-quasi-cyclic codes $C_{1},C_{2},\dots$ over $F$ (hence ${\rm R}(C_{i})=\frac{1}{2}$) such that: (1) the relative minimum distance $\Delta(C_{i})>\delta$  for $i=1,2,\dots$; and (2) the code length $2n_{i}$ of $C_{i}$ satisfies that every $n_{i}$ is odd and coprime to $q$, and $\lim\limits_{i\to\infty}\frac{\log_{q}(n_{i})}{\mu(n_{i})}=0$ (in particular, $2n_{i}\to\infty$).
Definition 2.7.
Let $C\subseteq F^{n}=F^{I}$ with $I=\{0,1,\dots,n-1\}$ being the index set. If there are positive integers $s,k,t$ and subsets (repetition is allowed) $I_{1},\dots,I_{s}$ of $I$ with $|I_{j}|=k$ for $j=1,\dots,s$ satisfying the following two conditions: (1) for each $I_{j}=\{i_{1},\cdots,i_{k}\}$ ($1\leq j\leq s$) the projection $\rho_{I_{j}}$: $F^{I}~{}\to~{}F^{I_{j}}$, $(a_{0},a_{1},\dots,a_{n-1})\mapsto(a_{i_{1}},\dots,a_{i_{k}})$, maps $C$ bijectively onto $F^{I_{j}}$; and (2) for any $i$ ($0\leq i<n$) the number of the subsets $I_{j}$ which contains $i$ (i.e., $i\in I_{j}$) equals $t$; then we say that $C$ is a balanced code over $F$ of length $n$, and $I_{1},\dots,I_{s}$ are called information index sets of the code $C$.
An important result (see [13, Corollary 3.4]) is that: if $C\subseteq F^{n}$ is a balance code with cardinality $|C|=q^{k}$, then
$|C^{\leq\delta}|\leq q^{kh_{q}(\delta)},~{}~{}~{}\mbox{for}~{}~{}0\leq\delta\leq 1-q^{-1},$
where $h_{q}(\delta)$ is defined in Eq.(2.22) and
$C^{\leq\delta}=\{\,{\bf c}\,|\,{\bf c}\in C,\,{{\rm w}({\bf c})}\leq\delta n\,\}.$ (2.23)
Constacyclic codes are balanced, see [12, Lemma II.8]. By [13, Corollary 3.4 and Corollary 3.5], we can easily obtain the following lemma.
Lemma 2.8.
If $A$ is an ideal of ${\cal R}_{\kern-1.0pt\lambda}$, then the ${\cal R}_{\kern-1.0pt\lambda}$-submodule $A\times A$ of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ is a balanced code, hence for $0\leq\delta\leq 1-q^{-1}$,
$|(A\times A)^{\leq\delta}|\leq q^{2\dim_{F}(A)\cdot h_{q}(\delta)}.$
3 $2$-quasi constacyclic codes over finite fields
In this sections we are primarily concerned with the algebra properties of $2$-quasi $\lambda$-constacyclic codes over $F$. In the following, we always assume that
$\lambda\in F^{\times}~{}~{}~{}\mbox{and}~{}~{}~{}t={\rm ord}_{F^{\times}}(\lambda),$ (3.1)
where ${\rm ord}_{F^{\times}}(\lambda)$ denotes the order of $\lambda$ in the multiplication group $F^{\times}$.
As remarked in Remark 2.5, most of this section discusses ${\cal R}_{\kern-1.0pt\lambda}$ for any $q$ and $n$, only Theorem 3.9 and its corollaries consider the semisimple case (i.e., $\gcd(n,q)=1$).
Remark 3.1.
For $F^{n}\times F^{n}$, there are two projections $\rho_{1}$, $\rho_{2}$: $F^{n}\times F^{n}\to F^{n}$ as follows
$\rho_{1}({\bf a},{\bf a}^{\prime}\big{)}={\bf a},~{}~{}~{}\rho_{2}({\bf a},{\bf a}^{\prime})={\bf a}^{\prime},~{}~{}~{}~{}~{}~{}\forall~{}({\bf a},\,{\bf a}^{\prime})\in F^{n}\times F^{n}.$
By the linear isomorphism Eq.(2.11), the projections $\rho_{1}$ and $\rho_{2}$ are also defined on ${\cal R}_{\kern-1.0pt\lambda}^{2}={\cal R}_{\kern-1.0pt\lambda}\times{\cal R}_{\kern-1.0pt\lambda}$: for $\big{(}a(X),\,a^{\prime}(X)\big{)}\in{\cal R}_{\kern-1.0pt\lambda}^{2}$,
$\rho_{1}\big{(}a(X),\,a^{\prime}(X)\big{)}=a(X),~{}~{}~{}~{}~{}\rho_{2}\big{(}a(X),\,a^{\prime}(X)\big{)}=a^{\prime}(X).$
For any ${\cal R}_{\kern-1.0pt\lambda}$-submodule $C$ of ${{\cal R}_{\kern-1.0pt\lambda}}\times{{\cal R}_{\kern-1.0pt\lambda}}$, restricting $\rho_{1}$ to $C$, we have an ${{\cal R}_{\kern-1.0pt\lambda}}$-homomorphism $\rho_{1}|_{C}:C\to\rho_{1}(C)$. Observe that the kernel of the restricted homomorphism is
${\rm Ker}(\rho_{1}|_{C})=C\cap(0\times{{\cal R}_{\kern-1.0pt\lambda}})=\big{\{}\big{(}0,c^{\prime}(X)\big{)}\in C\,\big{|}\,c^{\prime}(X)\in{{{\cal R}_{\kern-1.0pt\lambda}}}\big{\}}.$ (3.2)
It is known that for $\lambda_{1}\neq\lambda_{2}\in F^{\times}$, if a linear code $C\subseteq F^{n}$ is both $\lambda_{1}$-constacyclic and $\lambda_{2}$-constacyclic, then either $C=\{{\bf 0}\}$ or $C=F^{n}$ (cf. [9]). Extending the result to $F^{n}\times F^{n}$, we get the following lemma.
Lemma 3.2.
Let $\lambda_{1},\lambda_{2}\in F^{\times}$ such that $\lambda_{1}\neq\lambda_{2}$. If a subspace $C$ of $F^{n}\times F^{n}$ is both a $2$-quasi $\lambda_{1}$-constacyclic code and a $2$-quasi $\lambda_{2}$-constacyclic code, then either $\rho_{1}(C)=\{{\bf 0}\}$ or $\rho_{1}(C)=F^{n}$; and it is the same for $\rho_{2}(C)$.
Proof.
Suppose $\rho_{1}(C)\neq\{{\bf 0}\}$. There is a codeword
$({\bf c},\,{\bf c}^{\prime})=(c_{0},\dots,c_{n-1},\;c^{\prime}_{0},\dots,c^{\prime}_{n-1})\,\in\,C$
with some $c_{j}\neq 0$, $0\leq j<n$; so we can assume that $c_{n-1}\neq 0$. By the double $\lambda$-constacyclic permutation in Eq.(1.4), then $C$ contains the following word
$\frac{1}{(\lambda_{1}-\lambda_{2})c_{n-1}}\Big{(}\big{(}\mathbf{c}P_{\!\lambda_{1}},\,\mathbf{c}^{\prime}P_{\!\lambda_{1}}\big{)}-\big{(}\mathbf{c}P_{\!\lambda_{2}},\,\mathbf{c}^{\prime}P_{\!\lambda_{2}}\big{)}\Big{)}=\big{(}1,0,\dots,0,\,d^{\prime}_{0},\dots,d^{\prime}_{n-1}\big{)},$
for some $d^{\prime}_{i}\in{F}$, $i=0,1,\dots,n-1$. Using the double $\lambda$-constacyclic permutation $i$ times, we see that $C$ contains such words
$\big{(}\mathop{0,\dots,0,\,1,\,0,\dots,0,}\limits_{\mbox{\scriptsize index\,$i$}}\,b^{\prime}_{0},\dots,b^{\prime}_{n-1}\big{)},\qquad i=0,1,\dots,n-1,$
for some $b^{\prime}_{i}\in{F}$, $i=0,1,\dots,n-1$. It follows that $\rho_{1}(C)$ contains a basis of $F^{n}$, hence $\rho_{1}(C)=F^{n}$.
By the same argument, either $\rho_{2}(C)=\{{\bf 0}\}$ or $\rho_{2}(C)=F^{n}$. ∎
For any matrix $A=\big{(}a_{ij}\big{)}_{m\times n}$ with $a_{ij}\in F$, we denote $A^{(p^{h})}=\big{(}a_{ij}^{p^{h}}\big{)}_{m\times n}$ (cf. Example 2.1(2)). By $A^{T}$ we denote the transpose of $A$. Let
$A^{*h}=(A^{(p^{h})})^{T}=\big{(}a_{ji}^{p^{h}}\big{)}_{n\times m}=\begin{pmatrix}a_{11}^{p^{h}}&\cdots&a_{m1}^{p^{h}}\\ \dots&\dots&\dots\\ a_{1n}^{p^{h}}&\cdots&a_{mn}^{p^{h}}\end{pmatrix}$ (3.3)
be the transpose of the matrix $A^{(p^{h})}$. We call $A^{*h}$ the Galois $p^{h}$-transpose of $A$. If $h=0$, $A^{*0}=A^{T}$ is just the transpose matrix of $A$. If $\ell$ is even and $h=\frac{\ell}{2}$, $A^{*\frac{\ell}{2}}$ is the Hermitian transpose of $A$.
For $B=(b_{ij})_{m\times n}$, $C=(c_{ij})_{n\times k}$ and $\alpha\in F$, it is easy to check that
$(AC)^{*h}=C^{*h}A^{*h},~{}~{}(A+B)^{*h}=A^{*h}+B^{*h},~{}~{}(\alpha A)^{*h}=\alpha^{p^{h}}A^{*h}.$ (3.4)
[…]
With the identification Eq.(2.9) and the operator “$*h$”, we can compute the Galois $p^{h}$-inner product on ${{\cal R}_{\kern-1.0pt\lambda}}$ (see Eq.(1.5)) in a matrix version:
$\big{\langle}a,\,b\big{\rangle}_{h}=\mathbf{a}\cdot\mathbf{b}^{*h},\qquad\forall~{}a,b\in{{\cal R}_{\kern-1.0pt\lambda}}.$ (3.5)
And for any $\lambda$-constacyclic code $C$ (i.e., any idea of ${\cal R}_{\kern-1.0pt\lambda}$), the Galois $p^{h}$-dual code of $C$ is as follows:
$C^{\bot h}=\big{\{}\,a\in{{{\cal R}_{\kern-1.0pt\lambda}}}~{}\big{|}~{}\big{\langle}c,\,a\big{\rangle}_{h}=0,~{}\forall\,c\in{C}\,\big{\}}.$
Similarly, with the identification Eq.(2.11) the Galois $p^{h}$-inner product on ${\cal R}_{\kern-1.0pt\lambda}^{2}$ is computed in a matrix version:
$\big{\langle}(a,a^{\prime}),\,(b,b^{\prime})\big{\rangle}_{h}=\big{(}\mathbf{a},\mathbf{a}^{\prime}\big{)}\cdot\big{(}\mathbf{b},\mathbf{b}^{\prime}\big{)}^{*h},\qquad\forall~{}\big{(}a,a^{\prime}\big{)},\,\big{(}b,b^{\prime}\big{)}\in{\cal R}_{\kern-1.0pt\lambda}^{2}.$ (3.6)
And for any $2$-quasi $\lambda$-constacyclic code $C$ (${\cal R}_{\kern-1.0pt\lambda}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$), the Galois $p^{h}$-dual code of $C$ is
$C^{\bot h}=\big{\{}\,(a,a^{\prime})\in{{\cal R}_{\kern-1.0pt\lambda}^{2}}~{}\big{|}~{}\big{\langle}(c,c^{\prime}),\,(a,a^{\prime})\big{\rangle}_{h}=0,~{}\forall\,(c,c^{\prime})\in{C}\,\big{\}}.$
If $C=C^{\bot h}$, then we say that $C$ is Galois $p^{h}$-self-dual, or Galois self-dual.
Lemma 3.3.
If $C\subseteq F^{n}\times F^{n}$ is a $2$-quasi $\lambda$-constacyclic code, then $C^{\bot h}$ is a $2$-quasi $\lambda^{\!-\!p^{\ell-h}}\!$-constacyclic code. In particular, if $h=0$, $C^{\bot 0}=C^{\bot}$ is a $2$-quasi $\lambda^{\!-\!1}$-constacyclic code.
Proof.
Assume that $\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}\in C^{\bot h}$. By Lemma 2.2(2), it is enough to prove that for any $\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)}\in C$ we have
$\left\langle\big{(}\mathbf{c},\,\mathbf{c}\big{)},\,\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}\begin{pmatrix}P_{\lambda^{\!-\!p^{\ell-h}}}\!\\ &\!P_{\lambda^{\!-\!p^{\ell-h}}}\end{pmatrix}\right\rangle_{\!h}=0.$
Applying Eq.(3.4) and Eq.(3.6) yields
$\left\langle\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)},\,\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}\begin{pmatrix}P_{{\lambda^{\!-\!p^{\ell-h}}}}\!\\ &\!P_{{\lambda^{\!-\!p^{\ell-h}}}}\end{pmatrix}\right\rangle_{\!h}$
$=\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)}\cdot\left(\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}\begin{pmatrix}P_{{\lambda^{\!-\!p^{\ell-h}}}}\!\\ &\!P_{{\lambda^{\!-\!p^{\ell-h}}}}\end{pmatrix}\right)^{*h}$
$=\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)}\begin{pmatrix}P_{\lambda^{\!-\!p^{\ell-h}}}\!\\ &\!P_{\lambda^{\!-\!p^{\ell-h}}}\end{pmatrix}^{*h}\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}^{*h}$
$=\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)}\begin{pmatrix}\big{(}P_{\lambda^{\!-\!p^{\ell-h}}}\big{)}^{*h}\!\\ &\!\big{(}P_{\lambda^{\!-\!p^{\ell-h}}}\big{)}^{*h}\end{pmatrix}\big{(}\mathbf{a},\,\mathbf{a}^{\prime}\big{)}^{*h}.$
Because $(\lambda^{\!-\!p^{\ell-h}})^{p^{h}}=(\lambda^{p^{\ell}})^{-1}=\lambda^{-1}$, it is easy to check that
$P_{\!\lambda}\big{(}P_{\lambda^{\!-\!p^{\ell-h}}}\big{)}^{*h}=\begin{pmatrix}&E_{n-1}\\ \lambda E_{1}\end{pmatrix}\begin{pmatrix}&\!(\lambda^{\!-\!p^{\ell-h}})^{p^{h}}E_{1}\\ E_{n-1}\!\end{pmatrix}=E_{n}.$
Since $P_{\!\lambda}^{n}=\lambda E_{n}$ and ${\rm ord}_{F^{\times}}(\lambda)=t$ (see Eq(3.1)), we have $P_{\!\lambda}^{nt}=E_{n}$, hence $P_{\!\lambda}^{-1}=P_{\!\lambda}^{nt-1}$. So
$\big{(}P_{\lambda^{\!-\!p^{\ell-h}}}\big{)}^{*h}=P_{\!\lambda}^{-1}=P_{\!\lambda}^{nt-1}.$ (3.7)
By Lemma 2.2(2), $\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)}\begin{pmatrix}P_{\!\lambda}\!\\ &\!P_{\!\lambda}\end{pmatrix}^{nt-1}\in C$. We get that
$\bigg{\langle}\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)},\,\big{(}\mathbf{a},\mathbf{a}^{\prime}\big{)}\begin{pmatrix}P_{\!\lambda^{\!-\!p^{\ell-h}}}\!\!\\ &\!\!P_{\!\lambda^{\!-\!p^{\ell-h}}}\end{pmatrix}\bigg{\rangle}_{\!h}=\bigg{\langle}\big{(}\mathbf{c},\mathbf{c}^{\prime}\big{)}\begin{pmatrix}P_{\!\lambda}\!\\ &\!P_{\!\lambda}\end{pmatrix}^{nt-1}\!,\,\big{(}\mathbf{a},\mathbf{a}^{\prime}\big{)}\bigg{\rangle}_{\!h}=0.$
We are done. ∎
Remark 3.4.
We note that $\lambda=\lambda^{-p^{\ell-h}}$ if and only if $\lambda^{1+p^{h}}=1$, since
$\lambda=\lambda^{-p^{\ell-h}}\;\iff\;\lambda^{p^{h}}=(\lambda^{-p^{\ell-h}})^{p^{h}}=\lambda^{-p^{\ell}}=\lambda^{-1}\;\iff\;\lambda^{1+p^{h}}=1.$
If $\lambda=\lambda^{-p^{\ell-h}}$, then for any $2$-quasi $\lambda$-constacyclic code $C$, its Galois $p^{h}$-dual code $C^{\bot h}$ is still a $2$-quasi $\lambda$-constacyclic code. Otherwise (i.e., $\lambda\neq\lambda^{-p^{\ell-h}})$, Lemma 3.2 implies that, for many $2$-quasi $\lambda$-constacyclic codes, their Galois $p^{h}$-dual codes are no longer $2$-quasi $\lambda$-constacyclic codes.
Remark 3.5.
For $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}\in{{\cal R}_{\kern-1.0pt\lambda}}$, by $C_{a}$ we denote the ideal of ${{\cal R}_{\kern-1.0pt\lambda}}$ generated by $a(X)$. Similarly, for $(a(X),a^{\prime}(X))\in{\cal R}_{\kern-1.0pt\lambda}^{2}$, by $C_{a,a^{\prime}}$ we denote the ${{\cal R}_{\kern-1.0pt\lambda}}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ generated by $\big{(}a(X),\,a^{\prime}(X)\big{)}$. Note that any ideal of ${\cal R}_{\kern-1.0pt\lambda}$ is generated by one element (cf. Remark 2.4(1) for semisimple case, and cf. [11, Lemma 4.3] for general case). However, some ${\cal R}_{\kern-1.0pt\lambda}$-submodules of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ can not be generated by one element. For example, as an ${\cal R}_{\kern-1.0pt\lambda}$-submodule ${\cal R}_{\kern-1.0pt\lambda}^{2}$ can not be generated by one element (because any ${\cal R}_{\kern-1.0pt\lambda}$-module generated by one element is a quotient of the regular module).
Definition 3.6.
For any $a(X)=a_{0}+a_{1}X+\cdots+a_{n-1}X^{n-1}\in{F[X]}$ with $\deg(a(X))<n$, we have an $n\times n$ matrix
$a(P_{\!\lambda})=a_{0}E+a_{1}P_{\!\lambda}+\cdots+a_{n-1}P_{\!\lambda}^{n-1}=\begin{pmatrix}a_{0}&a_{1}&\cdots&a_{n-1}\\ \lambda a_{n-1}&a_{0}&\cdots&a_{n-2}\\ \cdots&\cdots&\cdots&\cdots\\ \lambda a_{1}&\lambda a_{2}&\cdots&a_{0}\end{pmatrix},$ (3.8)
whose first row is the vector $\mathbf{a}=(a_{0},a_{1},\dots,a_{n-1})$, and each next row is obtained by $\lambda$-constacyclically permuting the present row (cf. Eq.(2.5)). We call $a(P_{\!\lambda})$ the $\lambda$-consta circulant matrix associated with the polynomial $a(X)$.
Lemma 3.7.
Let $a(X),a^{\prime}(X)\in{\cal R}_{\kern-1.0pt\lambda}$. Then we have:
(1)  $C_{a}$ is linearly generated by the rows of the $n\times n$ matrix $a(P_{\!\lambda})$.
(2)  $C_{a,a^{\prime}}$ is linearly generated by the rows of the $n\times 2n$ matrix $\big{(}a(P_{\!\lambda}),a^{\prime}(P_{\!\lambda})\big{)}$.
Proof.
(1). Let $C_{a}$ be the ideal of ${\cal R}_{\kern-1.0pt\lambda}$ generated by $a(X)$, i.e.,
$C_{a}=\Big{\{}f(X)a(X)~{}\Big{|}~{}f(X)=\sum_{i=0}^{n-1}f_{i}X^{i}\in{\cal R}_{\kern-1.0pt\lambda},f_{i}\in F\Big{\}}.$
Obviously, $f(X)a(X)=f_{0}\cdot a(X)+f_{1}\cdot Xa(X)+\dots+f_{n-1}\cdot X^{n-1}a(X)$. So $C_{a}$ is the subspace of ${\cal R}_{\kern-1.0pt\lambda}$ linearly generated by $a(X),Xa(X),\dots,X^{n-1}a(X)$. Eq.(2.10) and Eq.(3.8) imply that $a(X)$ is identified with the row vector ${\bf a}$ which is just the first row of the matrix $a(P_{\lambda})$; for $i=1,\dots,n-1$, $X^{i}a(X)$ is identified with the row vector ${\bf a}\cdot P_{\lambda}^{\,i}$, which is just the $i$’th row of the matrix $a(P_{\lambda})$. Therefore, $C_{a}$ is linearly generated by the rows of the matrix $a(P_{\lambda})$.
Obviously, (2) is proved in a similar way. ∎
Corollary 3.8.
For $a(X)\in{{\cal R}_{\kern-1.0pt\lambda}}$, the $2$-quasi $\lambda$-constacyclic code $C_{1,a}$ has a generating matrix $\big{(}E_{n},\,a(P_{\!\lambda})\big{)}$.
Proof.
By Lemma 3.7, the rows of the matrix $\big{(}E_{n},\,a(P_{\!\lambda})\big{)}$ linearly generate the code $C_{1,a}$. And the rows of the matrix $\big{(}E_{n},\,a(P_{\!\lambda})\big{)}$ are linearly independent. ∎
In the rest of this section, we turn to the semisimple case, i.e., $\gcd(n,q)=1$. Extending [17, Theorem 3.2] which characterized the algebraic structure of $2$-quasi-cyclic codes in the semisimple case, we characterize the algebraic structure of $2$-quasi constacyclic codes as follows.
Theorem 3.9.
Assume that $\gcd(n,q)=1$. If $C$ is an ${\cal R}_{\kern-1.0pt\lambda}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$, then there are ideals $C_{1},C_{2},C_{12}$ of ${{\cal R}_{\kern-1.0pt\lambda}}$ satisfying that $C_{1}\cap C_{12}=C_{2}\cap C_{12}=0$ and an element $g\in C_{12}^{\times}$ such that
$C=(C_{1}\times C_{2})\oplus\widehat{C}_{12},~{}~{}\mbox{where}~{}~{}\widehat{C}_{12}=\big{\{}\big{(}c_{12},c_{12}g\big{)}\,\big{|}\,c_{12}\in C_{12}\big{\}}\cong C_{12}.$ (3.9)
Conversely, if there are ideals $C_{1},C_{2},C_{12}$ of ${{\cal R}_{\kern-1.0pt\lambda}}$ with $C_{1}\cap C_{12}=C_{2}\cap C_{12}=0$ and an element $g\in C_{12}^{\times}$, then $C$ in Eq.(3.9) is an ${\cal R}_{\kern-1.0pt\lambda}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$.
Proof.
The “conversely” part is obviously true because both $C_{1}\times C_{2}$ and $\widehat{C}_{12}$ are ${\cal R}_{\kern-1.0pt\lambda}$-submodules of ${\cal R}_{\kern-1.0pt\lambda}^{2}$, and $(C_{1}\times C_{2})\cap\widehat{C}_{12}=0$.
Assume that $C$ is an ${\cal R}_{\kern-1.0pt\lambda}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$, and $\rho_{1}$, $\rho_{2}$ are defined in Remark 3.1. We consult the module version of Goursat Lemma (see[17, Remark 3.1]). Take
$\begin{array}[]{l}\tilde{C_{1}}=\rho_{1}(C)=\big{\{}a_{1}\in{{{\cal R}_{\kern-1.0pt\lambda}}}~{}|~{}(a_{1},a_{2})\in{C}~{}\mbox{for some}~{}a_{2}\in{{{\cal R}_{\kern-1.0pt\lambda}}}\big{\}},\\[2.0pt] {C_{1}}=\rho_{1}(C\cap({{\cal R}_{\kern-1.0pt\lambda}}\times 0))=\big{\{}a_{1}\in{{{\cal R}_{\kern-1.0pt\lambda}}}~{}|~{}(a_{1},0)\in{C}\big{\}},\\[2.0pt] \tilde{C_{2}}=\rho_{2}(C)=\big{\{}a_{2}\in{{{\cal R}_{\kern-1.0pt\lambda}}}~{}|~{}(a_{1},a_{2})\in{C}~{}\mbox{for some}~{}a_{1}\in{{{\cal R}_{\kern-1.0pt\lambda}}}\big{\}},\\[2.0pt] {C_{2}}=\rho_{2}(C\cap(0\times{{\cal R}_{\kern-1.0pt\lambda}}))=\big{\{}a_{2}\in{{{\cal R}_{\kern-1.0pt\lambda}}}~{}|~{}(0,a_{2})\in{C}\big{\}}.\end{array}$ (3.14)
Then $C_{i}\subseteq\tilde{C_{i}}$ are ideals of ${{\cal R}_{\kern-1.0pt\lambda}}$, $i=1,2$. For any $c_{1}+C_{1}\in{\tilde{C_{1}}/C_{1}}$, there is a unique $c_{2}+C_{2}\in{\tilde{C_{2}}/C_{2}}$ such that $(c_{1},c_{2})\in{C}$, hence we have the map
$\varphi:\tilde{C}_{1}/C_{1}\to\tilde{C}_{2}/C_{2},~{}c_{1}+C_{1}\mapsto c_{2}+C_{2},$ (3.15)
which is an ${{\cal R}_{\kern-1.0pt\lambda}}$-isomorphism, and
$C=\big{\{}(c_{1},c_{2})\;\big{|}\;c_{i}\in\tilde{C}_{i}~{}{\rm for}~{}i=1,2;~{}\varphi(c_{1}+C_{1})=c_{2}+C_{2}\big{\}}.$ (3.16)
Since ${{\cal R}_{\kern-1.0pt\lambda}}$ is semisimple and ${C}_{1}\subseteq\tilde{C}_{1}$ are ideals of ${{\cal R}_{\kern-1.0pt\lambda}}$, there is an ideal $C_{12}$ of ${\cal R}_{\kern-1.0pt\lambda}$ such that $\tilde{C}_{1}=C_{1}\oplus C_{12}$; see Remark 2.4(1). Similarly, we have an ideal $C_{12}^{\prime}$ of ${{\cal R}_{\kern-1.0pt\lambda}}$ such that $\tilde{C}_{2}=C_{2}\oplus C_{12}^{\prime}$. Then $\tilde{C}_{1}/C_{1}\cong C_{12}$ and $\tilde{C}_{2}/C_{2}\cong C^{\prime}_{12}$. The ${{\cal R}_{\kern-1.0pt\lambda}}$-isomorphism $\varphi$ in Eq.(3.15) induces an ${{\cal R}_{\kern-1.0pt\lambda}}$-isomorphism $\varphi^{\prime}:C_{12}\to C_{12}^{\prime}$ such that $\varphi\big{(}c+C_{1}\big{)}=\varphi^{\prime}(c)+C_{2}$ for all $c\in C_{12}$. Thus the image $C_{12}^{\prime}=C_{12}$, and there is a $g\in C_{12}^{\times}$ such that $\varphi^{\prime}(c)=cg$ for all $c\in C_{12}$; cf. Remark 2.4(2). In conclusion, we have an ideal $C_{12}$ of ${\cal R}_{\kern-1.0pt\lambda}$ such that $\tilde{C}_{1}=C_{1}\oplus C_{12},~{}\tilde{C}_{2}=C_{2}\oplus C_{12};$ and a $g\in C_{12}^{\times}$ such that
$C=\big{\{}\,(c_{1}+c_{12},\;c_{2}+c_{12}g)\;\big{|}\;c_{1}\in C_{1},\,c_{2}\in C_{2},\,c_{12}\in C_{12}\,\big{\}}.$
Obviously, $(c_{1}\!+\!c_{12},\;c_{2}\!+\!c_{12}g)=(c_{1},c_{2})+(c_{12},c_{12}g)$. Thus Eq.(3.9) holds. ∎
Lemma 3.10.
Assume that $\gcd(n,q)=1$. Then for $a,b\in{{\cal R}_{\kern-1.0pt\lambda}}$, $C_{a}\cap C_{b}=0$ if and only if $ab=0$.
Proof.
If $C_{a}\cap C_{b}=0$ then $ab\in C_{a}\cap C_{b}=0$. Conversely, assume that $ab=0$. Note that $C_{b}={\cal R}_{\kern-1.0pt\lambda}\,e_{b}$ for an idempotent $e_{b}$ which is the identity of the ring $C_{b}$ (cf. Remark 2.4(1)). There is an element $f\in{\cal R}_{\kern-1.0pt\lambda}$ such that $e_{b}=fb$. For any element $c\in{C_{a}\cap C_{b}}$, we can write $c=ra=sb$ with $r,\,s\in{{\cal R}_{\kern-1.0pt\lambda}}$. Then $c=ra=rae_{b}=rafb=0$. Thus $C_{a}\cap C_{b}=0$. ∎
Corollary 3.11.
Keep the notation in Theorem 3.9 (in particular, ${\gcd(n,q)=1}$). Any ${\cal R}_{\kern-1.0pt\lambda}$-submodule $C$ of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ can be written as
$C=(C_{a}\times C_{a^{\prime}})\oplus C_{b,bg}=C_{a,0}\oplus C_{0,a^{\prime}}\oplus C_{b,bg},$ (3.17)
where $a,a^{\prime},b\in{\cal R}_{\kern-1.0pt\lambda}$ and $g\in C_{b}^{\times}$ satisfy that $ab=a^{\prime}b=0$.
Proof.
Take $C_{1}=C_{a}$, $C_{2}=C_{a^{\prime}}$ and $C_{12}=C_{b}$, hence $\widehat{C}_{12}=C_{b,bg}$. By the above Lemma, we get Eq.(3.17) immediately. ∎
4 Galois self-dual $2$-quasi constacyclic codes
In this section we investigate the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$. By Lemma 3.3 and Remark 3.4, any Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic code (i.e., $C=C^{\bot h}$) is also $\lambda^{-p^{\ell-h}}$-constacyclic. We study them in two cases: $\lambda\neq\lambda^{-p^{\ell-h}}$ (i.e., $\lambda^{1+p^{h}}\neq 1$), or $\lambda=\lambda^{-p^{\ell-h}}$ (i.e., $\lambda^{1+p^{h}}=1$).
Still, this section discusses ${\cal R}_{\kern-1.0pt\lambda}$ for any $q$ and $n$ except for Theorem 4.10 which considers the semisimple case (i.e., $\gcd(n,q)=1$).
4.1 The case that $\lambda^{1+p^{h}}\neq 1$
Our concern in this subsection is the Galois $p^{h}$-self-dual 2-quasi $\lambda-$constacyclic codes over $F$ under the assumption that $\lambda^{1+p^{h}}\neq 1$.
Theorem 4.1.
Assume that $\lambda^{1+p^{h}}\neq 1$ (i.e., $\lambda\neq\lambda^{-p^{\ell-h}}$). The following three are equivalent to each other:
(1) $C$ is a Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic code over $F$ of length $2n$.
(2) $C=C_{1,\alpha}$ is a ${{\cal R}_{\kern-1.0pt\lambda}}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ generated by $(1,\alpha)$, where $\alpha\in F^{\times}$ with $\alpha^{1+p^{h}}=-1$, and $1,\alpha$ are viewed as the constant polynomials of ${{\cal R}_{\kern-1.0pt\lambda}}$.
(3) $C$ is an $F$-linear code of length $2n$ with a generating matrix $\big{(}E_{n},\alpha E_{n}\big{)}$, where $\alpha\in F^{\times}$ with $\alpha^{1+p^{h}}=-1$.
Proof.
Observe that by Corollary 3.8, the statements (2) and (3) are equivalent. Suppose that (3) holds. Then $\big{(}E_{n},\,\alpha E_{n}\big{)}\big{(}E_{n},\,\alpha E_{n}\big{)}^{*h}=(1+\alpha^{p^{h}+1})E_{n}=0,$ which implies $\langle C,C\rangle_{h}=0$, cf. Eq.(3.6). Since the rank of the generating matrix $\big{(}E_{n},\,\alpha E_{n}\big{)}$ is $n$, we have $\dim_{F}C=n$, and so the statement (1) follows. Therefore, it suffices to show that (1) implies (3).
Assume that (1) holds, i.e., $C=C^{\bot h}$. By Lemma 3.3, $C$ is both a $2$-quasi $\lambda$-constacyclic code and a $2$-quasi $\lambda^{\!-\!p^{\ell-h}}$-constacyclic code. Since $\lambda\neq\lambda^{-p^{\ell-h}}$, we deduce from Lemma 3.2 that either $\rho_{1}(C)=0$ or $\rho_{1}(C)={{\cal R}_{\kern-1.0pt\lambda}}$. Suppose that $\rho_{1}(C)=0$. Since $\dim_{F}C=n$, by Eq.(3.2), we have that $C=0\times{{\cal R}_{\kern-1.0pt\lambda}}$, which is impossible because $0\times{{\cal R}_{\kern-1.0pt\lambda}}$ is not Galois $p^{h}$-self-dual. So, it must be the case that $\rho_{1}(C)={{\cal R}_{\kern-1.0pt\lambda}}$. By the same argument, we have $\rho_{2}(C)={{\cal R}_{\kern-1.0pt\lambda}}$.
As $\rho_{1}(C)={{\cal R}_{\kern-1.0pt\lambda}}$, we can take $(1,a)\in C$ with $a=a(X){=\sum_{i=0}^{n-1}a_{i}X^{i}\in{{{\cal R}_{\kern-1.0pt\lambda}}}}$. Let $C_{1,a}$ be the ${{\cal R}_{\kern-1.0pt\lambda}}-$submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ generated by $(1,a)$. Obviously, $C_{1,a}\subseteq C$. By Corollary 3.8, $C_{1,a}$ has the generating matrix
$\big{(}E_{n},\;a(P_{\lambda})\big{)}=\begin{pmatrix}1&&&&&a_{0}&a_{1}&\cdots&a_{n-1}\\ &1&&&&\lambda a_{n-1}&a_{0}&\cdots&a_{n-2}\\ &&\ddots&&&\cdots&\cdots&\cdots&\cdots\\ &&&1&&\lambda a_{1}&\lambda a_{2}&\cdots&a_{0}\end{pmatrix}_{n\times 2n}.$ (4.1)
The rank of the matrix is $n$, hence $\dim_{F}C_{1,a}=n$. In a word, $C=C_{1,a}$ is linearly generated by the $n$ rows of the matrix in Eq.(4.1).
Because $C$ is also a $2$-quasi $\lambda^{\!-\!p^{\ell-h}}\!$-constacyclic code, by the same way we can also get that $C$ is linearly generated by the $n$ rows of the matrix $\big{(}E_{n},a(P_{\!\lambda^{-p^{\ell-h}}})\big{)}$. Thus any row of $\big{(}E_{n},a(P_{\!\lambda^{-p^{\ell-h}}})\big{)}$ is a linear combination of the rows of the matrix in Eq.(4.1). So there is an $n\times n$ matrix $A$ such that
$\big{(}E_{n},\;a(P_{\!\lambda^{-p^{\ell-h}}})\big{)}=A\Big{(}E_{n},\;a(P_{\lambda})\Big{)}=\Big{(}AE_{n},\;A\cdot a(P_{\lambda})\Big{)}.$
It follows that $A=E_{n}$, and so $a(P_{\!\lambda^{\!-\!p^{\ell-h}}})=A\cdot a(P_{\!\lambda})=a(P_{\!\lambda})$. The latter equality is as follows:
$\begin{pmatrix}a_{0}&a_{1}&\cdots&a_{n-1}\\ \lambda^{\!-\!p^{\ell-h}}a_{n-1}&a_{0}&\cdots&a_{n-2}\\ \cdots&\cdots&\cdots&\cdots\\ \lambda^{\!-\!p^{\ell-h}}a_{1}&\lambda^{\!-\!p^{\ell-h}}a_{2}&\cdots&a_{0}\end{pmatrix}=\begin{pmatrix}a_{0}&a_{1}&\cdots&a_{n-1}\\ \lambda a_{n-1}&a_{0}&\cdots&a_{n-2}\\ \cdots&\cdots&\cdots&\cdots\\ \lambda a_{1}&\lambda a_{2}&\cdots&a_{0}\end{pmatrix}.$
Since $\lambda^{-p^{\ell-h}}\neq\lambda$, it follows that $a_{i}=0$ for $i=1,\dots,n-1$; hence the polynomial $a(X)$ is a constant polynomial: $a(X)=\alpha$ for some $\alpha\in{F}$, and $C$ has a generating matrix $\big{(}E_{n},\,\alpha E_{n}\big{)}$. Because $C$ is Galois $p^{h}$-self-dual,
$0=\big{(}E_{n},\,\alpha E_{n}\big{)}\cdot\big{(}E_{n},\,\alpha E_{n}\big{)}^{*h}=E_{n}+\alpha^{1+\alpha^{p^{h}}}E_{n}=(1+\alpha^{1+p^{h}})E_{n},$
hence $01+\alpha^{1+p^{h}}=$ 0. In conclusion, (3) holds. ∎
Corollary 4.2.
Assume that $\lambda^{1+p^{h}}\neq 1$. The Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes of length $2n$ exist if and only if the polynomial $X^{1+p^{h}}+1$ has roots in $F$; and in that case, the ${{\cal R}_{\kern-1.0pt\lambda}}$-submodules $C_{1,\alpha}$ of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ with $\alpha\in F$ being a root of $X^{1+p^{h}}+1$ are all the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes of length $2n$. ∎
Note that in the Euclidean case “$h=0$”, $\lambda^{1+p^{h}}\neq 1$ if and only if $\lambda\neq\pm 1$.
Corollary 4.3.
[…]
Proof.
[…]
As a comparison, the Hermitian self-dual ones always exist.
Corollary 4.4.
Assume that $\ell$ is even and $\lambda^{1+p^{\ell/2}}\neq 1$. Then the Hermitian self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ of length $2n$ always exist; and in that case, the ${{\cal R}_{\kern-1.0pt\lambda}}$-submodules $C_{1,\alpha}$ of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ with $\alpha\in F$ satisfying $\alpha^{1+p^{\ell/2}}=-1$ are the all Hermitian self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ of length $2n$.
Proof.
If $q=p^{\ell}$ is even, then the polynomial $X^{1+p^{\ell/2}}+1=X^{1+p^{\ell/2}}-1$ always has roots in $F$. Assume that $p$ is odd. The order of the multiplication group
$|F^{\times}|=p^{\ell}-1=(p^{\ell/2}+1)(p^{\ell/2}-1),$
and $2\,\big{|}\,(p^{\ell/2}-1)$. So, there is a subgroup $H$ of the multiplication group $F^{\times}$ with order $|H|=2(p^{\ell/2}+1)$. Then any generator of the group $H$ is a root of the polynomial $X^{1+p^{\ell/2}}+1$. ∎
As a consequence, we get the following.
Theorem 4.5.
Assume that $\lambda^{1+p^{h}}\neq 1$ (i.e., $\lambda\neq\lambda^{-p^{\ell-h}}$). Then the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ are asymptotically bad.
Proof.
If the polynomial $X^{1+p^{h}}+1$ has no root in $F$, then Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ do not exist, hence Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ are asymptotically bad.
Otherwise, any Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic code $C_{1,\alpha}$ has minimum weight ${\rm w}(C_{1,\alpha})=2$, because any row of the generating matrix $\big{(}E_{n},\alpha E_{n}\big{)}$ has weight $2$. The relative minimum distance $\Delta(C_{1,\alpha})=\frac{2}{2n}\to 0$ while ${n\to\infty}$. So Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ are asymptotically bad. ∎
4.2 The case that $\lambda^{1+p^{h}}=1$
We start with a general result about the $\lambda$-consta circulant matrices (cf. Definition 3.6). As in Example 2.1(2), ${\rm M}_{n}(F)$ denotes the $n\times n$ matrix algebra over $F$. We consider its subset consisting of all the $\lambda$-consta circulant matrices of degree $n$:
${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)=\big{\{}\,a(P_{\lambda})\;\big{|}\;a(X)\in{\cal R}_{\kern-1.0pt\lambda}\,\big{\}},$
where $P_{\lambda}$ is defined in Eq.(2.4) and $a(P_{\lambda})$ is defined in Eq.(3.8).
Lemma 4.6.
With the notation as above. ${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$ is a subalgebra of ${\rm M}_{n}(F)$ and the following is an algebra isomorphism:
$\nu:\;{{\cal R}_{\kern-1.0pt\lambda}}\;\mathop{\longrightarrow}\;{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F),\quad a(X)\,\longmapsto\,a(P_{\!\lambda}).$ (4.2)
Proof.
The following is obviously an $F$-algebra homomorphism:
$\delta:~{}~{}F[X]\to{\rm M}_{n}(F),~{}~{}~{}f(X)\mapsto f(P_{\!\lambda}).$
Since $P_{\!\lambda}^{n}=\lambda E_{n}$, the kernel of the homomorphism $\delta$ is the ideal $\langle X^{n}-\lambda\rangle$ of $F[X]$ generated by $X^{n}-\lambda$. Note that the quotient algebra $F[X]/\langle X^{n}-\lambda\rangle={{\cal R}_{\kern-1.0pt\lambda}}$, see Eq.(2.7). By Homomorphism Theorem, the homomorphism $\delta$ induces an injective homomorphism:
$\hat{\delta}:~{}~{}{\cal R}_{\kern-1.0pt\lambda}=F[X]/\langle X^{n}-\lambda\rangle~{}\to~{}{\rm M}_{n}(F),~{}~{}~{}a(X)~{}\mapsto~{}a(P_{\!\lambda}).$
The image of this homomorphism $\hat{\delta}$ is exactly ${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$, cf. Definition 3.6. Thus ${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$ is a subalgebra of ${\rm M}_{n}(F)$ and the homomorphism $\hat{\delta}$ induces the algebra isomorphism $\nu$ in Eq.(4.2). ∎
In the rest of this subsection we assume that $\lambda^{1+p^{h}}=1$ i.e., $\lambda=\lambda^{-p^{\ell-h}}$, cf. Remark 3.4. By Eq.(3.3), we have the following map
$\tilde{\tau}:~{}{\rm M}_{n}(F)\,\longrightarrow\,{\rm M}_{n}(F),~{}~{}~{}A\,\longmapsto\,A^{*h};$
and by Eq.(3.4), for any $A,B\in{\rm M}_{n}(F)$ and $\alpha\in F$,
$(AB)^{*h}=B^{*h}A^{*h},~{}~{}(A+B)^{*h}=A^{*h}+B^{*h},~{}~{}(\alpha A)^{*h}=\alpha^{p^{h}}\!A^{*h}.$
So $\tilde{\tau}$ is a $p^{h}$-anti-automorphism of the matrix algebra ${\rm M}_{n}(F)$.
Lemma 4.7.
Keep the notation as above. Assume that $\lambda^{1+p^{h}}=1$. Then $\tilde{\tau}\big{(}{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)\big{)}\subseteq{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$, and the restricted map $\tau=\tilde{\tau}|_{{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)}$ as follows is a $p^{h}$-automorphism of the $F$-algebra ${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$:
$\tau:\;{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)\;\mathop{\longrightarrow}\;{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F),\quad a(P_{\lambda})\,\longmapsto\,\big{(}a(P_{\lambda})\big{)}^{*h}.$ (4.3)
Proof.
Since $\lambda=\lambda^{-p^{\ell-h}}$, by Eq.(3.7) we deduce that
$P_{\lambda}^{\,*h}=P_{\!\lambda}^{-1}=P_{\!\lambda}^{nt-1}\in{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F).$ (4.4)
By Eq.(4.2), any $a(P_{\lambda})\in{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$ is associated with $a(X)\in{{\cal R}_{\kern-1.0pt\lambda}}$, where $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}\in{\cal R}_{\kern-1.0pt\lambda}$, so
$\big{(}a(P_{\lambda})\big{)}^{*h}=\Big{(}\sum_{i=0}^{n-1}a_{i}P_{\lambda}^{\,i}\Big{)}^{*h}=\sum_{i=0}^{n-1}a_{i}^{p^{h}}(P_{\lambda}^{\,*h})^{i}\in{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F).$
Thus $\tilde{\tau}\big{(}{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)\big{)}\subseteq{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$. Restricting the $p^{h}$-anti-automorphism $\tilde{\tau}$ to ${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)$, we get the $p^{h}$-anti-automorphism Eq.(4.3), which is in fact a $p^{h}$-automorphism because ${\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)\cong{\cal R}_{\kern-1.0pt\lambda}$ is a commutative algebra. ∎
Next, we introduce an operator “*” on ${\cal R}_{\kern-1.0pt\lambda}$, which is the key to obtaining the necessary and sufficient conditions for $2$-quasi $\lambda$-constacyclic codes being Galois $p^{h}$-self-dual. With the isomorphism $\nu$ in Eq.(4.2), inspiring by Eq.(4.3) and Eq.(4.4), for $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}\in{\cal R}_{\kern-1.0pt\lambda}$ we define
$a^{*}(X)=a^{(p^{h})}(X^{nt-1})~{}({\rm mod}~{}X^{n}-\lambda),$ (4.5)
where $a^{(p^{h})}(X)=\sum_{i=0}^{n-1}a_{i}^{p^{h}}X^{i}$, cf. Example 2.1(1).
Lemma 4.8.
Assume that $\lambda^{1+p^{h}}=1$. Let $a^{*}(X)$, $\nu$ and $\tau$ be as in Eq.(4.5), Eq.(4.2) and Eq.(4.3) respectively. Then
$a^{*}(X)=\nu^{-1}\tau\nu\big{(}a(X)\big{)},~{}~{}~{}~{}\forall\;a(X)\in{\cal R}_{\kern-1.0pt\lambda};$ (4.6)
and the following map is a $p^{h}$-automorphism of the algebra ${\cal R}_{\kern-1.0pt\lambda}$:
$*:~{}{\cal R}_{\kern-1.0pt\lambda}\;\mathop{\longrightarrow}\;{\cal R}_{\kern-1.0pt\lambda},~{}~{}~{}a(X)\,\longmapsto\,a^{*}(X)\,.$ (4.7)
Proof.
For $a(X)\in{\cal R}_{\kern-1.0pt\lambda}$, by the definition of $a^{*}(X)$ in Eq.(4.5) and by Eq.(4.4),
$a^{*}(P_{\!\lambda})$ $=\sum_{i=0}^{n-1}a_{i}^{p^{h}}(P_{\!\lambda}^{nt-1})^{i}=\sum_{i=0}^{n-1}a_{i}^{p^{h}}(P_{\!\lambda}^{*h})^{i}=\sum_{i=0}^{n-1}a_{i}^{p^{h}}(P_{\!\lambda}^{i})^{*h}$
$=\sum_{i=0}^{n-1}(a_{i}P_{\!\lambda}^{i})^{*h}=\Big{(}\sum_{i=0}^{n-1}a_{i}P_{\!\lambda}^{i}\Big{)}^{*h}=\big{(}a(P_{\!\lambda})\big{)}^{*h}.$
That is, $\nu\big{(}a^{*}(X)\big{)}=a^{*}(P_{\!\lambda})=\big{(}a(P_{\!\lambda})\big{)}^{*h}=\tau\nu\big{(}a(X)\big{)}$. So Eq.(4.6) holds; equivalently, the following diagram is commutative:
$\begin{array}[]{ccc}{\cal R}_{\kern-1.0pt\lambda}&\mathop{\longrightarrow}\limits^{\nu}&{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)\\[5.0pt] *\,\big{\downarrow}{}&&\big{\downarrow}\,\tau\\[5.0pt] {\cal R}_{\kern-1.0pt\lambda}&\mathop{\longrightarrow}\limits_{\nu}&{\rm M}_{n\mbox{-}\lambda\mbox{-}{\rm circ}}(F)\end{array}$ (4.11)
Because $\tau$ is a $p^{h}$-automorphism and both $\nu$ and $\nu^{-1}$ are algebra isomorphisms, by Eq.(4.6) we see that Eq.(4.7) is a $p^{h}$-automorphism of ${\cal R}_{\kern-1.0pt\lambda}$. ∎
Recall that the $2$-quasi $\lambda$-constacyclic code $C_{a,a^{\prime}}$ generated by $(a,a^{\prime})\in{\cal R}_{\kern-1.0pt\lambda}^{2}$ has been defined in Remark 3.5. For the ${\cal R}_{\kern-1.0pt\lambda}$-submodules of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ generated by one element, we have the following Galois self-duality criteria.
Lemma 4.9.
Assume that $\lambda^{1+p^{h}}=1$.
(1) The $2$-quasi $\lambda$-constacyclic code $C_{a,a^{\prime}}$ is Galois $p^{h}$-self-dual if and only if $aa^{*}+a^{\prime}a^{\prime*}\!=\!0$ and the rate ${\rm R}(C_{a,a^{\prime}})=1/2$.
(2) The $2$-quasi $\lambda$-constacyclic code $C_{1,g}$ is Galois $p^{h}$-self-dual if and only if $gg^{*}=-1$.
Proof.
(1) The $C_{a,a^{\prime}}$ is Galois $p^{h}$-self-dual if and only if $\langle C_{a,a^{\prime}},C_{a,a^{\prime}}\rangle_{h}=0$ and $\dim_{F}C_{a,a^{\prime}}=n$. What remains is to show that
$\langle C_{a,a^{\prime}},C_{a,a^{\prime}}\rangle_{h}=0\;\iff\;aa^{*}+a^{\prime}{a^{\prime}}^{*}=0.$ (4.12)
Observe that by Lemma 3.7, $C_{a,a^{\prime}}$ is linearly generated by the rows of the matrix $\big{(}a(P_{\!\lambda}),\,a^{\prime}(P_{\!\lambda})\big{)}$. Thus $\langle C_{a,a^{\prime}},C_{a,a^{\prime}}\rangle_{h}=0$ if and only if
$\big{(}a(P_{\!\lambda}),\,a^{\prime}(P_{\!\lambda})\big{)}\cdot\big{(}a(P_{\!\lambda}),\,a^{\prime}(P_{\!\lambda})\big{)}^{*h}=a(P_{\!\lambda})\cdot a(P_{\!\lambda})^{*h}+a^{\prime}(P_{\!\lambda})\cdot a^{\prime}(P_{\!\lambda})^{*h}=0.$
Thus, Eq.(4.12) follows from Lemma 4.8 (cf. Eq.(4.11)) immediately.
(2)  By Corollary 3.8, $C_{1,g}$ has the generating matrix $\big{(}E_{n},\,g(P_{\!\lambda})\big{)}$. In particular, $\dim_{F}C_{1,g}=n$, i.e., ${\rm R}(C_{1,g})=1/2$. Similarly to the proof of $(1)$, we have $\langle C_{1,g},C_{1,g}\rangle_{h}=0$ if and only if $01+gg^{*}=$ 0. ∎
In the semisimple case, extending [17, Theorem 4.2], we have the following theorem to characterize the Galois self-dual $2$-quasi $\lambda$-constacyclic codes. If $\gcd(n,q)=1$, by Corollary 3.11 any ${\cal R}_{\kern-1.0pt\lambda}$-submodule $C$ of ${\cal R}_{\kern-1.0pt\lambda}^{2}$ can be written as
$C=C_{a,0}\oplus C_{0,a^{\prime}}\oplus C_{b,bg},$ (4.13)
where $a,a^{\prime},b\in{{\cal R}_{\kern-1.0pt\lambda}}$ with $ab=a^{\prime}b=0$ and $g\in{C_{b}^{\times}}$.
Theorem 4.10.
Assume that $\lambda^{1+p^{h}}=1$ and $\gcd(n,q)=1$. Let $C$ in Eq.(4.13) be any $2$-quasi $\lambda$-constacyclic code. Then $C$ is Galois $p^{h}$-self-dual if and only if the following two hold:
(1)  $aa^{*}=a^{\prime}{a^{\prime}}^{*}=ab^{*}=a^{*}b=a^{\prime}b^{*}=a^{\prime*}b=bb^{*}(1+gg^{*})=0;$
(2)  $\dim_{F}C=n$.
Proof.
The $C$ is Galois $p^{h}$-self-dual if and only if $\langle C,C\rangle_{h}=0$ and $\dim_{F}C=n$. Note that the inner product $\langle-,-\rangle_{h}$ is linear for the first variable, and it is $p^{h}$-linear for the second variable, but it is not symmetric in general. So $\langle C,C\rangle_{h}=0$ is equivalent to the following
$\langle C_{a,0},C_{0,a^{\prime}}\rangle_{h}=\langle C_{0,a^{\prime}},C_{a,0}\rangle_{h}=0,$
$\langle C_{a,0},C_{a,0}\rangle_{h}=\langle C_{0,a^{\prime}},C_{0,a^{\prime}}\rangle_{h}=0,$
$\langle C_{a,0},C_{b,bg}\rangle_{h}=\langle C_{b,bg},C_{a,0}\rangle_{h}=0,$
$\langle C_{0,a^{\prime}},C_{b,bg}\rangle_{h}=\langle C_{b,bg},C_{0,a^{\prime}}\rangle_{h}=0,$
$\langle C_{b,bg},C_{b,bg}\rangle_{h}=0.$
The first line holds obviously. By Eq.(4.12), the second and the third lines are equivalent to that $aa^{*}=a^{\prime}a^{\prime*}=0$ and $ab^{*}=a^{*}b=0$, respectively. The last line is equivalent to that $bb^{*}(1+gg^{*})=0$.
Turn to the forth line which is equivalent to $a^{\prime}b^{*}g^{*}=a^{\prime*}bg=0$. Note that $C_{b}={\cal R}_{\kern-1.0pt\lambda}\,e_{b}$ is a ring with identity $e_{b}$ which is an idempotent, cf. Remark 2.4(1), hence “$g\in C_{b}^{\times}$” implies that $gg^{\prime}=e_{b}$ for a $g^{\prime}\in C_{b}$. So $a^{\prime*}b=a^{\prime*}be_{b}=a^{\prime*}bgg^{\prime}=0g^{\prime}=0$. Similarly, $a^{\prime}b^{*}=a^{\prime}b^{*}e_{b}^{*}=a^{\prime}b^{*}g^{*}g^{\prime*}=0g^{\prime*}=0$.
The theorem is proved. ∎
5 Hermitian self-dual $2$-quasi-cyclic codes
In this section, we always assume that $\ell$ is even and $h=\ell/2$, and ${\gcd(n,q)=1}$. The map $\sigma_{\ell/2}:F\to F$, $\alpha\mapsto\alpha^{p^{\ell/2}}$, is a Galois automorphism of order $2$, and
$\langle{\bf a},{\bf a^{\prime}}\rangle_{\ell/2}=\sum_{i=0}^{n-1}a_{i}{a^{\prime}_{i}}^{p^{\ell/2}},~{}~{}~{}~{}\forall\,{\bf a}=(a_{0},\cdots,a_{n-1}),\,{\bf a^{\prime}}=(a^{\prime}_{0},\cdots,a^{\prime}_{n-1})\in F^{n},$
is the Hermitian inner product on $F^{n}$. In this section we consider ${\cal R}={\cal R}_{1}=F[X]/\langle X^{n}-1\rangle$ and ${\cal R}^{2}={\cal R}\times{\cal R}$, cf. Eq.(2.13); and prove that the Hermitian self-dual $2$-quasi-cyclic codes are asymptotically good.
5.1 The operator “*” on ${\cal R}=F[X]/\langle X^{n}-1\rangle$
Since $11^{1+p^{\ell/2}}=$ 1, the results in Subsection 4.2 can be quoted freely for ${\cal R}$ and ${\cal R}^{2}$. In particular, the operator “*” in Lemma 4.8 is a $p^{\ell/2}$-automorphism of ${\cal R}$:
$*:~{}~{}{\cal R}\;\longrightarrow\;{\cal R},~{}~{}~{}a(X)\;\longmapsto\;a^{*}(X),$ (5.1)
where $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}$ and $a^{*}(X)=a^{(p^{\ell/2})}(X^{n-1})~{}({\rm mod}~{}X^{n}-1)$, i.e., $a^{*}(X)=\sum_{i=0}^{n-1}a_{i}^{p^{\ell/2}}(X^{n-1})^{i}~{}({\rm mod}~{}X^{n}-1)$, cf. Eq.(4.5). By Eq.(2.14), we have the identification:
${\cal R}\cong FG,~{}~{}~{}~{}\sum_{i=0}^{n-1}a_{i}X^{i}~{}\mapsto~{}\sum_{i=0}^{n-1}a_{i}x^{i},$
where $G=\langle\,x\,|\,x^{n}=1\rangle$ is the cyclic group of order $n$ and $FG$ is the cyclic group algebra. So the symbol “$X$” can be identified with the element $x$ of the group $G$, and we can write $a(x)=\sum_{i=0}^{n-1}a_{i}x^{i}\in{\cal R}$, and the expressions $x^{-1}$, $a(x^{-1})$ etc. make sense. Hence $x^{n-1}=x^{-1}$ and
$a^{*}(x)=\sum_{i=0}^{n-1}a_{i}^{p^{\ell/2}}(x^{-1})^{i}={a}^{(p^{\ell/2})}(x^{-1}),~{}~{}~{}~{}\forall\;a(x)\in{\cal R}.$ (5.2)
Lemma 5.1.
The $p^{\ell/2}$-automorphism “*” of ${\cal R}$ in Eq.(5.1) is of order 2.
Proof.
Since $x^{n-1}=x^{-1}$, for any $a(x)\in{\cal R}$ by Eq.(5.2) we have
$\Big{(}\sum_{i=0}^{n-1}a_{i}x^{i}\Big{)}^{**}=\Big{(}\sum_{i=0}^{n-1}a_{i}^{p^{\ell/2}}(x^{-1})^{i}\Big{)}^{*}=\sum_{i=0}^{n-1}(a_{i}^{p^{\ell/2}})^{p^{\ell/2}}((x^{-1})^{-1})^{i}=\sum_{i=0}^{n-1}a_{i}x^{i}.$
Thus the order of the operator “*” equals $1$ or $2$. There is an $\alpha\in F$ such that $\alpha^{p^{\ell/2}}\!\neq\alpha$. Then in ${\cal R}$ we have $(\alpha 1)^{*}\neq\alpha 1$. So the order of the operator “*” equals $2$. ∎
Corollary 5.2.
If $C$ is a non-zero ideal of ${\cal R}$ which is invariant by the operator “*” (i.e., $C^{*}=C$), then the restriction of the operator “*” to $C$ induces a $p^{\ell/2}$-automorphism of $C$ of order $2$.
Proof.
By Lemma 5.1, the restriction of “*” to $C$ is of order $1$ or $2$. Let $c\in C$ with $c\neq 0$. If $c^{*}\neq c$, then the restriction of “*” to $C$ is not the identity. Otherwise, $c^{*}=c$; there is an $\alpha\in F$ with $\alpha^{p^{\ell/2}}\!\neq\alpha$; so $(\alpha c)^{*}=\alpha^{p^{\ell/2}}\!c^{*}=\alpha^{p^{\ell/2}}c\neq\alpha c$. In conclusion, the restriction of “*” to $C$ is of order $2$. ∎
Note that “$\gcd(n,q)=1$” is assumed in this section. Recall from Eq.(2.15) that $e_{0}=\frac{1}{n}\sum_{i=0}^{n-1}x^{i}$, $e_{1},\dots,e_{m}$ are all primitive idempotents of ${\cal R}$. Then the $p^{h}$-automorphism “*” permutes the primitive idempotents, i.e., every $e_{i}^{*}$ is still a primitive idempotent. Note that $e_{0}^{*}=e_{0}$. We can reorder the other primitive idempotents
$e_{1},\dots,e_{r},\;e_{r+1},e_{r+1}^{*},\dots,e_{r+s},e_{r+s}^{*},$ (5.3)
such that $e_{i}^{*}=e_{i}$ for $i=1,\dots,r$, $e_{r+j}^{*}\neq e_{r+j}$ but $e_{r+j}^{**}=e_{r+j}$ for $j=1,\dots,s$.
Remark 5.3.
(1) For $i=0,1,\dots,r$, we get $({\cal R}e_{i})^{*}={\cal R}e_{i}^{*}={\cal R}e_{i}$. The restriction of the map “*” in Eq.(5.1) to ${\cal R}e_{i}$ induces a $p^{\ell/2}$-automorphism of ${\cal R}e_{i}$ of order 2 (cf. Corollary 5.2) as follows
$*\!|_{{\cal R}e_{i}}:~{}{\cal R}e_{i}\rightarrow{\cal R}e_{i},~{}~{}~{}a\mapsto a^{*}.$
Let $d_{i}=\dim_{F}{\cal R}e_{i}$. By Eq.(2.19), the ideal ${\cal R}e_{i}$ is a field extension over $F$ with cardinality $|{\cal R}e_{i}|=p^{d_{i}\ell}$. So $a^{*}=a^{p^{d_{i}\ell/2}}$ for $a\in{\cal R}e_{i}$.
(2) For $j=1,\dots,s$, $({\cal R}e_{r+j})^{*}={\cal R}e_{r+j}^{*}\neq{\cal R}e_{r+j}$, and the restriction of the map “*” in Eq.(5.1) to ${\cal R}e_{r+j}$ induces a $p^{\ell/2}$-isomorphism as follows
$*\!|_{{\cal R}e_{r+j}}:~{}{\cal R}e_{r+j}\rightarrow{\cal R}e_{r+j}^{*},~{}~{}~{}a\mapsto a^{*}.$
Let $d_{r+j}=\dim_{F}{\cal R}e_{r+j}$, then $\dim_{F}{\cal R}e_{r+j}^{*}=\dim_{F}{\cal R}e_{r+j}=d_{r+j}$. Denote $\widehat{e}_{r+j}=e_{r+j}+e_{r+j}^{*}$, by Eq.(2.19) we have that
${\cal R}\widehat{e}_{r+j}={\cal R}e_{r+j}\oplus{\cal R}e_{r+j}^{*}=\{a^{\prime}+a^{\prime\prime}\,|\,a^{\prime}\in{\cal R}e_{r+j},\,a^{\prime\prime}\in{\cal R}e_{r+j}^{*}\},$
$\dim_{F}{\cal R}\widehat{e}_{r+j}=2d_{r+j}.$
It is easy to check that $\widehat{e}_{r+j}^{\,*}=(e_{r+j}+e_{r+j}^{*})^{*}=e_{r+j}^{*}+e_{r+j}=\widehat{e}_{r+j}$. So $({\cal R}\widehat{e}_{r+j})^{*}={\cal R}\widehat{e}_{r+j}^{\;*}={\cal R}\widehat{e}_{r+j}$. The restriction of the map “*” in Eq.(5.1) to ${\cal R}\widehat{e}_{r+j}$ induces a $p^{\ell/2}$-automorphism of ${\cal R}\widehat{e}_{r+j}$ of order 2 (cf. Corollary 5.2) as follows
$*\!|_{{\cal R}\widehat{e}_{r+j}}:~{}{\cal R}\widehat{e}_{r+j}\rightarrow{\cal R}\widehat{e}_{r+j},~{}~{}~{}a^{\prime}+a^{\prime\prime}\,\mapsto\,a^{\prime\prime*}+a^{\prime*}.$ (5.4)
For convenience, in the following we denote $\widehat{e}_{i}=e_{i}$ for $i=0,1,\dots,r$. Then $\widehat{e}_{i}^{\,*}=\widehat{e}_{i}$ for $i=0,1,\dots,r+s$, and
$1=\widehat{e}_{0}+\widehat{e}_{1}+\dots+\widehat{e}_{r+s},~{}~{}~{}~{}~{}\widehat{e}_{i}\widehat{e}_{j}=\begin{cases}\widehat{e}_{i},&i=j;\\ 0,&i\neq j.\end{cases}$ (5.5)
Thus, ${\cal R}$ can be rewritten as
${\cal R}={\cal R}\widehat{e}_{0}\oplus{\cal R}\widehat{e}_{1}\oplus\dots\oplus{\cal R}\widehat{e}_{r}\oplus{\cal R}\widehat{e}_{r+1}\oplus\dots\oplus{\cal R}\widehat{e}_{r+s}.$ (5.6)
Any $a\in{\cal R}$ is decomposed into
$a=a_{0}+a_{1}+\cdots+a_{r+s},~{}~{}~{}~{}~{}\mbox{where}~{}~{}a_{i}=a\widehat{e}_{i},~{}i=0,1,\dots,r+s.$ (5.7)
The $a_{i}=a\widehat{e}_{i}$ is called the $\widehat{e}_{i}$-component of $a$. For $a=\sum_{i=0}^{r+s}a_{i},b=\sum_{i=0}^{r+s}b_{i}\in{\cal R}$, by Eq.(5.5) it is trivial to check that
$ab=a_{0}b_{0}+a_{1}b_{1}+\cdots+a_{r+s}b_{r+s}.$ (5.8)
Keep the notation in Remark 5.3, we have $d_{i}=\dim_{F}{\cal R}\widehat{e}_{i}$ for $i=0,1,\dots,r$; and $2d_{i}=\dim_{F}{\cal R}\widehat{e}_{i}$ for $i=r+1,\dots,r+s$. Thus,
$\dim_{F}{{\cal R}}=d_{0}+d_{1}+\dots+d_{r}+2d_{r+1}+\dots+2d_{r+s}=n,$ (5.9)
where $d_{0}=1$ and $d_{i}\geq\mu(n)$ for $i=1,\dots,r+s$; cf. Eq.(2.21).
5.2 A class of Hermitian self-dual $2$-quasi-cyclic codes
Recall that the $2$-quasi-cyclic code $C_{1,g}$ of ${\cal R}^{2}$ (defined in Remark 3.5) is Hermitian self-dual if and only if $gg^{*}=-1$, see Lemma 4.9. So we denote
${\cal D}=\big{\{}\,g\;\big{|}\;g\in{\cal R},\;gg^{*}=-1\,\big{\}}.$ (5.10)
Any $g\in{\cal D}$ corresponds to a Hermitian self-dual $2$-quasi-cyclic code $C_{1,g}$. Set
${\cal D}_{i}=\{\,z\;|\;z\in{\cal R}\widehat{e}_{i},\;zz^{*}=-\widehat{e}_{i}\},~{}~{}~{}~{}~{}i=0,1,\dots,r+s.$ (5.11)
By Eq.(5.5), Eq.(5.7) and Eq.(5.8),
$g=g_{0}+g_{1}+\dots+g_{r+s}\in{\cal D}~{}\iff~{}g_{i}\in{\cal D}_{i},~{}\forall\;i=0,1,\dots,r+s.$ (5.12)
Lemma 5.4.
(1) If $0\leq i\leq r$ (i.e. $\widehat{e}_{i}=e_{i}=e_{i}^{*}$), then $|{\cal D}_{i}|=p^{d_{i}\ell/2}+1$.
(2) If $r<i\leq r+s$ (i.e. $e_{i}\neq e_{i}^{*}$ and $\widehat{e}_{i}=e_{i}+e_{i}^{*}$), then $|{\cal D}_{i}|=p^{d_{i}\ell}-1$.
Proof.
(1). By Remark 5.3(1), ${\cal R}\widehat{e}_{i}$ is the field with $|{\cal R}\widehat{e}_{i}|=p^{d_{i}\ell}$. For $z\in{\cal R}\widehat{e}_{i}$, $zz^{*}=zz^{p^{d_{i}\ell/2}}\!=z^{p^{d_{i}\ell/2}+1}$. So, $zz^{*}=-\widehat{e}_{i}$ if and only if ${z^{p^{d_{i}\ell/2}+1}+\widehat{e}_{i}=0}$. Hence $|{\cal D}_{i}|$ equals the number of the roots in $F_{p^{d_{i}\ell}}$ of the $F_{p^{d_{i}\ell}}$-polynomial $X^{p^{d_{i}\ell/2}+1}+1$, where $F_{p^{d_{i}\ell}}$ denotes the finite field with cardinality $p^{d_{i}\ell}$.
If $p=2$, then $1=-1$. The order of the multiplicative group $F_{p^{d_{i}\ell}}^{\times}$ is
$\big{|}F_{p^{d_{i}\ell}}^{\times}\big{|}=p^{d_{i}\ell}-1=\big{(}p^{d_{i}\ell/2}-1\big{)}\big{(}p^{d_{i}\ell/2}+1\big{)}.$ (5.13)
Thus the multiplication group $F_{p^{d_{i}\ell}}^{\times}$ has a subgroup $H$ of order $p^{d_{i}\ell/2}+1$, and all elements of $H$ are roots of the polynomial $X^{p^{d_{i}\ell/2}+1}+1$. Hence $|{\cal D}_{i}|=p^{d_{i}\ell/2}+1$.
Otherwise $p$ is odd, then $p^{d_{i}\ell/2}-1$ is even. By Eq.(5.13), we get that $2(p^{d_{i}\ell/2}+1)\,\big{|}\,|F_{p^{d_{i}\ell}}^{\times}|$. So $F_{p^{d_{i}\ell}}^{\times}$ has a subgroup $H$ of order $2(p^{d_{i}\ell/2}+1)$. The elements of $H$ are just all roots of the polynomial $X^{2(p^{d_{i}\ell/2}+1)}-1$. Since
$X^{2(p^{d_{i}\ell/2}+1)}-1=(X^{p^{d_{i}\ell/2}+1}-1)(X^{p^{d_{i}\ell/2}+1}+1),$
all roots of the polynomial $X^{p^{d_{i}\ell/2}+1}+1$ are inside $F_{p^{d_{i}\ell}}$. Thus, $|{\cal D}_{i}|=p^{d_{i}\ell/2}+1$.
(2). For $r<i\leq r+s$, $e_{i}\neq e_{i}^{*}$ and $\widehat{e}_{i}=e_{i}+e_{i}^{*}$. By Remark 5.3(2), ${\cal R}\widehat{e}_{i}={\cal R}e_{i}\oplus{\cal R}e_{i}^{*}$. For $z=z^{\prime}+z^{\prime\prime}\in{\cal R}e_{i}\oplus{\cal R}e_{i}^{*}$ with $z^{\prime}\in{\cal R}e_{i}$ and $z^{\prime\prime}\in{\cal R}e_{i}^{*}$, by Eq.(5.4) we have $z^{*}=z^{\prime\prime*}+z^{\prime*}$, and so $zz^{*}=(z^{\prime}+z^{\prime\prime})(z^{\prime\prime*}+z^{\prime*})=z^{\prime}z^{\prime\prime*}+z^{\prime*}z^{\prime\prime}$. It follows that: $zz^{*}=-\widehat{e}_{i}=-e_{i}-e_{i}^{*}$ if and only if $z^{\prime}z^{\prime\prime*}=-e_{i}$ and $z^{\prime*}z^{\prime\prime}=-e_{i}^{*}$. We take $z^{\prime}\in({\cal R}e_{i})^{\times}$, then $z^{\prime\prime*}=-z^{\prime-1}$ (where $z^{\prime-1}$ is the inverse of $z^{\prime}$ in ${\cal R}e_{i}$, not in ${\cal R}$), hence $z^{\prime\prime}=z^{\prime\prime**}=-(z^{\prime*})^{-1}$ is uniquely determined. Thus,
$zz^{*}=-\widehat{e}_{i}~{}\iff~{}z=z^{\prime}-(z^{\prime*})^{-1}~{}~{}\mbox{for a}~{}z^{\prime}\in({\cal R}e_{i})^{\times}.$ (5.14)
Since ${\cal R}e_{i}$ is a field with cardinality $p^{d_{i}\ell}$, $|{\cal D}_{i}|=|({\cal R}e_{i})^{\times}|=p^{d_{i}\ell}-1$. ∎
For any subset $\omega\subseteq\{0,1,\dots,r\!+\!s\}$, refining the notation in Eq.(5.6) and in Remark 5.3, we define an ideal $A_{\omega}$ of ${\cal R}$ and an integer $d_{\omega}$ as follows
$\textstyle A_{\omega}=\bigoplus\limits_{i\in\omega}{\cal R}\widehat{e}_{i},~{}~{}~{}~{}d_{\omega}=\dim_{F}A_{\omega}.$
Obviously, $\omega$ can be written as a disjoint union:
$\omega=\omega^{\prime}\cup\omega^{\prime\prime},~{}~{}~{}\mbox{where }\omega^{\prime}\!=\!\omega\cap\{0,1,\dots,r\},~{}\omega^{\prime\prime}\!=\!\omega\cap\{r\!+\!1,\dots,r\!+\!s\}.$
Similarly to Eq.(5.9), we have
$A_{\omega}=A_{\omega^{\prime}}\oplus A_{\omega^{\prime\prime}},~{}~{}~{}~{}d_{\omega}=\sum_{i\in\omega^{\prime}}d_{i}+2\sum_{i\in\omega^{\prime\prime}}d_{i}.$ (5.15)
It is known that ([23, Lemma 4.7]) for integers $k_{1},\dots,k_{v}$, if $k_{i}\geq\log_{p}(v)$ for $i=1,\dots,v$, then
$\begin{array}[]{l}(p^{k_{1}}-1)\dots(p^{k_{v}}-1)\geq p^{k_{1}+\dots+k_{v}-2};\\[3.0pt] (p^{k_{1}}+1)\dots(p^{k_{v}}+1)\leq p^{k_{1}+\dots+k_{v}+2}.\end{array}$ (5.18)
Lemma 5.5.
For a subset $\omega\subseteq\{0,1,\dots,r+s\}$, if $\mu(n)\geq\log_{p}(n)$ (where $\mu(n)$ is defined in Eq.(2.21)), then
$p^{-2}p^{d_{\omega}\ell/2}\leq\prod_{i\in\omega}|{\cal D}_{i}|\leq p^{3}p^{d_{\omega}\ell/2}.$
Proof.
For $1\leq i\leq r+s$, we deduce that $d_{i}\geq\mu(n)\geq\log_{p}(r+s)$ since $n\geq r+s$. By Lemma 5.4 we have that
$\prod_{i\in\omega}|{\cal D}_{i}|$ $=\prod_{i\in\omega^{\prime}}|{\cal D}_{i}|\cdot\prod_{i\in\omega^{\prime\prime}}|{\cal D}_{i}|=\prod\limits_{i\in\omega^{\prime}}(p^{d_{i}\ell/2}+1)\cdot\prod\limits_{i\in\omega^{\prime\prime}}(p^{d_{i}\ell}-1).$
If $0\notin\omega$, using Eq.(5.18) we get
$\prod_{i\in\omega}|{\cal D}_{i}|$ $\geq\prod\limits_{i\in\omega^{\prime}}(p^{d_{i}\ell/2}-1)\cdot\prod\limits_{i\in\omega^{\prime\prime}}(p^{d_{i}\ell}-1)\geq p^{\big{(}\sum_{i\in\omega^{\prime}}d_{i}\ell/2+\sum_{i\in\omega^{\prime\prime}}d_{i}\ell\big{)}-2}$
$=p^{-2}p^{\big{(}\sum_{i\in\omega^{\prime}}d_{i}+\sum_{i\in\omega^{\prime\prime}}2d_{i}\big{)}\ell/2}=p^{-2}p^{d_{\omega}\ell/2},$
where the last equality follows by Eq.(5.15); and
$\prod_{i\in\omega}|{\cal D}_{i}|$ $\leq\prod\limits_{i\in\omega^{\prime}}(p^{d_{i}\ell/2}+1)\cdot\prod\limits_{i\in\omega^{\prime\prime}}(p^{d_{i}\ell}+1)\leq p^{\big{(}\sum_{i\in\omega^{\prime}}d_{i}\ell/2+\sum_{i\in\omega^{\prime\prime}}d_{i}\ell\big{)}+2}$
$=p^{2}p^{\big{(}\sum_{i\in\omega^{\prime}}d_{i}+\sum_{i\in\omega^{\prime\prime}}2d_{i}\big{)}\ell/2}=p^{2}p^{d_{\omega}\ell/2}.$
That is, if $0\notin\omega$ then
$p^{-2}p^{d_{\omega}\ell/2}\leq\prod_{i\in\omega}|{\cal D}_{i}|\leq p^{2}p^{d_{\omega}\ell/2}.$ (5.19)
If $0\in\omega$, we set $\tilde{\omega}=\omega\setminus\{0\}$, and so $d_{\omega}=d_{0}+d_{\tilde{\omega}}$, where $d_{0}=1$. By Lemma 5.4(1), we see that $|{\cal D}_{0}|=p^{d_{0}\ell/2}+1$, and
$\prod_{i\in\omega}|{\cal D}_{i}|$ $=(p^{d_{0}\ell/2}+1)\prod\limits_{i\in\tilde{\omega}}|{\cal D}_{i}|.$
It follows by Eq.(5.19) that
$\prod_{i\in\omega}|{\cal D}_{i}|\geq p^{d_{0}\ell/2}\cdot p^{-2}p^{d_{\tilde{\omega}}\ell/2}$ $=p^{-2}p^{\big{(}d_{0}+d_{\tilde{\omega}}\big{)}\ell/2}=p^{-2}p^{d_{\omega}\ell/2};$
and that
$\prod_{i\in\omega}|{\cal D}_{i}|\leq p^{1+d_{0}\ell/2}\cdot p^{2}p^{d_{\tilde{\omega}}\ell/2}=p^{3}p^{\big{(}d_{0}+d_{\tilde{\omega}}\big{)}\ell/2}=p^{3}p^{d_{\omega}\ell/2}.$
Thus the lemma holds. ∎
By Eq.(5.12), $|{\cal D}|=\prod_{i=0}^{r+s}|{\cal D}_{i}|$. We have the following at once.
Corollary 5.6.
If $\mu(n)\geq\log_{p}(n)$ (where $\mu(n)$ is defined in Eq.(2.21)), then
$p^{-2}p^{n\ell/2}\leq|{\cal D}|\leq p^{3}p^{n\ell/2}.$
Any element $a\in{\cal R}$ can be written as $a=a_{0}\hat{e}_{0}+a_{1}\hat{e}_{1}+\dots+a_{r+s}\hat{e}_{r+s}$ (cf. Eq.(5.7)). We denote
$\omega_{a}=\big{\{}i\,\big{|}\,a_{i}=a\widehat{e}_{i}\neq 0\big{\}}\subseteq\{0,1,\dots,r+s\};~{}~{}d_{a}=d_{\omega_{a}}=\dim_{F}A_{\omega_{a}}.$
Obviously ${\cal R}a\subseteq A_{\omega_{a}}$. For $a,b\in{\cal R}$ it is easy to check that
${\cal R}a={\cal R}b~{}~{}\implies~{}~{}A_{\omega_{a}}=A_{\omega_{b}}~{}(\mbox{equivalently, }\omega_{a}=\omega_{b});$ (5.20)
but the converse is not true in general.
Lemma 5.7.
For $(a,b)\in{\cal R}^{2}$, we denote
${\cal D}_{(a,b)}=\big{\{}g\>\big{|}\;g\in{\cal D},\;(a,b)\in C_{1,g}\big{\}},$ (5.21)
where $C_{1,g}$ is defined in Remark 3.5 and ${\cal D}$ is defined in Eq.(5.10).
(1) If ${\cal D}_{(a,b)}\neq\emptyset$, then ${\cal R}a={\cal R}b$, hence $\omega_{a}=\omega_{b}$.
(2) If $\mu(n)\geq\log_{p}(n)$, then $|{\cal D}_{(a,b)}|\leq p^{5}p^{(n-d_{a})\ell/2}.$
Proof.
(1). Assume that $g\in{\cal D}_{(a,b)}$, then $gg^{*}=-1$ (cf. Eq.(5.10)) and $u(1,g)=(a,b)$ for an element $u\in{\cal R}$. The former equality implies that $g$ is invertible with $g^{-1}=-g^{*}$. The latter equality implies that $a=u$ and $b=ug=ag$. We deduce that ${\cal R}b={\cal R}ag={\cal R}ga={\cal R}a$ since $g$ is invertible.
(2). By Eq.(5.7), we write $a=a_{0}+\dots+a_{r+s}$, $b=b_{0}+\dots+b_{r+s}$ and $g=g_{0}+\dots+g_{r+s}$ with $g_{i}\in{\cal D}_{i}$. By Eq.(5.8),
$ag=b~{}~{}\iff a_{i}g_{i}=b_{i}~{}\mbox{for}~{}~{}i=0,1,\dots,r+s.$
We count the number of such $g_{i}$ in two cases.
Case 1: $a_{i}=0$, i.e., $i\notin\omega_{a}$. Since ${\cal R}a={\cal R}b$, we have that ${\cal R}b_{i}={\cal R}a_{i}=0$, hence $b_{i}=0$. Then any $g_{i}\in{\cal D}_{i}$ satisfies that $a_{i}g_{i}=b_{i}$.
Case 2: $a_{i}\neq 0$, i.e., $i\in\omega_{a}$. There are two subcases.
Subcase 2.1: $0\leq i\leq r$. Since ${\cal R}\widehat{e}_{i}$ is a field, $a_{i}$ is invertible in ${\cal R}\widehat{e}_{i}$, and so there is a unique $z\in{\cal R}\widehat{e}_{i}$ such that $a_{i}z=b_{i}$. We see that there is at most one $g_{i}\in{\cal D}_{i}$ such that $a_{i}g_{i}=b_{i}$.
Subcase 2.2: $r<i\leq r+s$. By Remark 5.3(2), ${\cal R}\widehat{e}_{i}={\cal R}e_{i}\oplus{\cal R}e_{i}^{*}$. We can write $a_{i}=\alpha^{\prime}+\alpha^{\prime\prime}$ and $b_{i}=\beta^{\prime}+\beta^{\prime\prime}$, where $\alpha^{\prime},\beta^{\prime}\in{\cal R}e_{i}$ and $\alpha^{\prime\prime},\beta^{\prime\prime}\in{\cal R}e_{i}^{*}$. Since $a_{i}\neq 0$, at least one of $\alpha^{\prime}$ and $\alpha^{\prime\prime}$ is nonzero. We may assume that $\alpha^{\prime}\neq 0$. Take $g_{i}=z^{\prime}-(z^{\prime*})^{-1}\in{\cal D}_{i}$ as in Eq.(5.14), then $a_{i}g_{i}=b_{i}$ implies that $\alpha^{\prime}z^{\prime}=\beta^{\prime}$ in ${\cal R}e_{i}$. Hence $z^{\prime}=\alpha^{\prime-1}\beta^{\prime}$ is uniquely determined.
In a word, if $a_{i}\neq 0$ (i.e., $i\in\omega_{a}$), then there is at most one $g_{i}\in{\cal D}_{i}$ such that $a_{i}g_{i}=b_{i}$. Thus
$|{\cal D}_{(a,b)}|\leq\prod_{i\notin\omega_{a}}|{\cal D}_{i}|=\prod_{i=0}^{r+s}|{\cal D}_{i}|\Big{/}\prod_{i\in\omega_{a}}|{\cal D}_{i}|=|{\cal D}|\Big{/}\prod_{i\in\omega_{a}}|{\cal D}_{i}|.$
By Lemma 5.5 and Corollary 5.6,
$|{\cal D}_{(a,b)}|\leq p^{3}p^{n\ell/2}\Big{/}p^{-2}p^{d_{\omega_{a}}\ell/2}=p^{5}p^{(n-d_{\omega_{a}})\ell/2}.$
We are done. ∎
5.3 Hermitian self-dual $2$-quasi-cyclic codes are asymptotically good
Keep the notation in Subsection 5.2. From now on, let $\delta$ be a real number satisfying that (where $h_{q}(\delta)$ is the $q$-entropy function defined in Eq.(2.22))
$\delta\in(0,1-q^{-1})\quad\mbox{and}\quad h_{q}(\delta)<1/4.$ (5.22)
And set
$\begin{array}[]{l}{\cal D}^{\leq\delta}=\big{\{}\,g\;\big{|}\;g\in{\cal D},\;\Delta(C_{1,g})\leq\delta\,\big{\}};\\[3.0pt] \Omega_{d}=\big{\{}A~{}|~{}A~{}{\rm is~{}an~{}ideal~{}of~{}}{\cal R}{\rm~{}with~{}}\dim_{F}A=d\big{\}}.\end{array}$ (5.25)
Recall that $\mu(n)$ is defined in Eq.(2.21), $(A\times A)^{\leq\delta}$ is defined in Eq.(2.23) and ${\cal D}_{(a,b)}$ is defined in Eq.(5.21).
Lemma 5.8.
${\cal D}^{\leq\delta}\subseteq\bigcup_{d=\mu(n)}^{n}~{}\bigcup_{A\in\Omega_{d}}~{}\bigcup_{(a,b)\in(A\times A)^{\leq\delta}}{\cal D}_{(a,b)}$.
Proof.
Assume that $g\in{\cal D}^{\leq\delta}$. By Eq.(5.25), we have an $(a,b)\in{\cal R}^{2}$ such that $0<{\rm w}(a,b)\leq 2\delta n$ and $(a,b)\in C_{1,g}$, i.e., $g\in{\cal D}_{(a,b)}$. From Lemma 5.7(1) and Eq.(5.20), we deduce that $A_{\omega_{a}}=A_{\omega_{b}}$; hence $d_{a}\!=\!\dim_{F}A_{\omega_{a}}\!=\!\dim_{F}A_{\omega_{b}}=d_{b}$, and $(a,b)\in A_{\omega_{a}}\times A_{\omega_{a}}$ and $A_{\omega_{a}}\in\Omega_{d_{a}}$. If $A_{\omega_{a}}={\cal R}e_{0}$, then $0\neq a,b\in{\cal R}e_{0}$, hence ${\rm w}(a,b)=2n$ (cf. Eq.(2.20)), which contradicts that ${\rm w}(a,b)\leq 2\delta n$ and $\delta<1-q^{-1}$ (see Eq.(5.22)). Thus $A_{\omega_{a}}\neq{\cal R}e_{0}$. By the definition of $\mu(n)$ in Eq.(2.21), we obtain that $d_{a}\geq\mu(n)$. So the lemma is proved. ∎
Lemma 5.9.
If $\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{\mu(n)\ell}>0$ and $\mu(n)\geq\log_{p}(n)$, then
$|{\cal D}^{\leq\delta}|\leq p^{5}p^{\frac{n\ell}{2}-2\mu(n)\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{\mu(n)\ell}\big{)}}.$
Proof.
Note that any ideal $A$ of ${\cal R}$ is a direct sum of some of ${\cal R}e_{0},{\cal R}e_{1},\dots,{\cal R}e_{m}$ (cf. Eq.(2.19)). For $1\leq i\leq m$, the dimension $\dim_{F}{\cal R}e_{i}\geq\mu(n)$, where $\mu(n)$ is defined in Eq.(2.21). Suppose that $A\in{\Omega_{d}}$ where $\mu(n)\leq d\leq n$. If ${\cal R}e_{0}$ is not a direct summand of $A$, then the number of such $A$ is at most $m^{\frac{d}{\mu(n)}}$; otherwise ${\cal R}e_{0}$ is a direct summand of $A$, then the number of such $A$ is at most $m^{\frac{d-1}{\mu(n)}}<m^{\frac{d}{\mu(n)}}$. Hence
$|\Omega_{d}|\leq m^{\frac{d}{\mu(n)}}\leq n^{\frac{d}{\mu(n)}}.$ (5.26)
Applying Lemma 5.8 and Lemma 5.7(2), we obtain
$|{\cal D}^{\leq\delta}|$ $\leq\sum_{d=\mu(n)}^{n}~{}\sum_{A\in\Omega_{d}}~{}\sum_{(a,b)\in(A\times A)^{\leq\delta}}|{\cal D}_{(a,b)}|$
$\leq\sum_{d=\mu(n)}^{n}~{}\sum_{A\in\Omega_{d}}~{}\sum_{(a,b)\in(A\times A)^{\leq\delta}}p^{5}p^{(n-d)\ell/2},$
where the number $p^{5}p^{(n-d)\ell/2}$ is independent of the choice of $(a,b)\in(A\times A)^{\leq\delta}$. By Lemma 2.8, for $A\in\Omega_{d}$ we have $|(A\times A)^{\leq\delta}|\leq(p^{\ell})^{2dh_{q}(\delta)}$. So
$|{\cal D}^{\leq\delta}|\leq\sum_{d=\mu(n)}^{n}~{}\sum_{A\in\Omega_{d}}~{}|(A\times A)^{\leq\delta}|\cdot p^{5}p^{(n-d)\ell/2}$
$\leq\sum_{d=\mu(n)}^{n}\sum_{A\in\Omega_{d}}p^{2d\ell h_{q}(\delta)}p^{5}p^{(n-d)\ell/2}=p^{5}\!\sum_{d=\mu(n)}^{n}\sum_{A\in\Omega_{d}}p^{\frac{n\ell}{2}+2d\ell\big{(}h_{q}(\delta)-\frac{1}{4}\big{)}},$
where the number $p^{\frac{n\ell}{2}+2d\ell\big{(}h_{q}(\delta)-\frac{1}{4}\big{)}}$ is independent of the choice of $A\in\Omega_{d}$. Using Eq.(5.26) yields
$|{\cal D}^{\leq\delta}|\leq p^{5}\!\sum_{d=\mu(n)}^{n}\!|\Omega_{d}|\!\cdot\!p^{\frac{n\ell}{2}+2d\ell\big{(}h_{q}(\delta)-\frac{1}{4}\big{)}}\leq p^{5}\!\sum_{d=\mu(n)}^{n}\!n^{\frac{d}{\mu(n)}}p^{\frac{n\ell}{2}+2d\ell\big{(}h_{q}(\delta)-\frac{1}{4}\big{)}}$
$~{}=p^{5}\!\sum_{d=\mu(n)}^{n}\!p^{\frac{n\ell}{2}+2d\ell\big{(}h_{q}(\delta)-\frac{1}{4}\big{)}+\frac{d\log_{p}(n)}{\mu(n)}}=p^{5}\!\sum_{d=\mu(n)}^{n}p^{\frac{n\ell}{2}-2d\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{2\mu(n)\ell}\big{)}}.$
Note that $\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{2\mu(n)\ell}>0$ (since $\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{\mu(n)\ell}>0$) and $d\geq\mu(n)$. We further get
$|{\cal D}^{\leq\delta}|$ $\leq p^{5}\sum_{d=\mu(n)}^{n}p^{\frac{n\ell}{2}-2\mu(n)\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{2\mu(n)\ell}\big{)}}$
$\leq p^{5}\cdot n\cdot p^{\frac{n\ell}{2}-2\mu(n)\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{2\mu(n)\ell}\big{)}}.$
That is, $|{\cal D}^{\leq\delta}|\leq p^{5}p^{\frac{n\ell}{2}-2\mu(n)\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n)}{\mu(n)\ell}\big{)}}$. ∎
By [3, Lemma 2.6] (or [14, Lemma II.6]), there are odd positive integers $n_{1},n_{2},\dots$ coprime to $q=p^{\ell}$ such that $\lim\limits_{i\to\infty}\frac{\log_{q}(n_{i})}{\mu(n_{i})}=0$, where $\mu(n)$ is defined in Eq.(2.21). Since $\log_{q}(n_{i})=\log_{p}(n_{i})/\log_{p}(q)=\log_{p}(n_{i})/\ell$, we see that there are odd positive integers $n_{1},n_{2},\dots$ coprime to $p$ such that
$\lim\limits_{i\to\infty}\frac{\log_{p}(n_{i})}{\mu(n_{i})}=0,$ (5.27)
which implies that $\mu(n_{i})\to\infty$, hence $n_{i}\to\infty$. Obviously, we can assume that $\mu(n_{i})\geq\log_{p}(n_{i})$ for $i=1,2,\dots$.
Theorem 5.10.
Assume that $0<\delta<1-q^{-1}$ and $h_{q}(\delta)<\frac{1}{4}$ as in Eq.(5.22). Then there are Hermitian self-dual $2$-quasi-cyclic codes $C_{1},C_{2},\dots$ over $F$ (hence ${\rm R}(C_{i})=\frac{1}{2}$) such that the code length $2n_{i}$ of $C_{i}$ goes to infinity and the relative minimum distance $\Delta(C_{i})>\delta$ for $i=1,2,\dots$.
Proof.
Take $n_{1},n_{2},\dots$ as in Eq.(5.27). There is a positive real number $\varepsilon>0$ such that $\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n_{i})}{\mu(n_{i})\ell}\geq\varepsilon$ for large enough index $i$. So we can further assume that
$\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n_{i})}{\mu(n_{i})\ell}\geq\varepsilon,~{}~{}~{}~{}i=1,2,\dots.$ (5.28)
Taking $n=n_{i}$ in Corollary 5.6 and Lemma 5.9, and denoting ${\cal D}$ by ${\cal D}(i)$, we get
$\frac{|{\cal D}(i)^{\leq\delta}|}{|{\cal D}(i)|}\leq\frac{p^{5}p^{\frac{n_{i}\ell}{2}-2\mu(n_{i})\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n_{i})}{\mu(n_{i})\ell}\big{)}}}{p^{-2}p^{\frac{n_{i}\ell}{2}}}=p^{7}p^{-2\mu(n_{i})\ell\big{(}\frac{1}{4}-h_{q}(\delta)-\frac{\log_{p}(n_{i})}{\mu(n_{i})\ell}\big{)}}.$
By Eq.(5.28) and that $\mu(n_{i})\to\infty$, we get that
$\lim_{i\to\infty}|{\cal D}(i)^{\leq\delta}|\big{/}|{\cal D}(i)|\leq\lim_{i\to\infty}p^{7}p^{-2\mu(n_{i})\ell\varepsilon}=0.$
Therefore, we can further assume that $|{\cal D}(i)^{\leq\delta}|<|{\cal D}(i)|$ for $i=1,2,\dots$. So we can take $g_{i}\in{\cal D}(i)\setminus{\cal D}(i)^{\leq\delta}$. Then $C_{i}=C_{1,g_{i}}$ is a Hermitian self-dual $2$-quasi-cyclic code of length $2n_{i}$ with $\Delta(C_{i})>\delta$. ∎
6 Hermitian (Euclidean) self-dual $2$-quasi
constacyclic codes
[…]
Lemma 6.1.
Assume that ${\lambda^{1+p^{h}}=1}$ and $\gcd(n,t)=1$, where $t={\rm ord}_{F^{\times}}(\lambda)$ as in Eq.(3.1). Then there is a $\gamma=\lambda^{s}\in F^{\times}$ such that $\gamma^{n}=\lambda^{-1}$ and the map
$\eta:~{}{\cal R}\,\longrightarrow\,{\cal R}_{\kern-1.0pt\lambda},\quad\sum_{i=0}^{n-1}a_{i}X^{i}\,\longmapsto\,\sum_{i=0}^{n-1}a_{i}(\gamma X)^{i}~{}({\rm mod}~{}X^{n}-\lambda),$ (6.1)
satisfies the following:
(1) $\eta$ is an algebra isomorphism.
(2) The weight ${\rm w}\big{(}\eta(a(X))\big{)}={\rm w}\big{(}a(X)\big{)}$,   $\forall$ $a(X)\in{\cal R}$.
(3) $\big{\langle}\eta(a(X)),\,\eta(b(X))\big{\rangle}_{h}=\big{\langle}a(X),\,b(X)\big{\rangle}_{h}$,   $\forall$ $a(X),b(X)\in{\cal R}$.
Proof.
Since $\gcd(n,t)=1$, there are integers $s,s^{\prime}$ such that $ns+ts^{\prime}=-1$. We deduce that $\lambda^{-1}=\lambda^{ns+ts^{\prime}}=\lambda^{ns}\lambda^{ts^{\prime}}=\lambda^{ns}$ since $\lambda^{t}=1$. Take $\gamma=\lambda^{s}$, then $\gamma^{n}=\lambda^{-1}$. It is known that (1) and (2) has been proved in [5, Theorem 3.2]. Therefore, it suffices to show that (3).
Let $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}\in{\cal R}$ and $b(X)=\sum_{i=0}^{n-1}b_{i}X^{i}\in{\cal R}$. Then $\eta\big{(}a(X)\big{)}=\sum_{i=0}^{n-1}a_{i}\gamma^{i}X^{i}$ and $\eta\big{(}b(X)\big{)}=\sum_{i=0}^{n-1}b_{i}\gamma^{i}X^{i}$. By Eq.(3.5), we have that
$\big{\langle}\eta(a(X)),\eta(b(X))\big{\rangle}_{h}=\sum_{i=0}^{n-1}(a_{i}\gamma^{i})\cdot(b_{i}\gamma^{i})^{p^{h}}=\sum_{i=0}^{n-1}a_{i}b_{i}^{p^{h}}\cdot\gamma^{(1+p^{h})i}.$
Applying the equality $\gamma=\lambda^{s}$ and the assumption $\lambda^{1+p^{h}}=1$, we get
${\gamma^{1+p^{h}}=(\lambda^{s})^{1+p^{h}}=(\lambda^{1+p^{h}})^{s}}=1.$
Thus
$\big{\langle}\eta(a(X)),\eta(b(X))\big{\rangle}_{h}=\sum_{i=0}^{n-1}a_{i}b_{i}^{p^{h}}=\big{\langle}a(X),b(X)\big{\rangle}_{h},$
which proves (3). ∎
Corollary 6.2.
Assume that ${\lambda^{1+p^{h}}=1}$ and $\gcd(n,t)=1$, where $t={\rm ord}_{F^{\times}}(\lambda)$ as in Eq.(3.1). Let
$\eta^{(2)}:~{}{\cal R}^{2}\,\longrightarrow\,{\cal R}_{\kern-1.0pt\lambda}^{2},~{}~{}\big{(}a(X),a^{\prime}(X)\big{)}\,\longmapsto\,\big{(}\eta(a(X)),\eta(a^{\prime}(X))\big{)},$ (6.2)
where $\eta$ is defined in Lemma 6.1. Then the following hold.
(1) $\eta^{(2)}$ is a module isomorphism.
(2) ${\rm w}\big{(}\eta^{(2)}(a(X),a^{\prime}(X))\big{)}={\rm w}\big{(}a(X),a^{\prime}(X)\big{)}$$\forall\;(a(X),a^{\prime}(X))\in{\cal R}^{2}$.
(3) $\!\big{\langle}\eta^{(2)}(a(X),a^{\prime}(X)),\eta^{(2)}(b(X),b^{\prime}(X))\big{\rangle}_{h}\!=\!\big{\langle}(a(X),a^{\prime}(X)),\!(b(X),b^{\prime}(X))\big{\rangle}_{h}$,  $\forall\;(a(X),a^{\prime}(X)),(b(X),b^{\prime}(X))\in{\cal R}^{2}$.
Proof.
(1)  For $\big{(}a(X),a^{\prime}(X)\big{)}\in{{\cal R}}^{2}$ and $f(X)\in{{\cal R}}$, by Lemma 6.1(1),
$\eta^{(2)}\big{(}f(X)\cdot(a(X),a^{\prime}(X))\big{)}=\big{(}\eta(f(X)a(X)),\eta(f(X)a^{\prime}(X))\big{)}$
$=\big{(}\eta(f(X))\eta(a(X)),\eta(f(X))\eta(a^{\prime}(X))\big{)}=\eta(f(X))\cdot\eta^{(2)}\big{(}a(X),a^{\prime}(X)\big{)}.$
Thus Eq.(6.2) is a module homomorphism. Observe that by Lemma 6.1(1), $\eta^{(2)}$ must be bijective, hence it is a module isomorphism.
(2)  By Lemma 6.1(2), we get
${\rm w}\big{(}\eta^{(2)}(a(X),a^{\prime}(X))\big{)}$ $={\rm w}\big{(}\eta(a(X)),\eta(a^{\prime}(X))\big{)}={\rm w}\big{(}\eta(a(X))\big{)}+{\rm w}\big{(}\eta(a^{\prime}(X))\big{)}$
$={\rm w}\big{(}a(X)\big{)}+{\rm w}\big{(}a^{\prime}(X)\big{)}={\rm w}\big{(}a(X),a^{\prime}(X)\big{)}.$
(3)  For $(a(X),a^{\prime}(X))\in{{\cal R}^{2}}$, $(b(X),b^{\prime}(X))\in{{\cal R}^{2}}$, where $a(X)\!=\sum_{i=0}^{n-1}a_{i}X^{i}$, $a^{\prime}(X)\!=\sum_{i=0}^{n-1}a^{\prime}_{i}X^{i}$ and $b(X)=\sum_{i=0}^{n-1}b_{i}X^{i}$, $b^{\prime}(X)=\sum_{i=0}^{n-1}b^{\prime}_{i}X^{i}$, it is obvious that
$\big{\langle}(a(X),a^{\prime}(X)),(b(X),b^{\prime}(X))\big{\rangle}_{h}$ $=\sum_{i=0}^{n-1}a_{i}{b_{i}}^{p^{h}}+\sum_{i=0}^{n-1}a^{\prime}_{i}{b^{\prime}_{i}}^{p^{h}}$
$=\big{\langle}a(X),b(X)\big{\rangle}_{h}+\big{\langle}a^{\prime}(X),b^{\prime}(X)\big{\rangle}_{h}.$
Then
$\big{\langle}\eta^{(2)}(a(X),a^{\prime}(X)),\;\eta^{(2)}(b(X),b^{\prime}(X))\big{\rangle}_{h}$
$=$ $\big{\langle}\big{(}\eta(a(X)),\eta(a^{\prime}(X))\big{)},\;\big{(}\eta(b(X)),\eta(b^{\prime}(X))\big{)}\big{\rangle}_{h}$
$=$ $\big{\langle}\eta(a(X)),\eta(b(X))\big{\rangle}_{h}+\big{\langle}\eta(a^{\prime}(X)),\eta(b^{\prime}(X))\big{\rangle}_{h}$
[…]
$=$ $\big{\langle}(a(X),a^{\prime}(X)),\;(b(X),b^{\prime}(X))\big{\rangle}_{h}.$
We are done. ∎
Theorem 6.3.
Assume that ${\lambda^{1+p^{h}}=1}$ and $\gcd(n,t)=1$, where $t={\rm ord}_{F^{\times}}\!(\lambda)$ as in Eq.(3.1). Let $\eta^{(2)}$ be as in Eq.(6.2). Then $C$ is an ${\cal R}$-submodule of ${\cal R}^{2}$ if and only if $\eta^{(2)}(C)$ is an ${\cal R}_{\kern-1.0pt\lambda}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$. At that case, the following hold.
(1) The rate ${\rm R}\big{(}\eta^{(2)}(C)\big{)}={\rm R}(C)$.
(2) The relative minimum distance $\Delta\big{(}\eta^{(2)}(C)\big{)}=\Delta(C)$.
(3) $\eta^{(2)}(C)$ is Galois $p^{h}$-self-dual if and only if $C$ is Galois $p^{h}$-self-dual.
Proof.
By Corollary 6.2(1), the map $\eta^{(2)}$ in Eq.(6.2) is a module isomorphism. So $C$ is an ${\cal R}$-submodule of ${\cal R}^{2}$ if and only if $\eta^{(2)}(C)$ is an ${\cal R}_{\kern-1.0pt\lambda}$-submodule of ${\cal R}_{\kern-1.0pt\lambda}^{2}$. Assume that it is this case.
(1) Since $\eta^{(2)}$ is an isomorphism, $\dim_{F}\eta^{(2)}(C)=\dim_{F}C$, and so
${\rm R}\big{(}\eta^{(2)}(C)\big{)}=\frac{\dim_{F}\eta^{(2)}(C)}{2n}=\frac{\dim_{F}C}{2n}={\rm R}(C).$
(2)  It holds by Corollary 6.2(2) obviously.
(3)  Observe that by Corollary 6.2(3), $\big{\langle}\eta^{(2)}(C),\eta^{(2)}(C)\big{\rangle}_{h}=0$ in ${\cal R}_{\kern-1.0pt\lambda}^{2}$ if and only if $\big{\langle}C,\,C\big{\rangle}_{h}=0$ in ${\cal R}^{2}$. Further, applying the above conclusion (1) yields that ${\rm R}\big{(}\eta^{(2)}(C)\big{)}=\frac{1}{2}$ if and only if ${\rm R}(C)=\frac{1}{2}$. Thus (3) holds. ∎
In the following, we consider the asymptotic property of Hermitian (Euclidean) self-dual $2$-quasi $\lambda$-constacyclic codes.
Theorem 6.4.
Assume that $\ell$ is even. The Hermitian self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ are asymptotically good if and only if $\lambda^{1+p^{\ell/2}}=1$.
Proof.
If $\lambda^{1+p^{\ell/2}}\neq 1$, then Hermitian self-dual $2$-quasi $\lambda$-constacyclic codes over $F$ are asymptotically bad, see Theorem 4.5.
Assume that $\lambda^{1+p^{\ell/2}}=1$. Let $0<\delta<1-q^{-1}$ and $h_{q}(\delta)<\frac{1}{4}$. By Theorem 5.10, there are Hermitian self-dual $2$-quasi-cyclic codes $C_{1},C_{2},\dots$ over $F$ such that:
  • the code length $2n_{i}$ of $C_{i}$ satisfy that $n_{i}$ is odd and coprime to $q$ and $\lim\limits_{i\to\infty}\frac{\log_{p}(n_{i})}{\mu(n_{i})}=0$; in particular, $\mu(n_{i})\to\infty$, and hence $n_{i}\to\infty$;
  • the rate ${\rm R}(C_{i})=\frac{1}{2}$ and the relative minimum distance $\Delta(C_{i})>\delta$  for $i=1,2,\dots$.
Let $\tilde{t}=\max\big{\{}\mu(r)\,\big{|}\,\mbox{$r$ runs over the prime divisors of $t$}\big{\}}$, where $t={\rm ord}_{F^{\times}}\!(\lambda)$ as in Eq.(3.1). By [23, Lemma II.2],
$\mu(n_{i})=\min\big{\{}\mu(r)\;|\;\mbox{$r$ runs over the prime divisors of $n_{i}$}\big{\}}.$
If $\gcd(n_{i},t)\neq 1$, then $\mu(n_{i})\leq\tilde{t}$. Since $\mu(n_{i})\to\infty$, there are only finitely many $n_{i}$ such that $\gcd(n_{i},t)\neq 1$. Removing such $n_{i}$, we can further assume that $\gcd(n_{i},t)=1$ for $i=1,2,\dots$. Thus, applying the isomorphism Eq.(6.2) to $C_{i}$ yields $\tilde{C}_{i}=\eta^{(2)}(C_{i})$ in ${\cal R}_{\kern-1.0pt\lambda}^{2}$, $i=1,2,\dots$. We get the code sequence
$\tilde{C}_{1},~{}\tilde{C}_{2},~{}\dots.$
By Theorem 6.3, each $\tilde{C}_{i}$ is Hermitian self-dual $2$-quasi $\lambda$-constacyclic codes over $F$, and the code length $2n_{i}$ goes to infinity, the rate ${\rm R}(\tilde{C}_{i})=\frac{1}{2}$ and the relative minimum distance $\Delta(\tilde{C}_{i})=\Delta(C_{i})>\delta$ for $i=1,2,\dots$. ∎
For Euclidean case, the “Euclidean self-dual” is referred to as “self-dual”.
Theorem 6.5.
[…]
Proof.
By Theorem 4.5, if $\lambda^{2}\neq 1$ then the self-dual $2$-quasi $\lambda$-constacyclic codes are asymptotically bad. In the following we assume that $\lambda=\pm 1$
If $\lambda=1$, it has been shown in [23] (cf. Remark 2.6) that the self-dual $2$-quasi-cyclic codes over $F$ are asymptotically good.
[…]
  • the code length $2n_{i}$ of $C_{i}$ satisfy that $n_{i}$ is odd and coprime to $q$, and $\lim\limits_{i\to\infty}\frac{\log_{p}(n_{i})}{\mu(n_{i})}=0$; in particular, $\mu(n_{i})\to\infty$, and hence $n_{i}\to\infty$;
  • the rate ${\rm R}(C_{i})=\frac{1}{2}$ and the relative minimum distance $\Delta(C_{i})>\delta$  for $i=1,2,\dots$.
Similarly to the proof of Theorem 6.4, we get self-dual $2$-quasi negacyclic codes
$\tilde{C}_{1},~{}\tilde{C}_{2},~{}\dots$
over $F$ such that their code length $2n_{i}$ goes to infinity, ${\rm R}(\tilde{C}_{i})=\frac{1}{2}$ and their relative minimum distance $\Delta(\tilde{C}_{i})=\Delta(C_{i})>\delta$ for $i=1,2,\dots$. ∎
7 Conclusions
The purpose of this paper is to characterize the Galois self-dual $2$-quasi $\lambda$-constacyclic codes, and to investigate their asymptotic properties.
We first showed the algebraic structure of $2$-quasi $\lambda$-constacyclic codes. Then we found that the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes behave much differently according to whether $\lambda^{1+p^{h}}\!\neq 1$ ($\lambda\neq\lambda^{-p^{\ell-h}}$ equivalently) or ${\lambda^{1+p^{h}}\!=1}$ ($\lambda=\lambda^{-p^{\ell-h}}$ equivalently). In both the cases, we exhibited the necessary and sufficient conditions for the $2$-quasi $\lambda$-constacyclic codes being Galois $p^{h}$-self-dual, see Theorem 4.1 (for the former case), Lemma 4.9 and Theorem 4.10 (for the latter case). And in the former case we proved that the Galois $p^{h}$-self-dual $2$-quasi $\lambda$-constacyclic codes are asymptotically bad.
[…]
Acknowledgements
The authors are grateful to the editor and the anonymous referees for taking time to read and comment on the article very carefully and insightfully. Their nice comments and suggestions helped them to improve the article very much.